### v3

# Special Hamiltonian $S^{1}$-actions on symplectic 4-manifolds

Abstract
In this paper we consider symplectic 4-manifolds $(M,\omega)$ with $c_{1}(M,\omega)=0$ which admit a Hamiltonian $S^{1}$-action together with an equivariant Maslov condition on orbits of the group action. We call such spaces special Hamiltonian $S^{1}$-spaces. It turns out that there are no compact special Hamiltonian $S^{1}$-spaces. We classify all exact special Hamiltonian $S^{1}$-spaces and show that all of them admit the structure of a Stein surface. Moreover, we show that any special Hamiltonian $S^{1}$-action on a 1-connected Stein surface $\mathcal{W}_{n}$ is eventually symplectically equivariant to the standard one if $n=0,1$, and at least near the $A_{n}$-string of Lagrangian spheres associated to the standard special Hamiltonian $S^{1}$-action if $n\geq 2$.
1 Introduction and main results
Hamiltonian $S^{1}$-actions on compact symplectic 4-manifolds have been classified by Karshon in [8] where she proved that all these Hamiltonian $S^{1}$-spaces are Kähler. In this article we consider a type of Hamiltonian $S^{1}$-action $G:S^{1}\to Ham(M,\omega)$ on a symplectic 4-manifold $(M,\omega)$ with an extra property called the Maslov condition, which requests that along any nondegenerate orbit $C$, any $G$-invariant Lagrangian subbundle in the symplectic normal bundle of $C$, treated as a loop of Lagrangian planes along $C$, has zero Maslov number. We call such a group action a special Hamiltonian $S^{1}$-action and the triple $(M,\omega,G)$ a special Hamiltonian $S^{1}$-space. See Definition 2.1 in §2 for the precise definition of a special Hamiltonian $S^{1}$-action. To ensure the Maslov condition is permitted we assume that $(M,\omega)$ is connected with the first Chern class $c_{1}(M,\omega)=0$, and that $G$ acts effectively on $M$.
It turns out that the Maslov condition imposes restrictions on the topology of $M$ as well of the $G$-action.
Theorem 1.1.
Let $(M,\omega,G)$ be a connected special Hamiltonian $S^{1}$-space and $G$ acts effectively on $M$ with moment map $h$, then
  1. (i).
    $M$ is not compact (Proposition 3.3), and
  2. (ii).
    the fixed point set $Fix(G)$ of the $G$-action is a discrete set of points (Proposition 3.4).
  3. (iii).
    Moreover, if $(M,\omega=d\lambda)$ is exact then
    1. (a).
      $Fix(G)\subset h^{-1}(0)$ up to addition of a constant to $h$, $h^{-1}(c)$ is connected for $c\in\mathbb{R}$, and
    2. (b).
      the $G$-action is semi-freeI.e., the stabilizer $stab_{p}:=\{g\in G\mid g(p)=p\}$ is $\{1\}$, $\forall p\in M\setminus Fix(G)$.. (Proposition 3.5)
In this paper we classify all connected exact special Hamiltonian $S^{1}$-spaces $(M,\omega=d\lambda,G)$ with finite fixed point set $Fix(G)$. For the 1-connected case we obtain the following classification results.
Theorem 1.2.
Let $(M,\omega,G)$ be a connected, 1-connected exact special Hamiltonian $S^{1}$-space, then the $G$-action is semi-free with $n+1$ fixed points for some integer $n\geq 0$. If $n>0$ then $(M,\omega)$ is symplectically the linear plumbing of $n$ cotangent bundles of spheres $S_{i}\subset M$, $i=1,...,n$, where the union $\cup_{i=1}^{n}S_{i}$ is an $A_{n}$-string of $G$-invariant Lagrangian spheres. If $n=0$ then $(M,\omega)$ is symplectomorphic to the standard symplectic 4-space.
For $n\in\mathbb{N}\cup\{0\}$ Let $\mathcal{W}_{n}\subset\mathbb{C}^{3}$ denote the Stein surface defined by the equation
$z_{1}^{2}+z_{2}^{2}+z_{3}^{n+1}=1.$
Let $\omega_{n}$ denote the standard symplectic structure on $\mathbb{C}^{3}$ restricted to $\mathcal{W}_{n}$. It is known that for $n>0$ $\mathcal{W}_{n}$ is topologically the plumbing of $n$ copies of cotangent bundles $T^{*}S^{2}$ of type $A_{n}$ [16], and $\mathcal{W}_{0}$ is symplectomorphic to the standard symplectic 4-space.
Proposition 1.3.
Let $\tilde{\textsf{G}}$ denote the $S^{1}$-action on $\mathbb{C}^{3}$ defined by
$(z_{1},z_{2},z_{3})\to(z_{1}\cos\theta-z_{2}\sin\theta,z_{1}\sin\theta+z_{2}\cos\theta,z_{3}),\quad\theta\in\mathbb{R}/2\pi\mathbb{Z}.$
Then $\tilde{\textsf{G}}$ acts on $(\mathcal{W}_{n},\omega_{n})$ as a special Hamiltonian $S^{1}$-action with $n+1$ Fixed points.
Theorem 1.4.
A connected, 1-connected, exact special Hamiltonian $S^{1}$-space $(M,\omega=d\lambda,G)$ with $n+1$ fixed points is symplectomorphic to the Stein surface $(\mathcal{W}_{n},\omega_{n})$. In particular, the symplectic topology of $(M,\omega)$ is determined by the cardinality of the fixed point set $Fix(G)$.
Moreover, together with the following theorem by Wu [16], any $A_{n}$-string of Lagrangian spheres in $(M,\omega)$ is Hamiltonian isotopic to any given $A_{n}$-string of Lagrangian spheres associated to the $G$-action as in Theorem 1.4, up to a composition of Lagrangian Dehn twists along the latter:
Theorem 1.5 ([16]).
[…]
Let G denote the linear $S^{1}$-action of $\tilde{\textsf{G}}$ restricted to $\mathcal{W}_{n}$. In Section 5 we will construct an explicit G-invariant $A_{n}$-string of Lagrangian spheres $\textsf{S}:=\cup_{j=1}^{n}\textsf{S}_{j}$ for in $(\mathcal{W}_{n})$. Theorem 1.4 and Theorem 1.5 together implies that up to a symplectomorphism we can identify $(M,\omega,G)$ with $(\mathcal{W}_{n},\omega_{n},G)$ and S in $\mathcal{W}_{n}$ is preserved by the $S^{1}$-action of $G$. The classification of 1-connected special Hamiltonian $S^{1}$-space is then reduced to the classification of semi-free special Hamiltonian $S^{1}$-actions on $(\mathcal{W}_{n},\omega_{n})$.
We have the following result concerning the uniqueness of special Hamiltonian $S^{1}$-actions on $\mathcal{W}_{n}$:
Theorem 1.6.
Let $G$ be any semi-free special Hamiltonian $S^{1}$-action on $(\mathcal{W}_{n},\omega_{n})$.
  1. (i).
    The $S^{1}$-action of $G$ is symplectically equivalent to the $S^{1}$-action of G on some open neighborhood of the $A_{n}$-string of Lagrangian spheres associated to G. I.e., there is a $G$-invariant open neighborhood $U\subset\mathcal{W}_{n}$ of S (resp. the fixed point $(0,0,1)\in\mathcal{W}_{0}$ if $n=0$) and a $C^{\infty}$-embedding $\EuFrak{e}_{\textsf{S}}:U\to W_{n}$ on which $G=\EuFrak{e}_{\textsf{S}}\circ\textsf{G}\circ\EuFrak{e}_{\textsf{S}}^{-1}$ and hence $h=\textsf{h}\circ\EuFrak{e}_{\textsf{S}}^{-1}$, where $h$ and h are the moment maps of $G$ and G respectively.
  2. (ii).
    If $n=0$, $G$ is symplectically equivariant to G on any compactly supported $G$-invariant open domain containing the fixed point $(0,0,1)$ of the G-action on $\mathcal{W}_{0}$.
  3. (iii).
    If $n=1$, then $G$ is symplectically equivariant to G on any compactly supported $G$-invariant open domain containing the $A_{1}$-sphere $S\cong S^{2}$ associated to the G-action on $\mathcal{W}_{1}$.
For the case that $(M,\omega=d\lambda)$ is not 1-connected we assume that $Fix(G)$ is not empty (otherwise the $G$-action is free and the Maslov condition is vacuous). Then the reduced space $M_{red}$ is topologically an oriented noncontractible surface of genus $g\geq 0$ with $b\geq 1$ boundary components, $g+b\geq 2$, and $k:=|Fix(G)|=n+1\geq 1$ marked points.
Theorem 1.7.
Let $(M,\omega,G)$ be a connected exact special Hamiltonian $S^{1}$-space with $c_{1}(M,\omega)=0$ and $n+1$ fixed points, $n\geq 0$. Assume also that $M$ is not 1-connected. Then $M$ admits the structure of a Stein surface, which is topologically obtained by attaching $(2g+b-1)$-pairs of 1- and Lagrangian 2-handlebodies to the boundary of a Stein domain $W_{n}\subset\mathcal{W}_{n}$ diffeomorphic to $\mathcal{W}_{n}$. The homology groups of $M$ are
$H_{m}(M,\mathbb{Z})=\begin{cases}0&\text{ for }m>2,\\ \mathbb{Z}^{2g+b-1+n}\cong H_{2}(W_{n},\mathbb{Z})\oplus\mathbb{Z}^{2g+b-1}&\text{ for }m=2,\\ \mathbb{Z}^{2g+b-1}\cong H_{1}(M_{red},\mathbb{Z})&\text{ for }m=1.\end{cases}$
In particular, $H_{2}(M,\mathbb{Z})$ is generated by $n$ Lagrangian spheres and $12g+b-$ 1 monotone Lagrangian tori.
Remark 1.8 (Motivation on Maslov condition and relation with complexity one Hamiltonian space).
A motivation to study 4-dimensional special Hamiltonian $S^{1}$-spaces is to explore Hamiltonian $S^{1}$-spaces which possess monotone Lagrangian surfaces invariant under the given Hamiltonian $S^{1}$-actions. It turns out that the 4-dimensional exact special Hamiltonian $S^{1}$ spaces $(M,\omega=d\lambda,G)$ that we consider here are centered complexity one spaces as defined in [9] (see also [10, 11, 12]) except that we do not demand the moment maps of these spaces to be proper, as was requested in [9]. Ignoring the condition on properness, the Maslov condition that we imposed here provides a concrete subclass of centered complexity one spaces which contain various types of immersed monotone Lagrangian surfaces, including spheres, tori, and Whitney spheres (immersed spheres with one transversal self-intersection point), etc., all contained in the 00-level set of the moment map associated to the special Hamiltonian $S^{1}$ action. See Remark 4.2 after Proposition 4.1. It will be interesting to see how these Lagrangian surfaces, in addition to their topological types, are classified up to Hamiltonian isotopy.
The rest of this paper is organized as follows: In §2 we give the definition of a special Hamiltonian $S^{1}$-action on a symplectic 4-manifold $(M,\omega)$. As an example we show that an $S^{1}$-subgroup of $SU(2)$ which acts linearly on the standard symplectic 4-space $(\mathbb{R}^{4},\omega_{0})$ meets the special Hamiltonian criteria. Proof of Theorem 1.1 is given in §3 where we also establish various symplectic topological properties of special Hamiltonian $S^{1}$-spaces. Theorem 1.2 is proved in §4. Proposition 1.3 and Theorem 1.4 are verified in §5 where the $\tilde{\textsf{G}}$-action is analyzed and an $A_{n}$-string of Lagrangian spheres is given. This is followed by the proof of Theorem 1.6. Theorem 1.7 is proved in §6.
2 Special Hamiltonian $S^{1}$-action
Maslov index of a loop of Lagrangian planes (see e.g. Ch.2 of [13]).  Let $L$ be a Lagrangian plane field defined along an oriented closed curve $C$ in a symplectic 4-manifold $(M,\omega)$. Assume that there exists a smooth map $\psi:D\to M$ from the unit 2-disc $D$ into $M$ such that $C=\psi(\partial D)$. Fix a symplectic trivialization $\Phi$ of the pullback bundle $\psi^{*}TM$ over $D$, with which the Lagrangian bundle $\psi^{*}L$ over $\partial D$ can be identified with a loop of Lagrangian planes in the symplectic space $\mathbb{R}^{4}\cong\mathbb{C}^{2}$ or equivalently, the image of a map
$\tau:S^{1}\to\Lambda(\mathbb{C}^{2}),$
where $\Lambda(\mathbb{C}^{2})$ is the space of unoriented Lagrangian planes in $\mathbb{C}^{2}$.
The unitary group $U(2)$ acts transitively on $\Lambda(\mathbb{C}^{2})$ with isotropy group $O(2)$, i.e., $\Lambda(\mathbb{C}^{2})\cong U(2)/O(2)$. By picking a unitary basis field $L$ can be represented by a loop (or half of a loop if $L$ not orientable) of unitary matrices $A$ in $U(2)$. The Maslov index $\mu(\tau)=\mu(L,\Phi)$ is then defined to be the degree of the composed map
[…]
Several remarks on $\mu(L,\Phi)$ are in order:
  1. (i).
    If $M$ is 1-connected then $\mu(L,\Phi)$ is always defined since $C$ is the boundary of some disc.
  2. (ii).
    $\mu(L,\Phi)$ is independent of the choice of $\Phi$.
  3. (iii).
    If $c_{1}(M,\omega)=0$ Then $\mu(L,\Phi)$ is independent of the class $[\psi]\in\pi_{2}(M,C)$ and hence the choice of $\psi$.
Definition 2.1 (Special Hamiltonian $S^{1}$-action).
Let $G$ be an $S^{1}$-group of Hamiltonian diffeomorphisms acting on $M$ effectively with nonempty fixed point set $Fix(G)$. Let $h\in C^{\infty}(M,\mathbb{R})$ denote the moment map of $G$, and $X=X_{h}$ the Hamiltonian vector field on $M$ which generates the $G$-action so that
$\omega(X,\cdot)=-dh.$
$h:M\to\mathbb{R}$ is also called a Hamiltonian function of $(M,\omega)$. For $p\not\in Fix(G)$ let $C:=Orb_{G}(p)$ denote the $G$-orbit of $p$. Take a nonzero $v\in T_{p}M/T_{p}C$ so that the plane $X|_{p}\wedge v$ spanned by $X|_{p}$ and $v$ is Lagrangian. Extend $X|_{p}\wedge v$ to an oriented $G$-invariant Lagrangian plane field $L_{p}$ along $C$ via the $G$-action. The Maslov index $\mu(L_{p})$ is independent of the choice of nonzero $v\in(T_{p}M/T_{p}C)$ with $\omega_{p}(X,v)=0$. Also $\mu(L_{p})=\mu(L_{q})$ for $q\in Orb_{G}(p)$. We say that the $G$ is special if
$\mu(L_{p})=0,\quad p\not\in Fix(G).$
Example 2.2.
Let $M=\mathbb{R}^{4}=\mathbb{C}^{2}$ with the standard symplectic structure $\omega_{0}=\sum_{j=1}^{2}dx_{j}\wedge dy_{j}=d\lambda_{0}$, $\lambda_{0}=\frac{1}{2}(\sum_{j=1}^{2}x_{j}dy_{j}-y_{j}dx_{j})$ and the standard Euclidean inner product. We claim that
$\textsf{G}:=\Big{\{}\textsf{G}_{\theta}=\begin{pmatrix}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{pmatrix}\Big{|}\theta\in\mathbb{R}/2\pi\mathbb{Z}\Big{\}}\subset SU(2)$
is a semi-free special Hamiltonian $S^{1}$-group acting on $\mathbb{C}^{2}$ with $Fix(\textsf{G})=\{0\}$.
It is easy to see that this G-action is semi-free with $Fix(\textsf{G})=\{0\in\mathbb{C}^{2}\}$, and is Hamiltonian with moment map
$\textsf{h}:\mathbb{C}^{2}\to\mathbb{R},\quad\textsf{h}(z_{1},z_{2})=\Im(z_{1}\bar{z}_{2})=x_{2}y_{1}-x_{1}y_{2},$
and the corresponding Hamiltonian vector field is
$X=-x_{2}\partial_{x_{1}}+x_{1}\partial_{x_{2}}-y_{2}\partial_{y_{1}}+y_{1}\partial_{y_{2}}.$
Maslov condition.  We claim that G is special: Since the set of nontrivial orbits of G is connected and the Maslov index is discrete, it suffices to show that the Maslov condition holds for a typical orbit of G, say the loop $C:=Orb_{\textsf{G}}(p)$ with $p=(1,0)\in\mathbb{C}^{2}$,
$C(t)=(\cos t,\sin t)\subset\mathbb{C}^{2},t\in\mathbb{R}/2\pi\mathbb{Z}.$
Then $\dot{C}(0)=X_{\textsf{h}}|_{p}=\partial_{x_{2}}$. Take $v=-\partial_{x_{1}}|_{p}\in T_{p}\mathbb{C}^{2}$. Apply G to $X_{\textsf{h}}|_{p}\wedge v=\partial_{x_{2}}\wedge(-\partial_{x_{1}})|_{p}$. The resulting Lagrangian plane field along $C$ is $L:=\partial_{x_{1}}\wedge\partial_{x_{2}}$. The pair $(\partial_{x_{1}},\partial_{x_{2}})$ is a unitary basis field of the tangent bundle $T\mathbb{C}^{2}$, also a unitary basis field of $L$. Therefore under the map $\tau$ as in (1) $L$ can be represented by the constant loop $\begin{bmatrix}1&0\\ 0&1\end{bmatrix}\subset U(2)$ with constant determinant $1$, hence the Maslov index of $L$ is $\mu(L)=0$. This confirms that the G-action on $\mathbb{C}^{2}$ is special.
Morse index of h at 00.  Note that $d\textsf{h}$ vanishes at $\{0\}=Fix(\textsf{G})$. The Hessian matrix of h at the point 00, with respect to the basis $(\partial_{x_{1}},\partial_{x_{2}},\partial_{y_{1}},\partial_{y_{2}})$, is
$\text{Hess}_{0}(\textsf{h})=\begin{bmatrix}0&0&0&-1\\ 0&0&1&0\\ 0&1&0&0\\ -1&0&0&0\end{bmatrix},$
which is nondegenerate with eigenvalues $-1,-1,1,1$, so h is a Morse function on $\mathbb{C}^{2}$, and the Morse index of h at $0\in\mathbb{C}^{2}$ is 2. The corresponding (-1)- and 1-eigenspaces are
$E_{-1}=Span\{\partial_{x_{1}}+\partial_{y_{2}},\ \partial_{y_{1}}-\partial_{x_{2}}\},\quad E_{1}=Span\{\partial_{x_{1}}-\partial_{y_{2}},\ \partial_{y_{1}}+\partial_{x_{2}}\}$
respectively. Also the two 2-dimensional vector spaces $E_{-1}$ and $E_{1}$ can be identified with a pair of complex lines (and hence $\omega_{0}$-symplectic planes) in $\mathbb{C}^{2}$ intersecting transversally at $0\in\mathbb{C}^{2}$:
$E_{-1}=\{z_{1}+iz_{2}=0\},\quad E_{1}=\{z_{1}-iz_{2}=0\}.$
[…]
$\nabla\textsf{h}=-y_{2}\partial_{x_{1}}+x_{2}\partial_{y_{1}}+y_{1}\partial_{x_{2}}-x_{1}\partial_{y_{2}}.$
[…]
$W_{\textsf{h}}^{s}(0)=E_{-1}=\{z_{1}+iz_{2}=0\},\quad W_{\textsf{h}}^{u}(0)=E_{1}=\{z_{1}-iz_{2}=0\}.$
[…]
Remark 2.3.
For $m\in\mathbb{Z}\setminus\{0\}$
$\textsf{G}^{m}:=\Big{\{}\textsf{G}_{m\theta}=\begin{pmatrix}\cos m\theta&-\sin m\theta\\ \sin m\theta&\cos m\theta\end{pmatrix}\Big{|}\theta\in\mathbb{R}/(2\pi/m)\mathbb{Z}\Big{\}}\subset SU(2)$
is a special Hamiltonian $S^{1}$-group acting on $\mathbb{C}^{2}$ with weight $m$, $Fix(\textsf{G}^{m})=\{0\}$, but it is not semi-free on $\mathbb{C}^{2}$ as $stab_{p}=\mathbb{Z}_{|m|}$ if $p\neq 0$.
3 Some topological constraints imposed by Maslov condition
It turns out that at every isolated fixed point $p\in Fix(G)$ of a special Hamiltonian $S^{1}$-space $(M,\omega,G)$, the $G$-action is Hamiltonian isotopic to the standard G-action in a Darboux chart of $p$, up to a weight $m_{p}\in\mathbb{Z}\setminus\{0\}$.
Proposition 3.1.
Let $(M,\omega,G)$ be a special Hamiltonian $S^{1}$-space with moment map $h$, and $p\in Fix(G)$ an isolated fixed point of the $G$-action.
  1. (i).
    At $p$ the $G$-action induces a linear symplectic $S^{1}$-action on the tangent space $T_{p}M$ which can be identified with $\textsf{G}\subset SU(2)$ up to a weight factor $m\in\mathbb{Z}\setminus\{0\}$, with respect to a suitable symplectic Darboux chart of $p\in M$.
  2. (ii).
    There is a $G$-invariant chart $(U,x^{\prime}_{1},x^{\prime}_{2},y^{\prime}_{1},y^{\prime}_{2})$ centered at $p$ and $G$-equivariant with respect to a linear action of $G$ on $\mathbb{R}^{4}\cong\mathbb{C}^{2}$ such that $\omega|_{U}=\sum_{k=1}^{2}dx^{\prime}_{k}\wedge dy^{\prime}_{k}$.
  3. (iii).
    All isolated critical points of $h$ are nondegenerate with Morse index $2$.
Proof.
(i). For simplicity we assume that $G$ acts effectively near $p$ and with weight ±$\pm 1$. Parametrize $G$ by $t\in\mathbb{R}/2\pi\mathbb{Z}$ so that $G_{0}$ is the identity map. Fix $p\in Fix(G)$. By Darboux-Weinstein Theorem we can symplectically identify a small open neighborhood $U\subset M$ of $p$ with a open neighborhood of the origin $0\in(\mathbb{R}^{4},\omega_{0})$ of the standard symplectic $\mathbb{R}^{4}$ via a local diffeomorphism $\phi:(U,\omega)\to(\phi(U)\subset\mathbb{R}^{4},\omega_{0})$ with $\phi(p)=0$ and $\phi^{*}\omega_{0}=\omega$. Then the linear map $A_{t}:=(G_{t})_{*}|_{p}:T_{p}M\to T_{p}M$ is symplectic, i.e., $A_{t}\in Sp(4,\mathbb{R})$ for all $t$, with $A_{0}$ being the identity map. Since $G_{t}\circ G_{s}=G_{t+s}=G_{s}\circ G_{t}$ for all $s,t$, we have $A_{t}A_{s}=A_{s}A_{t}$ for all $s,t$ as a result. I.e., $A_{t}$, $t\in\mathbb{R}/2\pi\mathbb{Z}$, is an $S^{1}$-subgroup of $Sp(4,\mathbb{R})$. Up to conjugation (or equivalently a choice of $\omega_{0}$-compatible complex structure) we may assume that $A$ is an $S^{1}$-subgroup of the unitary group $U(2)=Sp(4,\mathbb{R})\cap O(4,\mathbb{R})$ with respect to the standard complex structure $J_{0}$ on $\mathbb{R}^{4}\cong\mathbb{C}^{2}$. Now $A$ is either the centralizer $S^{1}$-group $C$ of $U(2)$ or its inverse, or an $S^{1}$-subgroup of $SU(2)$.
If $A=C$, then $A=\Big{\{}A_{t}=\begin{bmatrix}e^{it}&0\\ 0&e^{it}\end{bmatrix}\Big{|}t\in\mathbb{R}/2\pi\mathbb{Z}\Big{\}}$. The linear $A$-action is generated by the Hamiltonian vector field $X^{A}:=\sum_{k=1}^{2}(-y_{k}\partial_{x_{k}}+x_{k}\partial_{y_{k}})$. To check the Maslov condition, pick a point $q\not\in Fix(A)$ say, $q=(-i,0)\in\mathbb{C}^{2}$ where $X^{A}_{q}=\partial_{x_{1}}$. Let $v_{q}=\partial_{x_{2}}$. The Lagrangian plane field $L=Orb_{A}(\partial_{x_{1}}\wedge\partial_{x_{2}})$ along $\gamma:=Orb_{A}(q)$ is $L(t)=(\cos t\partial_{x_{1}}+\sin t\partial_{y_{1}})\wedge(\cos t\partial_{x_{2}}+\sin t\partial_{y_{2}})$, $t\in\mathbb{R}/2\pi\mathbb{Z}$. With its unitary basis field $(\cos t\partial_{x_{1}}+\sin t\partial_{y_{1}},\cos t\partial_{x_{2}}+\sin t\partial_{y_{2}})$ $L$ can be represented by the loop of unitary matrices $\begin{bmatrix}e^{it}&0\\ 0&e^{it}\end{bmatrix}$. By (1) the Maslov index of $L$ is $\mu(L)=4\neq 0$. So $A\neq C$. Similarly $A\neq C^{-1}$ as otherwise we would have $\mu(Orb_{A})(\partial_{x_{1}}\wedge\partial_{x_{2}})=-4\neq 0$. So $A$ is an $S^{1}$-subgroup of $SU(2)$. Since all $S^{1}$-subgroups of $SU(2)$ are conjugate we have $A=\textsf{G}$ up to a change of the complex coordinates of $\mathbb{C}^{2}$ via a special unitary linear map given by some element of $SU(2)$.
As the set $Fix(G)$ is discrete without accumulation points we may choose for each $p\in Fix(G)$ a symplectic Darboux chart $U_{p}$ such that $U_{p}\cap U_{q}=\emptyset$ for $p,q\in Fix(G)$ with $p\neq q$. The above result about $A$ applies to every $p\in Fix(G)$ as a result. This finishes the proof of (i).
(ii) It follows from (i) that an equivariant version of Darboux Theorem [1] (see also [6] Theorem 22.2) applies to every isolated fixed point $p\in Fix(G)$. Namely, for each isolated fixed point $p\in Fix(G)$, there exists a $G$-invariant neighborhood $U$ of $p\in M$ and a $G$-equivariant diffeomorphism $f$ of $U$ into $(M,\omega)$ with $f(p)=p$, $f^{*}\omega=\omega_{0}$ is the standard symplectic form $\omega_{0}=\sum_{k=1}^{2}dx_{k}\wedge dy_{k}$ on $\mathbb{R}^{4}$, and $f^{-1}\circ G\circ f=\textsf{G}\subset SU(2)$. A proof of (ii) is given below.
Let $p\in Fix(G)$ be an isolated fixed point. By Darboux theorem we may identify $\omega$ with the standard symplectic form on $\omega_{0}:=\sum_{j=1}^{2}dx_{j}\wedge dy_{j}$ on $\mathbb{R}^{4}$ when restricted to some Darboux chart $U\subset M$ centered at $p$, on which $\omega=\omega_{0}$ is $G$-invariant.
For simplicity we assume that the $G$-action near $p$ is semi-free with $m_{p}=1$. Then the linear special Hamiltonian $S^{1}$ action $G^{L}=G_{*}:T_{p}M\to T_{p}M$ is
$G^{L}=\Big{\{}G^{L}_{\theta}:=\begin{pmatrix}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{pmatrix}\Big{|}\ \theta\in\mathbb{R}/2\pi\mathbb{Z}\Big{\}}=\textsf{G}\subset SU(2)$
up to conjugation. Identify $(T_{p}M,\omega_{p})$ with $(\mathbb{R}^{4},\omega_{0})$ and let $g_{0}$ denote the standard Euclidean metric on $(T_{p}M,\omega_{p})\simeq(\mathbb{R}^{4},\omega_{0})$. $g_{0}$ is $\omega_{0}$-compatible and is invariant under the G-action on $\mathbb{R}^{4}$. Also let $g$ denote a $\omega$-compatible Riemannian metric on $M$ which is also $G$-invariant at least on a $G$-invariant compact domain $D\subset M$ with $p$ as its central interior point, and $D\cap Fix(G)=\{p\}$ We may assume that $g=g_{0}$ when restricted to $T_{p}M$.
Consider the exponential map with respect to the Riemannian metric $g$ on $M$:
$\EuFrak{e}_{p}:=\exp_{p}:T_{p}M\cong\mathbb{R}^{4}\to M,\quad\EuFrak{e}_{p}(v):=\gamma_{v}(1)$
where $\gamma_{v}(t)$, $t\in\mathbb{R}$, is the geodesic emanated from $p$ with $\gamma_{v}(0)=p$, $\dot{\gamma}(0)=v$. There exists $\epsilon>0$ such that
$\EuFrak{e}_{p}:\ V:=\{v\in T_{p}M\mid|v|<\epsilon\}\to\EuFrak{e}_{p}(v)=\gamma_{v}(1)$
is a diffeomorphism from $V$ onto $U:=\EuFrak{e}_{p}(V)\subset M$. For $v_{0}\in V\subset T_{p}M$, $v_{0}\neq 0$, let $v_{\theta}:=G^{L}_{\theta}(v_{0})$, $|v_{\theta}|=|v_{0}|$ with respect to both $g_{0}$ and $g$. Since $g$ is $G$-invariant on $U$, the $G$-action preserves the space of $g$-geodesics $\gamma_{v}$, $v\in V$, with $\gamma_{v_{\theta}}(t)=\gamma_{tv_{\theta}}(1)=G_{\theta}(\gamma_{tv_{0}}(1))=G_{\theta}(\gamma_{v_{0}}(t))=G_{\theta}(\gamma_{tv_{0}}(1))$, $|t|\leq 1$, $\theta\in\mathbb{R}/2\pi\mathbb{Z}$.
Consider $\EuFrak{e}^{-1}_{p}\circ G\circ\EuFrak{e}_{p}:V\to V$. At $v_{0}\in V$,
$\EuFrak{e}^{-1}_{p}\circ G_{\theta}\circ\EuFrak{e}_{p}(v_{0})=\EuFrak{e}^{-1}_{p}\circ G_{\theta}\circ\gamma_{v_{0}}(1)=\EuFrak{e}^{-1}_{p}(\gamma_{v_{\theta}}(1))=v_{\theta}=G^{L}_{\theta}(v),$
$\Rightarrow\quad\EuFrak{e}^{-1}_{p}\circ G\circ\EuFrak{e}_{p}=\textsf{G}\quad\text{ on }V$
or, equivalently, $G=\EuFrak{e}_{p}\circ\textsf{G}\circ\EuFrak{e}_{p}^{-1}$ on $U$. Also let $\EuFrak{e}_{p}^{0}$ denote the exponential map with respect to $g_{0}$.
WLOG we may identify $(V,\omega_{0})$ with $(U_{0}:=\EuFrak{e}_{p}^{0}(V),\omega_{0})$. Then $\EuFrak{e}_{p}:U_{0}\to U$ is a diffeomorphism from $(U_{0},\omega_{0})$ onto $(U,\omega_{0})$. Let
$\omega_{1}:=(\EuFrak{e}_{p}^{-1})^{*}\omega_{0}=\sum_{j=1}^{2}x^{\prime}_{j}\wedge dy^{\prime}_{j},\quad x^{\prime}_{j}=x_{j}\circ\EuFrak{e}_{p}^{-1},\quad y^{\prime}_{j}=y_{j}\circ\EuFrak{e}_{p}^{-1}.$
Up to the difffeomorphism $\EuFrak{e}_{p}:(U_{0},\omega_{0})\to(U,\omega_{0})$ $G$ acts on $U$ symplectically and linearly with respect to the symplectic form $\omega_{1}$, with symplectic basis $\{\frac{\partial}{\partial x^{\prime}_{1}},\ \frac{\partial}{\partial y^{\prime}_{1}},\ \frac{\partial}{\partial x^{\prime}_{2}},\ \frac{\partial}{\partial y^{\prime}_{2}}\}$. Now
$G=\EuFrak{e}_{p}\circ\textsf{G}\circ\EuFrak{e}^{-1}_{p},$
[…]
$(X:=(\EuFrak{e}_{p})_{*}\textsf{X},\ h:=\textsf{h}\circ\EuFrak{e}^{-1}_{p})$
are the the corresponding Hamiltonian vector field and Hamiltonian function associated to the $G$-action on $U$.
We may extend $\EuFrak{e}_{p}$ to a diffeomorphism $\EuFrak{e}^{\prime}_{p}:M\to M$ from $M$ onto $M$ with $G$-invariant compact support containing $U_{0}\cup U$. Let $G^{\prime}:=\EuFrak{e}^{\prime}_{p}\circ\textsf{G}\circ(\EuFrak{e}_{p}^{\prime})^{-1}$. $G^{\prime}$ acts as a Hamiltonian $S^{1}$-action $M$, with $G^{\prime}=G$ on $U$ and outside the compact support of $\EuFrak{e}^{\prime}_{p}$.
In addition, we claim that there exists a smooth isotopy $\phi_{t}:U\to M$, $t\in[0,1]$, with $\phi_{0}=id$, $\phi_{t}(p)=p$ for all $t$, and $\phi_{1}^{*}\omega_{1}=\omega_{0}$.
Let $\omega_{t}:=(1-t)\omega_{0}+t\omega_{1}$, $t\in[0,1]$, then $\omega_{t}|_{p}=\omega_{0}|_{p}$ for all $t$. By shrinking $U$ if necessary we may assume that $\omega_{t}$ are symplectic on $U$ for all $t\in[0,1]$. Since the $G$-action is special Hamiltonian, $\omega_{t}$ are exact for $t\in[0,1]$,
$\omega_{t}=d\alpha_{t},\quad\frac{d\omega_{t}}{dt}=\omega_{1}-\omega_{0}=d(\alpha_{1}-\alpha_{0})$
for some $\alpha_{t}\in\Omega^{1}(M)$. The $G$-action preserves both $\omega_{1}$ and $\omega_{0}$ on $U$, hence all $\omega_{t}$ are $G$-invariant. Since $G$ is compact, by averaging we may assume that all $\alpha_{t}$ are $G$-invariant at least on $U$ by shrinking $U$ if necessary.
Assume there exists an isotopy $\phi_{t}$, $t\in[0,1]$, $\phi_{0}=id$, such that $\phi^{*}_{t}\omega_{t}=\omega_{0}$, and let $X_{t}$ be the vector field associated to $\phi_{t}$, then
$0=\frac{d}{dt}\Big{(}\phi^{*}_{t}\omega_{t}\Big{)}=\phi^{*}_{t}\Big{(}\frac{d\omega_{t}}{dt}+L_{X_{t}}\omega_{t}\Big{)}=\phi^{*}_{t}d\Big{(}\alpha_{1}-\alpha_{0}+\omega_{t}(X_{t}.\cdot))\Big{)}$
$\Rightarrow\omega_{t}(X_{t},\cdot)=\alpha-\alpha_{0}$ (2)
As $\omega_{t}$’s are symplectic on $U$, for each $t\in[0,1]$ there exists a unique vector field $X_{t}$ defined on $U$ with (2) satisfied . Both $\omega_{t}$ and $\alpha_{t}$ are $G$-invariant on $U$, so are $X_{t}$. Moreover, $X_{t}=0$ at $p$ since $\omega_{t}=\omega_{0}$ at $p$ for $t\in[0,1]$. Then $\phi_{t}:U\to M$ are defined and equivariant with respect to $G$ on some $G$-invariant neighborhood $U_{p}\subset U$ of $p$, with $\phi_{0}=id$, $\phi_{1}^{*}\omega_{1}=\omega_{0}$, $\phi_{1}\circ\textsf{G}\circ\phi_{1}^{-1}=G$. This finishes the proof of (ii).
For (iii), observe that $dh|_{p}=-\omega(X_{h}|_{p},\cdot)$ for $p\in M$. Since $\omega$ is nondegenerate, $\omega(X_{h}|_{p},\cdot)=0$ iff $X_{h}$ vanishes at $p$, i.e., iff $p\in Fix(G)$, i.e., $Crit(h)=Fix(G)$. By (ii) we may assume that $=\textsf{G}$ G and hence $=\textsf{h}$ h near each isolated critical point $p\in Fix(G)$ of $h$, therefore all isolated critical points of $h$ are nondegenerate with Morse index $2$. ∎
Forgetting the Maslov condition, 4-dimensional special Hamiltonian $S^{1}$-spaces are by themselves (not necessarily compact) Hamiltonian $S^{1}$-spaces. Recall that 4-dimensional compact Hamiltonian $S^{1}$-spaces have been classified by Karshon [8]:
Theorem 3.2 ([8], Theorem 6.3).
Every compact 4-dimensional Hamiltonian $S^{1}$-space $M$ can be obtained by a sequence of $S^{1}$-equivariant symplectic blow-ups from one of the following two types of ”minimal model spaces” at their fixed points:
  1. (i).
[…]
  2. (ii).
    a ruled manifold, i.e. a smooth manifold which is topologically an $S^{2}$-bundle over a closed surface, with a compatible $S^{1}$-action that fixes the base surfaces and rotates the fibers ([8] Definition 6.13, Lemma 6.15). A ruled manifold admits a compatible Kähler structure, has two fixed surfaces and no interior fixed points .
Recall that if $(M,\omega,G)$ is a 4-dimensional compact Hamiltonian $S^{1}$-space then each connected component of $Fix(G)$ is either a single point or a symplectic surface, and the maximum and minimum of the moment map $h$ is each attained on exactly one component of $Fix(G)$ (see e.g. [8] Appendix A).
The following two propositions together verify Theorem 1.1.
Proposition 3.3.
A closed compact symplectic 4-manifold $(M,\omega)$ does not admit a special Hamiltonian $S^{1}$-action.
Proof.
Assume in contrary that $(M,\omega)$ allows a special Hamiltonian $S^{1}$-action $G:S^{1}\to Ham(M,\omega)$. We may assume that the action is effective. Let $h$ denote its moment map. Then $Fix(G)=Crit(h)$. $M$ is compact, $h$ is bounded, and the sets $M_{-}:=\{p\in M\mid h(p)=\min h\}$ and $M_{+}:=\{p\in M\mid h(p)=\max h\}$ are both nonempty. Since every isolated fixed point of the $G$-action has Morse index 2, the set $M_{-}$ contains no connected component with isolated fixed points, and neither does $M_{+}$. Then both $M_{-}$ and $M_{+}$ are connected compact closed symplectic surfaces and, following [8], $M$ can be obtained from a ruled manifold with two fixed surfaces and no interior fixed points by a sequence of equivalent blowups at fixed points that are not minima for the moment map of the ruled manifold.
Suppose that $M$ is a ruled manifold, then the orbit $\gamma_{q}:=Orb_{G}(q)$ of $q\not\in Fix(G)$ is contained some symplectic $S^{2}$-fiber $S$ of $M$. But $c_{1}(M)([S])=S\cdot S+e(S)=0+2\neq 0$, $c_{1}(M.\omega)\neq 0$. Indeed, let $F_{p}$ denote the $S^{2}$ fiber over $p\in M_{-}$. Fix any $q\in F_{p}\setminus(M_{+}\cup M_{-})$. $Orb_{G}(q)$ divides $F_{p}\cong S^{2}$ into two discs $D_{-}\cup D_{+}$ with $D_{-}\cap M_{-}=\{p\}$. We may identity a $G$-invariant symplectic Darboux neighborhood of $p\in M$ as $D^{\prime}\times D_{\epsilon}\subset\mathbb{C}_{z_{1}}\times\mathbb{C}_{z_{2}}$ so that $D_{-}\subset D^{\prime}\times\{0\}$, and $D_{\epsilon}$ is a coordinate chart of $F_{p}$ centered at $p$. Up to a choice of orientation of $\gamma_{q}$ we may assume that the tangent vector field along $\gamma_{q}$ is $X:=-y_{1}\partial_{x_{1}}+x_{1}\partial_{y_{1}}$. Let $v:=\partial_{x_{2}}|_{\gamma_{q}}$ be the vector field $\partial_{x_{2}}$ restricted to $\gamma_{q}$. Then $L:=X\wedge v$ is a Lagrangian plane field along $\gamma_{q}$ with Maslov index $\mu(L,D_{-})=2$. On the other hand, if we choose $D_{+}$ as the disc bounded by $\gamma_{q}$, then the orientation of $\gamma_{q}$ as the boundary of $D_{+}$ is given by $-X$, and the corresponding Maslov index becomes $\mu(L=(-X)\wedge v,D_{+})=-2$. Either way the Maslov condition is not satisfied. So $M$ cannot be a ruled manifold, nor can it be obtained from any ruled manifold by a sequence of equivalent blowups at fixed points that are not minima for the moment map of the ruled manifold. We conclude that a 4-dimensional compact symplectic manifold $(M,\omega)$ does not admit an effective special Hamiltonian $S^{1}$-action. ∎
Proposition 3.4 (Fixed point set is discrete).
Let $(M,\omega)$ be a connected noncompact symplectic 4-manifold. Suppose that there is a special Hamiltonian $S^{1}$-action $G:S^{1}\to Ham(M,\omega)$ on $M$ with contractible orbits and nonempty fixed point set $Fix(G)$, then $Fix(G)$ consists of isolated points.
Proof.
Let $G:S^{1}\to Ham(M,\omega)$ be a special Hamiltonian $S^{1}$-action on $(M,\omega)$. Denote by $X$ the Hamiltonian vector field on $M$ which generates the $G$-action, and $h:M\to\mathbb{R}$ the corresponding moment map defined by $\omega(X,\cdot)=-dh$. By assumption all $G$-orbits are contractable and $Fix(G)$ is not empty. It is known that every connected component of $Fix(G)$ is either a single point or a symplectic surface ([8], Appendix A). If $Fix(G)$ contains a connected symplectic surface $\Sigma$ then $h$ is constant on $\Sigma$ and $h(\Sigma)$ is a local extremum of $h$.
Fix a point $p\in\Sigma$ and an open neighborhood $U_{p}\subset\Sigma$ of $p$ so that $G$ acts nontrivially near $U_{p}$ except on $U_{p}$, and with isomorphic stabilizer group $\mathbb{Z}_{n}$ for some $n\in\mathbb{N}$. There is a $G$-invariant tubular neighborhood of $U_{p}$ which is symplectomorphic to a trivial disc bundle $E\cong D\times U_{p}$ over $U_{p}$ on which $G$ acts nontrivially as rotations along fibers of $E$, fixing $U_{p}$ pointwise. Since $\Sigma$ is symplectic we may identify $E=D\times U_{p}$ with $D_{1}\times D_{2}\subset\mathbb{C}\times\mathbb{C}$ with $U_{p}=\{0\}\times D_{2}$ being the base disc centered at $p=(0,0)$, and $D_{1}\times\{pt\}$ as fibers, with $G$ acts $D_{1}\times D_{2}$ by $t\cdot(z_{1},z_{2})=(e^{\pm int}z_{1},z_{2})$, $t\in G\cong\mathbb{R}/2\pi\mathbb{Z}$.
Let $\gamma$ be a $G$-orbit in the disc fiber over $p$ and $D_{\gamma}\subset D_{1}\times\{0\}$ the fiber disc over $p$ bounding $\gamma$. Assume that $h(\Sigma)$ is a local minimum of $h$ then $h=\frac{n}{2}|z_{1}|^{2}+constant$ on $E$ and the tangent vector field of $\gamma$ is $X=n(-y_{1}\partial_{x_{1}}+x_{1}\partial_{y_{1}})$. Let $v:=\partial_{x_{2}}|_{\gamma}$. Then $L:=X\wedge v$ is a loop of Lagrangian planes along $\gamma$ with maslov index $\mu(L,D_{\gamma})=2n\neq 0$. On the other hand, if $h(\Sigma)$ is a local maximum of $h$ then $h=-\frac{n}{2}|z_{1}|^{2}+constant$ on $E$ and the tangent vector field of $\gamma$ is $X=-n(-y_{1}\partial_{x_{1}}+x_{1}\partial_{y_{1}})$. In this case $\mu(L,D_{\gamma})=-2n\neq 0$. Either way the Maslov condition is not met. So $Fix(G)$ does not contain any symplectic surface. We conclude that $Fix(G)$ is a discrete set of points. ∎
Proposition 3.5.
Let $(M,\omega=d\lambda,G)$ be a connected, exact special Hamiltonian $S^{1}$-space and $\lambda$ a $G$-invariant primitive 1-form of $\omega$. Then
  1. (i).
    $Fix(G)\subset h^{-1}(0)$ up to addition of a constant to $h$, $h^{-1}(c)$ is connected for $c\in\mathbb{R}$, and
  2. (ii).
    the $G$-action is semi-free.
Proof.
(i). Let $X$ be the Hamiltonian vector field of the moment map $h$ associated to the $G$-action. Observe that
$-dh=d\lambda(X_{h},\cdot)=L_{X_{h}}\lambda-d(\lambda(X_{h}))=-d(\lambda(X_{h}))$ (3)
with $L_{X_{h}}\lambda=0$ since $\lambda$ is $G$-invariant, so $\lambda(X_{h})-h$ is a constant. Up to an addition of a constant to $h$ we may assume that
$h=\lambda(X_{h})\quad\text{and hence \quad$\lambda(X_{h})=0$ on $h^{-1}(0)$},$ (4)
then
$Fix(G)\subset h^{-1}(0)$
since $\lambda(X_{h})$ is constant on level sets of $h$.
We claim that $h^{-1}(0)$ is connected. Let $\nabla h$ denote the gradient vector field of $h$ with respect to some $G$-invariant $\omega$-compatible Riemannian metic on $M$. Then the vector fields $-\nabla h|_{h>0}$ and $\nabla h|_{h<0}$ together induces a deformation retraction from $M$ onto $h^{-1}(0)$, preserving level sets of $h$, and with $h^{-1}(0)$ fixed all the time, so $M$ is deformation equivalent and hence homotopic to $h^{-1}(0)$. Since $M$ is connected by hypothesis, so is $h^{-1}(0)$. For $c\neq 0$, $h^{-1}(c)$ is a circle bundle over the reduced space $M_{red}\cong h^{-1}(c)/G\cong h^{-1}(0)/G$. Since $h^{-1}(0)/G$ is connected, so is $h^{-1}(c)$ for $c\in\mathbb{R}$.
(ii). Suppose that $(M,\omega=d\lambda,G)$ contains exceptional $G$-orbits. Since $\omega$ is exact there are no symplectic exceptional spheres in $M$, as all critical points are on the same level set of the associated moment map $h$ say, $h^{-1}(0)$. Now assume that the $G$-action on $M$ has $n\in\mathbb{N}$ fixed points $p_{1},p_{2},\dots,p_{n}$. Maslov condition implies that in a small Darboux chart $U_{k}$ of $p_{k}$ $G$ acts as the the $S^{1}$ subgroup
$\begin{pmatrix}\cos m_{k}\theta&-\sin m_{k}\theta\\ \sin m_{k}\theta&\cos m_{k}\theta\end{pmatrix},\quad m_{k}\in\mathbb{Z}\setminus\{0\},\quad|m_{k}|>1\ \text{for some}\ 0\leq k\leq n.$
As $h^{-1}(0)$ is $G$-invariant and connected all elements in $h^{-1}(0)\setminus Fix(G)$ and hence in $M\setminus Fix(G)$ have isomorphic stabilizers, hence all $m_{k}$ are equal, the $G$-action is not effective unless either $m_{k}=1$ for all $k$, or $m_{k}=-1$ for all $k$. So there are no exceptional orbits of an effective special Hamiltonian action on $(M,\omega)$ if $\omega$ is exact. We conclude that the $G$-action on $(M,\omega=d\lambda)$ is semi-free. ∎
4 Simply-connected Exact Special Hamiltonian $S^{1}$-spaces
The rest of this paper will focus on the classification of connected exact special Hamiltonian $S^{1}$-spaces. We assume that $Fix(G)$ is a nonempty finite set as before.
Let
$\EuFrak{r}:M\to M_{red}$
denote the symplectic reduction of the $G$-action, we also denote
$\EuFrak{r}_{0}=\EuFrak{r}|_{h^{-1}(0)}:h^{-1}(0)\to h^{-1}(0)/G=M_{red}$
the reduction of $\EuFrak{r}$ restricted to the level set $h^{-1}(0)$.
Proposition 4.1.
Let $(M,\omega=d\lambda,G)$ be a connected, 1-connected exact Hamiltonian $S^{1}$-space with $c_{1}(M,\omega)=0$, $\lambda$ being $G$-invariant, and $G$ acts semi-freely on $M$ with moment map $h:M\to\mathbb{R}$, and $Fix(G)\subset h^{-1}(0)$ is a finite set.
  1. (i).
    $h^{-1}(c)/G$ is diffeomorphic to a disc.
  2. (ii).
    $\lambda(X_{h})=h$ on $M$. Moreover $d\lambda(X_{h},\cdot)=0$ on $h^{-1}(0)$. In particular, for $q\in h^{-1}(0)\setminus Fix(G)$, the symplectic normal space $T^{\omega}_{q}(h^{-1}(0)):=\{v\in T_{q}M\mid\omega_{q}(v,u)=0,\ \forall u\in T_{q}(h^{-1}(0))\}$ is generated by $(X_{h})_{q}$.
Proof.
(i). Recall that $h^{-1}(c)$ is connected for all $c$. Observe that the gradient flows of ±$\pm\nabla h$ induce a deformation retract from $M$ to $h^{-1}(0)$, keeping $h^{-1}(0)$ fixed all the time. So $M$ is homotopic to $h^{-1}(0)$. Since $M$ is 1-connected, so is $h^{-1}(0)$. $\EuFrak{r}_{0}:h^{-1}(0)\to h^{-1}(0)$ is a singlar fibration over $h^{-1}(0)/G$, So $h^{-1}(0)/G$ is 1-connected. Moreover, the gradient flows of ±$\pm\nabla f$ induce a homeomorphism between the quotient spaces $h^{-1}(c)/G$ and $h^{-1}(0)/G$ for all $c\in\mathbb{R}$. So $h^{-1}(c)/G$ is 1-connected for all $c$ provided that $M$ is 1-connected.
(ii). $\lambda$ is a $G$-invariant primitive 1-form of $\omega$, hence $0=L_{X_{h}}\lambda=d(\lambda(X_{h}))+d\lambda(X_{h},\cdot)=d(\lambda(X_{h}))-dh$. Since $d(\lambda(X_{h}))-dh=0$ on $h^{-1}(0)$, $\lambda(X_{h})$ is constant on $h^{-1}(0)$. $h^{-1}(0)$ is connected and $\lambda(X_{h})=0$ at $p\in Fix(G)$ so $\lambda(X_{h})=0$ and hence $d\lambda(X_{h},\cdot)=0$ on $h^{-1}(0)$. Therefore $(X_{h})_{q}\in T^{\omega}_{q}(h^{-1}(0))$ all $q\in h^{-1}(0)\setminus Fix(G)$. Since $\dim(T^{\omega}_{q}(h^{-1}(0))+\dim T_{q}(h^{-1}(0))=\dim T_{q}M=4$, $X_{h}\subset T(h^{-1}(0))$ generates the field of symplectic normal spaces of $T(h^{-1}(0)\setminus Fix(G))$. ∎
Remark 4.2.
Note that Property (ii) above implies that for any embedded smooth curve $\gamma\subset M_{red}$, its preimage $L_{\gamma}:=\EuFrak{r}_{0}^{-1}(\gamma)\subset h^{-1}(0)$ is an immersed monotone Lagrangian surface of $(M,\omega)$. For example, if $\gamma$ is a circle missing the set $\EuFrak{r}_{0}(Fix(G))$ then $L_{\gamma}$ is an embedded monotone Lagrangian torus. If $\gamma$ is a circle which meets the set $\EuFrak{r}_{0}(Fix(G))$ at exactly one point, then $L_{\gamma}$ is a Lagrangian Whitney sphere, i.e., an immersed monotone Lagrangian torus with one transversal self-intersection point. If $\gamma$ is an embedded arc which connects exactly two distinct points in $\EuFrak{r}_{0}(Fix(G))$ and at its endpoint, then $L_{\gamma}$ is an embedded Lagrangian sphere.
It turns out that a 1-connected exact special Hamiltonian 4-space $(M,\omega,G)$ with finite $Fix(G)$ is symplectically the plumbing of a finite number of $T^{*}S^{2}$.
Proof of Theorem 1.2.
Let $\gamma:(-\infty,\infty)\to h^{-1}(0)/G=M_{red}\cong\mathbb{C}$ be a smoothly embedded curve so that the complement $(h^{-1}(0)/G)\setminus\gamma$ consists of two connected components, and $\gamma(t_{i})=p_{i}$ for some $0=t_{0}<t_{1}<t_{2}<\cdots<t_{n}<\infty$, where $\{p_{0},p_{1},\dots,p_{n}\}=Fix(G)$. Let $\gamma_{i}:=\gamma([t_{i-1},t_{i}])$, $\gamma_{0}=\gamma((-\infty,0])$, $\gamma_{n+1}=\gamma([t_{n},\infty))$. Let $S_{i}:=\EuFrak{r}_{0}^{-1}(\gamma_{i})\subset h^{-1}(0)$, $S_{i}$ is a $G$-invariant Lagrangian sphere which is smooth except perhaps at the two points $p_{i-1}$ and $p_{i}$.
Let $U_{i}\subset M$ be a $G$-invariant symplectic Darboux chart centered at $p_{i}$, $U_{i}\cap U_{j}=\emptyset$ for $i\neq j$. Let $U:=\cup_{i=0}^{n}U_{i}$. By shrinking $U_{i}$ if necessary we may assume that $=\textsf{G}$ G on $U$. Then $=\textsf{h}$ h and $X_{h}=X_{\textsf{h}}$ on $U$. Each $U_{i}$ is symplectically identified with a open neighborhood of $0\in\mathbb{C}^{2}$ with $p_{i}$ identified with the origin $0\in\mathbb{C}^{2}$. Recall the symplectic reduction $\EuFrak{r}:M\to M_{red}\cong\mathbb{C}$ of the $G$-action and its restriction $\EuFrak{r}_{0}=\EuFrak{r}|_{h^{-1}(0)}$ to $h^{-1}(0)$. We may assume that $\EuFrak{r}(z_{1},z_{2})=\frac{1}{2}(z_{1}^{2}+z_{2}^{2})+\gamma(t_{i})$ when restricted to each $U_{i}$, and, by perturbing $\gamma$ if necessary, that $\EuFrak{r}_{0}^{-1}(\gamma_{i})\cap(U_{i-1}\cap h^{-1}(0))=\{y_{1}=0=y_{2}\}\cap U_{i-1}$ and $\EuFrak{r}_{0}^{-1}(\gamma_{i})\cap(U_{i}\cap h^{-1}(0))=\{x_{1}=0=x_{2}\}\cap U_{i}$ are Lagrangian discs for each $i$. Then
$S_{i}:=\EuFrak{r}_{0}^{-1}(\gamma_{i}),\quad i=1,...,n,$
are smoothly embedded Lagrangian spheres,
$S_{i}\pitchfork S_{i+1}=\{p_{i}\}\ \text{ for }i=1,...,n-1,\quad S_{i}\cap S_{j}=\emptyset\ \text{ if }|i-j|\neq 1.$
$S_{i}$ can be oriented so that the intersection number is $S_{i}\cdot S_{i+1}=1$ for $i=1,...,n-1$. The union $\mathcal{S}:=\cup_{i=1}^{n}S_{i}$ is an $A_{n}$-string of Lagrangian spheres.
By Lagrangian Neighborhood Theorem [13, 15], a tubular neighborhood $V_{i}\subset M$ of the Lagrangian sphere $S_{i}$ is symplectomorphic to a tubular neighborhood of the 0-section of the cotangent bundle $(T^{*}S_{i},\omega_{can})$ of $S_{i}$ equipped with the canonical symplectic structure. $T^{*}S_{i}$ and $T^{*}S_{i+1}$ are plumbed together near $p_{i}$ so that tangent and cotangent spaces of $S_{i}$ are identified with the cotangent and tangent spaces of $S_{i+1}$ respectively for $i=1,...,n-1$:
$T_{p_{i}}S_{i}=\partial_{x_{1}}\wedge\partial_{x_{2}}=T^{*}_{p_{i}}S_{i+1},\quad T^{*}_{p_{i}}S_{i}=\partial_{y_{1}}\wedge\partial_{y_{2}}=T_{p_{i}}S_{i+1},$
with $T^{*}S_{i}$ and $T^{*}S_{i+1}$ plumbed near $p_{i}$ via the identification $T_{p_{i}}S_{i}\cong T^{*}_{p_{i}}S_{i+1}$ with the following correspondence on local coordinates at $p_{i}$:
$T^{*}_{p_{i}}S_{i}\ni(x,y)\to(-y,x)\in T^{*}_{p_{i}}(S_{i+1}),$
[…]
Indeed we can take an open cover of $\EuFrak{r}(M)=h^{-1}(0)/G\cong\mathbb{R}^{2}$ by 1-connected domains $V_{i}$, $i=1,\cdots,n$ so that $\EuFrak{r}(M)=\cup_{i}V_{i}$, $S_{i}\subset V_{i}$, $V_{i}\cap V_{j}=\emptyset$ if $|i-j|>1$, and $V_{i}\cap V_{i+1}$ is 1-connected. Then each $\EuFrak{r}^{-1}(V_{i})$ is symplectomorphic to an open neighborhood in $T^{*}S_{i}$ of the zero section $S_{i}$ of $T^{*}S_{i}$, and $V_{i}\cap V_{i+1}$ is where the plumbing of $T^{*}S_{i}$ and $T^{*}S_{i+1}$ takes place. This completes the proof of Theorem 1.2.
5 A special Hamiltonian $S^{1}$-action on $\mathcal{W}_{n}$
We start with a verification of Proposition 1.3:
For each $n\in\mathbb{N}\cup\{0\}$ consider the Stein surface in $\mathbb{C}^{3}$ given by
$\mathcal{W}_{n}:z_{1}^{2}+z_{2}^{2}+z_{3}^{n+1}=1,\quad(z_{1},z_{2},z_{3})\in\mathbb{C}^{3}.$
The G-action on $\mathbb{C}^{2}$ can be lifted to a special Hamiltonian $S^{1}$-action on $\mathbb{C}^{3}$ which acts trivially on the $z_{3}$-coordinate:
$\tilde{\textsf{G}}:=\Big{\{}\tilde{\textsf{G}}(\theta)=\begin{pmatrix}\cos\theta&-\sin\theta&0\\ \sin\theta&\cos\theta&0\\ 0&0&1\end{pmatrix}\mid\theta\in\mathbb{R}/2\pi\mathbb{Z}\Big{\}}\subset SU(3).$
The action of $\tilde{\textsf{G}}$ preserves $\mathcal{W}_{n}$ and restricts to a special Hamiltonian $S^{1}$-action on $\mathcal{W}_{n}$, denoted as G for simplicity, with moment map
$\textsf{h}=\Im(z_{1}\bar{z}_{2})=x_{2}y_{1}-x_{1}y_{2}\in C^{\infty}(\mathbb{C}^{3}),$
the corresponding Hamiltonian vector field is
$X=-x_{2}\partial_{x_{1}}+x_{1}\partial_{x_{2}}-y_{2}\partial_{y_{1}}+y_{1}\partial_{y_{2}},$
and the fixed point set is
$Fix(\textsf{G})=\{p_{k}:=(0,0,\xi^{k})|k=0,...,n\}\subset\textsf{h}^{-1}(0),\quad\xi:=e^{\frac{2\pi i}{n+1}}.$
The tangent space of $\mathcal{W}_{n}$ at $p_{k}$ is $T_{p_{k}}\mathcal{W}_{n}=\mathbb{C}^{2}_{z_{1}z_{2}}$ for each $k$. Since h is independent of $z_{3}$, the Morse index of h at $p_{k}$ is $2$ as in Example 2.2. It follows that the G-action on $\mathcal{W}_{n}$ is special Hamiltonian. This completes the proof of Proposition 1.3.
Consider the standard projection
$\pi^{\prime}:\mathcal{W}_{n}\subset\mathbb{C}^{3}\to\mathbb{C},\quad\pi^{\prime}(z_{1},z_{2},z_{3})=z_{3}\quad\text{ restricted to }\mathcal{W}_{n}.$
[…]
Consider the level set $\textsf{h}^{-1}(0)$. Without loss of generality, we may restrict to a G-invariant Stein domain
$W_{n}\subset\mathcal{W}_{n},\quad W_{n}=(\pi^{\prime})^{-1}(D)\cap h^{-1}([-c,c])$
for some $c>0$ and $D=\{|z|\leq r\}$ for some $r>1$.
On $W_{n}$ consider the Riemann surface
$\Sigma=\{z_{2}=0\}\cap W_{n}\subset\textsf{h}^{-1}(0)$
which contains $Fix(\textsf{G})$ and is transversal to G-orbits except at $Fix(\textsf{G})$. Let $\Sigma_{\theta}:=\textsf{G}(\theta)(\Sigma)$, then $\Sigma_{\theta+\pi}=\Sigma_{\theta}$ for each $\theta$, and $\Sigma_{\theta}\cap\Sigma_{\theta^{\prime}}=Fix(\textsf{G})$ if $\theta\neq\theta^{\prime}\mod\pi$. Moreover,
$\textsf{h}^{-1}(0)=\cup_{\theta\in\mathbb{R}/\pi\mathbb{Z}}\Sigma_{\theta}.$
For each $\theta$ the symplectic reduction map $\EuFrak{r}_{0}|_{\Sigma_{\theta}}:\Sigma_{\theta}\to\mathbb{C}$ is a $2:1$ branched covering map branching at $Fix(\textsf{G})$, with branching locus $\{\xi\in\mathbb{C}\mid\xi^{n+1}=1\}\subset\mathbb{C}$. Let $g$ and $b$ be the genus and boundary components of $\Sigma$ (hence each $\Sigma_{\theta}$), then $2-2g-b=2-(n+1)$. Observe that the subgroup $K:=\{\pm Id\}\subset\textsf{G}$ acts on $\partial\Sigma$, $\partial\Sigma/K$ is a simple closed circle which is diffeomorphic to $\partial D$ under the map $\EuFrak{r}_{0}$. Therefore $b=1$ or $2$, $\Sigma$ is a surface of genus $g=\lfloor\frac{n}{2}\rfloor$ with $b=(n+1)-2g$ boundary components. More explicitly,
$(g,b)=\begin{cases}(m,1)&\text{ if $n=2m\geq 0$ is even},\\ (m,2)&\text{ if $n=2m+1\geq 1$ is odd}.\end{cases}$
To see the $A_{n}$-string of Lagrangian spheres in $W_{n}\subset\mathcal{W}_{n}$, let
$\gamma_{k}=\Big{\{}e^{i\theta}\in\mathbb{C}_{z_{3}}\Big{|}\frac{2(k-1)\pi}{n+1}\leq\theta\leq\frac{2k\pi}{n+1}\Big{\}},\quad k=1,\dots,n,$ (5)
$\gamma_{k}$ is an arc between $\xi^{k-1}$ and $\xi^{k}$. Orient $\gamma_{k}$ so that its tangents point from $\xi^{k-1}$ to $\xi^{k}$. Then $\textsf{S}_{k}:=(\EuFrak{r}_{0})^{-1}(\gamma_{k})$ is a G-invariant Lagrangian sphere, $\hat{\gamma}_{k}:=\textsf{S}_{k}\cap\Sigma_{0}$ a simple closed curve, and $\textsf{S}_{k}\cap\Sigma_{\theta}=\textsf{G}(\theta)(\hat{\gamma}_{k})$, $\textsf{S}_{k}=\cup_{\theta\in\mathbb{R}/\pi\mathbb{Z}}\textsf{G}(\theta)(\hat{\gamma}_{k})$. The union $\cup_{k=1}^{n}\textsf{S}_{k}$ is an $A_{n}$-string of Lagrangian spheres. $\mathcal{W}_{n}$ is an $A_{n}$-surface singularity as in [16].
Proof of Theorem 1.4.
Let $(M,\omega,G,h)$ be a connected and 1-connected exact special Hamiltonian $S^{1}$ space with $n+1$ fixed points, $n\in\mathbb{N}\cup\{0\}$, and $X$ the Hamiltonian vector field associated to $h$. Let $S=S_{1}\cup S_{2}\cdots\cup S_{n}$ be an $A_{n}$-string of Lagrangian spheres in $h^{-1}(0)\subset(M,\omega,G)$ associated with $(G,h)$, and $\textsf{S}:=\textsf{S}_{1}\cup\textsf{S}_{2}\cdots\cup\textsf{S}_{n}$ the $A_{n}$-string of Lagrangian spheres in $(\mathcal{W}_{n},\omega_{n},\textsf{G})$, say, induced from (5). There exists a diffeomorphism $\phi:S\to\textsf{S}$ so that $\phi(S_{i})=\textsf{S}_{i}$ for $i=1,\cdots,n$, and $\textsf{G}\circ\phi=\phi\circ G$ on S. Identify $\mathcal{W}_{n}=\cup_{i=1}^{n}T^{*}\textsf{S}_{i}=\cup_{i=1}^{n}T^{*}S_{i}$ with the plumbing of $n$ copies of the cotangent bundle $T^{*}S^{2}$ of Lagrangian $2$-spheres. By Darboux theorem we may extend $\phi$ to a symplectomorphism from a neighborhood of $S_{i}\cap S_{i+1}$ to a neighborhood of $\textsf{S}_{i}\cap\textsf{S}_{i+1}$ for each $i$. Since a diffeomorphism between two Lagrangian surfaces always extends to a symplectomorphism between their cotangent bundles, we can extend $\phi$ to a symplectomorphism $\phi:M\to\mathcal{W}_{n}$, sending $T^{*}S_{i}$ to $T^{*}\textsf{S}_{i}$ so that $\phi^{*}\omega_{n}=\omega$.
Proof of Theorem 1.6(i).
Recall that the $n=0$ case follows from Proposition 3.1 (ii), so we may assume that $n\geq 1$. With $\phi$ in the proof of Theorem 1.4 we may identify $(M,\omega,G,h)$ with $(\mathcal{W}_{n},\omega_{n},G,h)$, so that $G$ acts on $(\mathcal{W}_{n},\omega_{n}=d\alpha_{n})$ as an $S^{1}$-group of Hamiltonian diffeomorphisms $G_{\theta}\in Ham(\mathcal{W}_{n},\omega_{n})$, $G_{0}=id$. Both $G$ and G actions preserve $\textsf{S}=\textsf{S}_{1}\cup\cdots\cup\textsf{S}_{n}$ with $Fix(G)=Fix(\textsf{G})=\{p_{0},p_{1},\cdots,p_{n}\}$, $p_{0}\in\textsf{S}_{1}$, $p_{i}=\textsf{S}_{i}\cap\textsf{S}_{i+1}$ for $i=1,\cdots,n-1$, and $p_{n}\in\textsf{S}_{n}$. Since both $G$ and G acts on each $\textsf{S}_{i}$ as rotations with $p_{i-1}$ and $p_{i}$ fixed, by modifying $\phi$ if necessary we may assume that $G_{\theta}=\textsf{G}_{\theta}$ on S for all $\theta$. The Euclidean metric $g_{0}$ on $\mathbb{C}^{3}$ restricts to a G-invariant, $\omega_{n}$-compatible Riemannian metric on $\mathcal{W}_{n}$. Also let $g$ denote a $\omega_{n}$-compatible metric on $\mathcal{W}_{n}$ which is invariant under the $G$-action. We may assume that $g=g_{0}$ on $T_{\textsf{S}_{i}}\mathcal{W}_{n}$ the tangent bundle of $\mathcal{W}_{n}$ over $\textsf{S}_{i}$ for all $i$. Let $N(\textsf{S}_{i})$ denote the normal bundle of $\textsf{S}_{i}\subset\mathcal{W}_{n}$ determined by the metric $g_{0}$, which can also be identified with the cotangent bundle $T^{*}\textsf{S}_{i}$. Note that
$N_{p}(\textsf{S}_{i})=T_{p}\textsf{S}_{j},\qquad T_{p}\textsf{S}_{i}=N_{p}(\textsf{S}_{j})\ \text{ at $p=\textsf{S}_{i}\cap\textsf{S}_{j}$ when $|i-j|=1$}.$ (6)
Recall that the G-action on $\mathcal{W}_{n}$ is linear. Since $G_{\theta}=\textsf{G}_{\theta}$ on $\textsf{S}_{i}$, the linearized $G$-action $G^{L}:=G_{*}$ on $N(S_{i})$ acts on $N(\textsf{S}_{i})$ in the same way as $\textsf{G}=\textsf{G}_{*}$. Denote $x_{\theta}:=G_{\theta}(x)=\textsf{G}_{\theta}(x)$ and $v_{\theta}:=(G_{\theta})_{*}(v)=\textsf{G}_{\theta}(v)\in N_{x_{\theta}}(\textsf{S}_{i})$ for $x\in(\textsf{S}_{i})$, $v\in N_{x}(\textsf{S}_{i})$. We have for $v\in N_{x}(\textsf{S}_{i})$
$(G_{\theta})_{*}=\textsf{G}_{\theta}:N_{x}(\textsf{S}_{i})\to N_{x_{\theta}}(\textsf{S}_{i}),\quad v\to v_{\theta}=\textsf{G}_{\theta}(v).$
At $p_{i}\in Fix(G)$ we have
$(G_{\theta})_{*}=\textsf{G}_{\theta}:T_{p_{i}}\mathcal{W}_{n}\to T_{p_{i}}\mathcal{W}_{n}$
and the action preserves the splitting $T_{p_{i}}\mathcal{W}_{n}=T_{p_{i}}\textsf{S}_{i}\oplus N_{p_{i}}(\textsf{S}_{i})$. Below we consider exponential maps associated to $\textsf{S}\subset\mathcal{W}_{n}$.
For each $i$ consider the exponential map over $\textsf{S}_{i}$ relative to $g$, so that for $x\in\textsf{S}_{i}$ and $v\in N_{x}(\textsf{S}_{i})$,
$\EuFrak{e}_{i}:=\exp_{i}:\ N(\textsf{S}_{i})\to\mathcal{W}_{n},\quad v\to\gamma_{v}(1)$
where $\gamma_{v}(1)$ is the time $t=1$ image of the geodesic $\gamma_{v}(t)$ relative to $g$ with $\gamma_{v}(0)=x$ and $\dot{\gamma}_{v}(0)=v$.
If $n>1$ then at each $p_{i}\in Fix(G)$ with $1\leq i<n$ we also include the exponential map
$\EuFrak{e}_{p_{i}}:=\exp_{p_{i}}:T_{p_{i}}\mathcal{W}_{n}\to\mathcal{W}_{n}$
relative to any $G$-invariant $\omega_{n}$-compatible Riemannian metric $g$: for $v\in T_{p_{i}}\mathcal{W}_{n}$
$\EuFrak{e}_{p_{i}}(v):=\gamma_{v}(1)$
is the time $t=1$ image of the geodesic $\gamma_{v}(t)$ relative to the metric $g$ with $\gamma_{v}(0)=p_{i}$ and $\dot{\gamma}_{v}(0)=v$.
With (6) the $\EuFrak{e}_{p_{i}}$’s and $\EuFrak{e}_{i}$’s piece together well to form a differentiable map $\EuFrak{e}_{\textsf{S}}$ from the cotangent/normal bundle
$N(\textsf{S}):=\begin{cases}N(\textsf{S}_{1})&\text{if $n=1$}\\ \cup_{i=1}^{n}N(\textsf{S}_{i})\cup_{i=1}^{n-1}T_{p_{i}}\mathcal{W}_{n}&\text{if $n>1$}\end{cases}$
into $\mathcal{W}_{n}$. There exists $\epsilon>0$ so that the map
$\EuFrak{e}_{\textsf{S}}:=\exp_{\textsf{S}}:\ V:=\{v\in N(\textsf{S})\mid|v|<\epsilon\}\to\mathcal{W}_{n}$
is a diffeomorphism from $V$ onto the open set $U:=\EuFrak{e}_{\textsf{S}}(V)$. We can also consider the exponential map $\EuFrak{e}^{0}_{\textsf{S}}:=\exp^{0}_{\textsf{S}}$ relative to the metric $g_{0}$, and for $v\in N(\textsf{S})$ based at $x\in\textsf{S}$, we write $\gamma^{0}_{v}(t)$ as the geodesic relative to $g_{0}$ with $\gamma^{0}_{v}(0)=x$ and $\dot{\gamma}^{0}_{v}(0)=v$.
Since $^{L}:=G_{*}=\textsf{G}$ G and $g=g_{0}$ on $T_{\textsf{S}}\mathcal{W}_{n}$, $\gamma_{v}$ and $\gamma^{0}_{v}$ are tangent at the base point $x$ of $v$. For $v\in N(\textsf{S})$ denote $v_{\theta}:=G^{L}_{\theta}(v)$, $v_{0}=v$, then
$G^{L}_{\theta}(tv)=tv_{\theta},\quad G^{L}_{\theta_{1}+\theta_{2}}(v)=G^{L}_{\theta_{1}}G^{L}_{\theta_{2}}(v)=G^{L}_{\theta_{1}}(v_{\theta_{2}})=v_{\theta_{1}+\theta_{2}}.$
Since $g$ is $G$-invariant, $G_{\theta}$ maps $g$-geodesics to $g$-geodesics, preserving the arc length. Also, $G_{\theta}=\textsf{G}_{\theta}$ when restricted to S, we have $G_{\theta}(\gamma_{v})=\gamma_{v_{\theta}}$, and for $v\in N(\textsf{G})$ based at $x\in\textsf{S}$, the map $\EuFrak{e}_{\textsf{S}}$ takes $\gamma^{0}_{v}(1)$ to $\gamma_{v}(1)$, and $\gamma^{0}_{v_{\theta}}(1)$ to $\gamma_{v_{\theta}}(1)$.
The domain $U:=\EuFrak{e}_{\textsf{S}}(V)$ is $G$-invariant since $V$ is $^{L}=\textsf{G}$ G-invariant and $g=g^{0}$ on $V\subset N(\textsf{S})$. Also we may identify $(V\subset N(\textsf{S}),\omega_{n})$ with its image $(\EuFrak{e}_{\textsf{S}}(V)=U,\omega_{n})$. Then on $U$
$G=\EuFrak{e}_{\textsf{S}}\circ\textsf{G}\circ(\EuFrak{e}_{\textsf{S}})^{-1}.$
Let
$\omega^{\prime}:=((\EuFrak{e}_{\textsf{S}})^{-1})^{*}\omega_{n}=((\EuFrak{e}_{\textsf{S}})^{-1})^{*}\sum_{j=1}^{3}dx_{j}\wedge dy_{j}=\sum_{j=1}^{3}dx^{\prime}_{j}\wedge dy^{\prime}_{j},$
[…]
$(X:=(\EuFrak{e}_{\textsf{S}})_{*}\textsf{X},\ h:=\textsf{h}\circ\EuFrak{e}^{-1}_{\textsf{S}})$
is the the corresponding Hamiltonian vector field and Hamiltonian function associated to the $G$-action on $U$. We may extend $\EuFrak{e}_{\textsf{S}}$ to a diffeomorphism $\EuFrak{e}^{\prime}_{\textsf{S}}:\mathcal{W}_{n}\to\mathcal{W}_{n}$ with $G$-invariant compact support containing $U$. Let $G^{\prime}:=\EuFrak{e}^{\prime}_{\textsf{S}}\circ\textsf{G}\circ(\EuFrak{e}^{\prime}_{\textsf{S}})^{-1}=p$. $G^{\prime}$ acts as a Hamiltonian $S^{1}$-action $\mathcal{W}_{n}$, with $G^{\prime}=G$ on $U$ and outside the compact support of $\EuFrak{e}^{\prime}_{\textsf{S}}$.
Note that $(\EuFrak{e}_{\textsf{S}})_{*}:=d\EuFrak{e}_{\textsf{S}}=id$ when restricted to S, so $\omega^{\prime}=\omega_{n}$ on S. Both $\omega^{\prime}=d\alpha^{\prime}$ and $\omega_{n}=d\alpha_{n}$ are $G$-invariant exact symplectic forms. Up to averaging via the $G$-action we may assume that both primitive 1-forms $\alpha_{n}$ and $\alpha^{\prime}$ are $G$-invariant. Then similar to the proof of Proposition 3.1 (ii), by shrinking $U$ if necessarily, there exists a smooth isotopy $\phi_{t}:U\to M$, $t\in[0,1]$, with $\phi_{0}=id$, $\phi_{t}=id$ on S for all $t$, and $\phi_{1}^{*}\omega^{\prime}=\omega_{n}$. Therefore, any effective special Hamiltonian $S^{1}$-action on $(\mathcal{W}_{n},\omega_{n})$ is linear near S up to an equivariant smooth isotopy $\phi_{t}$ with $\phi^{-1}_{1}\circ G\circ\phi_{1}=\textsf{G}$ on $U$. ∎
In the following we will show that on $(\mathcal{W}_{n},\omega_{n})$ with $n=0,1$, any semi-free Hamiltonian $S^{1}$-action, which we denote it as $G$, up to conjugations is symplectically equivariant to the G-action on any prescribed relatively compact domain containing (i) the fixed point of the G-action if $n=0$, and (ii) a G-invariant Lagrangian sphere if $n=1$.
Proof of Theorem 1.6(ii)(iii).
Case 1: $n=0$. We denote $q=(x,y)$ for $q\in(\mathbb{C}^{2},\omega=\sum_{j=1}^{2}dx_{j}\wedge dy_{j})$, where $x=(x_{1},x_{2})$, $y=(y_{1},y_{2})$, also $|x|=\sqrt{x_{1}^{2}+x_{2}^{2}}$, $|y|=\sqrt{y_{1}^{2}+y_{2}^{2}}$.
Recall that $\mathcal{W}_{0}:\ z_{1}^{2}+z_{2}^{2}+z_{3}=1$. The embedding
$\iota:\mathbb{C}^{2}\to\mathbb{C}^{3},\quad\iota(z_{1},z_{2})=(z_{1},z_{2},1-z_{1}^{2}-z_{2}^{2})$
is a symplectomorphism between $(\mathbb{C}^{2},\omega=\sum_{j=1}^{2}dx_{j}\wedge dy_{j})$ and its image $(\mathcal{W}_{0},\omega_{0})$, $\mathcal{W}_{0}=\iota(\mathbb{C}^{2})$. Conversely the projection map $\pi:\mathcal{W}_{0}\to\mathbb{C}^{2}$, $\pi(z_{1},z_{2},z_{3})=(z_{1},z_{2})$ is the inverse symplectomorphism, $\pi\circ\iota=id_{\mathbb{C}^{2}}$, $\iota\circ\pi=id_{\mathcal{W}_{0}}$, $\pi^{*}\omega=\omega_{0}$.
Consider The Liouville vector field $Y:=\frac{1}{2}\sum_{j=1}^{2}x_{j}\partial_{x_{j}}+y_{j}\partial_{y_{j}}$ on $\mathbb{C}^{2}$. The differential $\iota_{*}:T\mathbb{C}^{2}\to T\mathcal{W}_{0}$ induces a Liouville vector field $Y_{0}$ on $\mathcal{W}_{0}$ given by $Y_{0}:=\iota_{*}Y$. Both $Y$ and $Y_{0}$ are G-invariant. Recall from Proposition 3.1 $G$ and G are symplectically equivariant when restricted to some open neighborhood $U$ of the fixed point. Since $\mathbb{C}_{2}$ and $\mathcal{W}_{0}$ are symplectomorphic G-spaces, and any semi-free special Hamiltonian $S^{1}$-action on $\mathcal{W}_{0}$ descends to one in $\mathbb{C}^{2}$, so it suffices to compare $G$ and G in $(\mathbb{C}^{2},\omega)$. Then by Proposition 3.1 (ii) up to conjugation $G$-action is symplectically equivariant to G-action on some open neighborhood $U_{0}$ of $0\in\mathbb{C}^{2}$, i.e., there exists a diffeomorphism $\EuFrak{e}_{0}:U_{0}\to U$ from a G-invariant open neighborhood of $0\in\mathbb{C}^{2}$ onto a $G$-invariant open neighborhood $U$ of $0\in\mathbb{C}^{2}$, $\EuFrak{e}_{0}(0)=0\in\mathbb{C}^{2}$, and
$G_{\theta}=\EuFrak{e}_{0}\circ\textsf{G}_{\theta}\circ\EuFrak{e}_{0}^{-1}\quad\text{on $U$},\quad\forall\theta,$
or equivalently,
$\textsf{G}_{\theta}=\EuFrak{e}_{0}^{-1}\circ G_{\theta}\circ\EuFrak{e}_{0}\quad\text{on $U_{0}$},\quad\forall\theta.$
Let $\rho:[0,\infty)\to[0,1]$ be a $C^{\infty}$-function so that $\rho^{\prime}(t)\leq 0$ for $t\in[0,\infty)$, and for some $0<r_{1}<r_{0}<\infty$, $\rho(t)=1$ if $t\leq r_{1}$, $\rho(t)=0$ if $t\geq r_{0}$. Let $\tilde{\rho}:\mathbb{C}^{2}\to[0,1]$ be the $C^{\infty}$-function given by $\tilde{\rho}(q)=\rho(|q|^{2})$. Let $\tilde{Y}:=\tilde{\rho}Y$ and $\psi_{t}$ be the smooth isotopy on $\mathbb{C}^{2}$ associated to $\tilde{Y}$, $\psi_{0}=id$, $\frac{d\psi_{t}}{dt}=\tilde{Y}\circ\psi_{t}$. For $q=(x,y)\in\mathbb{C}^{2}$,
$\psi_{t}(q)=(e^{\frac{t}{2}\tilde{\rho}(q)}x,e^{\frac{t}{2}\tilde{\rho}(q)}y).$
For $r>0$ and $r<r_{1}$ there exists $t_{r}<0$ such that whenever $|q|^{2}\leq r$,
$\phi_{t_{r}}(x,y)=(e^{t_{r}/2}x,e^{t_{r}/2}y)\in U_{0},$
Let $\Psi_{r}:=\EuFrak{e}_{0}\circ\psi_{t_{r}}$, then for $q=(x,y)$ with $|q|^{2}\leq r$,
$\begin{split}(\Psi_{r}^{-1}\circ G_{\theta}\circ\Psi_{r})(x,y)&=\psi_{t_{r}}^{-1}\circ\EuFrak{e}_{0}^{-1}\circ G_{\theta}\circ\EuFrak{e}_{0}\circ\psi_{t_{r}}(x,y)\\ &=\psi_{t_{r}}^{-1}\circ\EuFrak{e}_{0}^{-1}\circ G_{\theta}\circ\EuFrak{e}_{0}(e^{t_{r}/2}x,e^{t_{r}/2}y)\\ &=\psi_{t_{r}}^{-1}\circ\textsf{G}_{\theta}(e^{t_{r}/2}x,e^{t_{r}/2}y)\\ &=\psi_{t_{r}}^{-1}(e^{t_{r}/2}x_{\theta},e^{t_{r}/2}y_{\theta})\\ &=(x_{\theta},y_{\theta})=\textsf{G}_{\theta}(x,y)\end{split}$ (7)
where
$\begin{split}x_{\theta}&:=(x_{1}\cos\theta+x_{2}\sin\theta,-x_{1}\sin\theta+x_{2}\cos\theta),\\ y_{\theta}&:=(y_{1}\cos\theta+y_{2}\sin\theta,-y_{1}\sin\theta+y_{2}\cos\theta).\end{split}$ (8)
So
$\Psi_{r}^{-1}\circ G_{\theta}\circ\Psi_{r}=\textsf{G}\quad\text{on }\{x^{2}+y^{2}\leq r<r_{1}\}\subset\mathbb{C}^{2}.$
By allowing $r_{1}$ and $r_{0}$ to be arbitrarily large we conclude that any semi-free special Hamiltonian $S^{1}$-action on $(\mathbb{C}^{2},\omega)$ is eventually symplectically equivariant to the standard linear action of G. This completes the proof of Case 1.
Case 2: $n=1$.  Consider $(\mathcal{W}_{1},\omega_{1})$ where $\mathcal{W}_{1}:\{z_{1}^{2}+z_{2}^{2}+z_{3}^{2}=1\}\subset\mathbb{C}^{3}$ is given by the equations
$\sum_{j=1}^{3}x_{j}^{2}=1+\sum_{j=1}^{3}y_{j}^{2},\quad\sum_{j=1}^{3}x_{j}y_{j}=0.$
For convenience we use the notations $x:=(x_{1},x_{2},x_{3})$, $y:=(y_{1},y_{2},y_{3})$, $x^{2}:=\sum_{j=1}^{3}x_{j}^{2}$ and $y^{2}:=\sum_{j=1}^{3}y_{j}^{2}$.
$\mathcal{W}_{1}$ is indeed the cotangent bundle of the sphere $S:=\{\sum_{j=1}^{3}x_{j}^{2}=1\}\subset\mathbb{R}^{3}$ which is Lagrangian viewed as a submanifold of $(\mathcal{W}_{1},\omega_{1})$, G acts on $S$ with two fixed points $(0,0,\pm 1)$. Consider the vector field on $\mathcal{W}_{1}$:
$Y:=\sum_{j=1}^{3}y_{j}\partial_{y_{j}}.$
The 1-form
$\alpha:=\iota_{Y}\omega_{1}=-\sum_{j=1}^{3}y_{j}\partial_{x_{j}}$
satisfies $d\alpha=\omega_{1}$, is a G-invariant primitive 1-form of $\omega_{1}$, so $Y$ is a Liouville vector field on $\mathcal{W}_{1}$ which vanishes precisely on the Lagrangian sphere $S$. Moreover, $Y$ is invariant under the G-action, which can be checked by direct computation.
Remark 5.1.
The vector field $Y:=\sum_{j=1}^{3}y_{j}\partial_{y_{j}}$ on $\mathcal{W}_{1}$ is indeed the gradient vector field $\nabla f$ of the $C^{\infty}$-function $f(x,y)=\sum_{j=1}^{3}y_{j}\partial_{y_{j}}$ with respect to the Riemannian metric which is the pullback of the Euclidean metric on $\mathbb{C}^{3}$ associated to the inclusion map $\iota:\mathcal{W}_{1}\to\mathbb{C}^{3}$.
It is known that all Lagrangian spheres in the cotangent bundle $T^{*}S^{2}$ are hamiltonian isotopic [7], so are those in $\mathcal{W}_{1}$. Recall from Proposition 3.1 that up to conjugation any semi-simple special Hamiltonian $S^{1}$-action on $\mathcal{W}_{1}$ are symplectically equivalent to the standard G-action when restricted to some open neighborhood $U$ of the Lagrangian sphere $S$. We may assume that $(x,y)\in U$ whenever $|y|:=\sqrt{y^{2}}<\delta$ for some $\delta>0$.
Let $\eta_{t}$ be the time $t$ map of the flow of $Y$ with $\eta_{0}=id_{\mathcal{W}_{1}}$. $\eta_{t}$ is G-invariant for all $t$. For $(x,y)\in\mathcal{W}_{1}$
$\eta_{t}(x,y)=(e^{s}x,e^{t}y),\quad e^{2s}x^{2}=1+e^{2t}y^{2},\quad s,t\in\mathbb{R}.$
Consider a $C^{\infty}$-function $\rho:[0,\infty)\to[0,1]$ such that $\rho^{\prime}\leq 0$, and for some $0<r_{1}<r_{0}<\infty$, $\rho(r)=1$ if $0\leq r\leq r_{1}$, $\rho(r)=0$ if $r_{0}\leq r$. let $\tilde{\rho}:\mathcal{W}_{1}\to[0,1]$ be defined as $\tilde{\rho}(x,y):=\rho(|y|^{2})$. Let $Y_{\rho}:=\tilde{\rho}Y$.
Let $\psi_{t}$ be the time $t$ map of the flow of $Y_{\rho}$. $\frac{d\psi_{t}}{dt}=Y_{\rho}(\psi_{t})$. $\phi_{0}=id_{\mathcal{W}_{1}}$. $\frac{d\psi_{t}}{dt}=0$ when restricted to $S$. $\psi_{t}$ is G-equivariant for all $t$. For $(x,y)\in\mathcal{W}_{1}$,
$\psi_{t}(x,y)=(e^{s}x,e^{\rho(|y|^{2})t}y),\quad e^{2s}x^{2}=1+e^{2\rho(|y|^{2})t}y^{2}.$
For $r>0$ and $r<r_{1}$ there exists $t_{r}<0$ such that $\psi_{t_{r}}(x,y)\in U$ whenever $|y|\leq r$,
$\psi_{t_{r}}(x,y)=(e^{s_{r}}x,e^{t_{r}}y)\quad e^{2s_{r}}x^{2}=1+e^{2t_{r}}y^{2}.$
As in the proof of Theorem 1.6(i) we may assume that $=\textsf{G}$ G when restricted to the Lagrangian sphere $S:x^{2}=1,y=0$ in $\mathcal{W}_{1}$, and consider the exponential map $\EuFrak{e}_{S}$ over $S$ relative to a $G$-invariant $\omega_{1}$-compatible Riemannian metric $g$ on $\mathcal{W}_{1}$:
$\EuFrak{e}_{S}:=\exp_{S}:N(S)\to\mathcal{W}_{1},\quad\EuFrak{e}_{S}(v)=\gamma_{v}(1).$
There exists $\epsilon>0$ so
$\EuFrak{e}_{S}:V=\{v\in N(S)||v|<\epsilon\}\to\mathcal{W}_{1}$
is a diffeomorphism from $V$ onto the open set $U:=\EuFrak{e}_{S}(V)$. Then on $U$
$G_{\theta}=\EuFrak{e}_{S}\circ\textsf{G}_{\theta}\circ(\EuFrak{e}_{S})^{-1},\quad\forall\theta.$
Let $F_{r}:=\EuFrak{e}_{S}\circ\psi_{t_{r}}$. Similar to the $n=0$ case, we have
$F_{r}^{-1}\circ G_{\theta}\circ F_{r}=\textsf{G}_{\theta}\quad\text{on $U_{r}:=\{(x,y)\in\mathcal{W}_{1}||y|^{2}\leq r\}$}.$
By allowing $r_{1}$ and $r_{0}$ to be arbitrarily large we conclude that any semi-free special Hamiltonian $S^{1}$-action on $(\mathcal{W}_{1},\omega_{1})$ is eventually symplectically equivariant to the standard linear action of G. This completes the proof of Case 2.
6 Non-simply connected cases
Now we consider the case that $(M,\omega)$ is a connected exact symplectic 4-manifold with $c_{1}(M,\omega)=0$ but not 1-connected. Assume that $(M,\omega)$ equips with a semi-free special Hamiltonian $S^{1}$-action which we denote as $G\subset\text{Ham}(M,\omega)$.
Since the Maslov condition applies only when nonconstant orbits of $G$ are homologically trivial, we assume that $Fix(G)$ is nonempty and finite to accommodate the Maslov condition . Let
$k=n+1:=|Fix(G)|\geq 1$
denote the number of fixed points of the $G$-action, $h:M\to\mathbb{R}$ the moment map associated to $G$, and $X=X_{h}$ the Hamiltonian vector field of $h$ defined by $\omega(X,\cdot)=-dh$.
Fix any $\lambda\in\Omega^{1}(M)$ with $d\lambda=\omega$. As in the 1-connected case we may assume that $\lambda$ is $G$-invariant and $Fix(G)\subset h^{-1}(0)$, then $\lambda$ satisfies $\lambda(X)=h$. Let
$M_{red}:=h^{-1}(0)/G\cong h^{-1}(c)/G$
denote the symplectic reduction of $(M,\omega,G)$ via the reduction map
$\EuFrak{r}:M\to M_{red}$
associated to the $G$-action on $M$. $M_{red}$ is topologically an oriented surface of genus $g$ with $k$ marked points ($\EuFrak{r}(Fix(G))$) and $b\geq 1$ connected boundary components at infinity. Since $G$ acts freely on $h^{-1}(c)$ for all $c\neq 0$, $M_{red}$ is not 1-connected and hence
$g+b\geq 2,\quad g\geq 0,\ b\geq 1.$
Proof of Theorem 1.7
Let $B\subset M_{red}$ be a bounded open domain of $M_{red}$ which is homotopic to $M_{red}$ and contains $\EuFrak{r}(Fix(G))$ in its interior. Up to homotopy equivalence we may identify $M$ with the corresponding $G$-invariant subdomain $\EuFrak{r}^{-1}(B)\cap h^{-1}((-r,r))$ of $M$ for some $r>0$.
Part 1: Handle decomposition of $B$ and Homology of $M$.  For each of the $k=1+n$ fixed points $p_{i}$, $0\leq i\leq n$, of the $G$-action on $M$, denote by $q_{i}:=\EuFrak{r}(p_{i})\in B$ the image of $p_{i}$ under the symplectic reduction map $\EuFrak{r}:M\to M_{red}$. In $B$ pick $n+1$ open discs $D_{i}$ centered at $q_{i}$, $i=0,1,\dots,n$, such that the closures of $D_{i}$’s are pairwise disjoint closed discs in $B$. As $B$ can be constructed by attaching 2-dimensional 1-handles to 2-dimensional 0-handles, we may view $D_{i}$ as the set of 2-dimensional 0-handles of $B$, and with $q_{i}$ as the core point of $D_{i}$. Then $B$ can be obtained by attaching the following two types of 1-handles to $\cup_{i}D_{i}$:
  1. (i).
    A disjoint union of $n$ 1-handles which connect $\cup_{i=1}^{n}D_{i}$ to $D_{0}$:
    Let $\tau_{i}\subset B\setminus D_{0}$, $i=1,\dots,n$, be a set of pairwise disjoint embedded arcs with endpoints such that $\tau_{i}$ is disjoint from $q_{l}$ if $l\neq i$, $q_{i}=\partial_{+}\tau_{i}$ is one endpoint of $\tau_{i}$, and $\tau_{i}$ intersects transversally with $\partial D_{0}$ at the other endpoint $\partial_{-}\tau_{i}$ of $\tau_{i}$. Thicken each of $\tau_{i}$ a bit to get $n$ mutually disjoint 2-dimensional 1-handles $H_{\tau_{i}}$ with $\tau_{i}$ as the core curve of $H_{\tau_{i}}$, $H_{\tau_{i}}\cap D_{l}=\emptyset$ if $l\neq i$ or 00, and $H_{\tau_{i}}\cap D_{0}$ is the end interval of $H_{\tau_{i}}$ which contains the endpoint $\partial_{-}\tau_{i}$ of $\tau_{i}$. WLOG we may assume that that $D_{i}\cup H_{\tau_{i}}$ is 1-connected for $i=1,2,\dots,n$.
    Let
    $D:=D_{0}\cup(\cup_{i=1}^{n}\tilde{H}_{\tau_{i}}),\quad\tilde{H}_{\tau_{i}}:=H_{\tau_{i}}\cup D_{i},$
    $D$ is diffeomorphic to a 2-disc. Note that $Fix(G)\subset\EuFrak{r}^{-1}(D)$, and $\EuFrak{r}^{-1}(D)$ can be identified with a 1-connected Stein domain $W_{n}\subset\mathcal{W}_{n}$ is a Stein domain.
  2. (ii).
    A disjoint union $(2g+b-1)$ 1-handles with both ends attached to $\partial D_{0}$, missing all of the handles $H_{\tau_{i}}$ in (i):
    Denote these 1-handles as $H_{C_{j}}$ for $j=1,2,\dots,2g+b-1$, where $C_{j}$ is the core arc of $H_{C_{j}}$. Denote the boundary of $H_{C_{j}}$ as $\partial H_{C_{j}}=\partial_{+}H_{C_{j}}\cup\partial_{-}H_{C_{j}}$, $H_{C_{j}}$ is attached to $\partial D$ along $\partial_{-}H_{C_{j}}$, where $\partial_{-}H_{C_{j}}$ is the disjoint union of two short intervals containing $\partial C_{j}$. By handle sliding along $\partial D$ (and thinning $\tilde{H}_{\tau_{i}}$ and $H_{C_{j}}$ if necessary) we may assume that all $\partial_{-}H_{C_{j}}$ are attached to $\partial D_{0}$ and not touching any part of $\partial H_{\tau_{i}}$.
Then the union
$D\cup\Big{(}\sum_{j=1}^{2g+b-1}H_{C_{j}}\Big{)}=B$
up to a diffeomorphism.
Recall $\EuFrak{r}_{0}:h^{-1}(0)\to h^{-1}(0)/G=B$ the standard projection from $h^{-1}(0)\subset M$ to the reduced space. Fix a $G$-invariant $\omega$-compatible Riemannian metric on $M$ and let $\nabla h$ denote the gradient vector field of the moment map $h$ with respect to the metric. $M$ deformation retracts to the hypersurface $h^{-1}(0)$ via flows of gradients ±$\pm\nabla h$ of $h$, therefore $H_{i}(M,\mathbb{Z})\cong H_{i}(h^{-1}(0),\mathbb{Z})$ for all $i$.
Now $h^{-1}(0)=\EuFrak{r}_{0}^{-1}(B)$. As $G$ acts freely on $\EuFrak{r}_{0}^{-1}(B)\setminus Fix(G)$, we have the following results.
  1. (i).
    $\EuFrak{r}_{0}^{-1}(\tau_{i})\subset h^{-1}(0)$ is Lagrangian disc centered at $p_{i}$, $1\leq i\leq n$. If we extend $\tau_{i}$ to an embedded arc $\tilde{\tau}_{i}\subset D$ with $q_{i}$ and $q_{0}$ as endpoints, then $\tilde{S}_{i}:=\EuFrak{r}_{0}^{-1}(\tilde{\tau}_{i})$ is a Lagrangian 2-sphere with points $p_{i}$ and $p_{0}$ as poles for $i=1,2,\dots n$. These $n$ spheres are homologically independent in $\EuFrak{r}^{-1}_{0}(D)$ and hence in $W_{n}=\EuFrak{r}^{-1}_{G}(D)$. Indeed $W_{n}$ is homotopic to the bouquet $\cup_{i=1}^{n}\tilde{S}_{i}$ of $n$ 2-spheres.
  2. (ii).
    $\EuFrak{r}_{0}^{-1}(C_{j})\subset h^{-1}(0)$ is Lagrangian cylinder with boundary attached to $\EuFrak{r}^{-1}_{0}(\partial D)$. If we extend $C_{j}$ to a simple closed curve $\tilde{C}_{j}\subset D\setminus\{q_{i}\mid i=0,\dots n\}$ then we get an embedded Lagrangian torus $\tilde{T}_{j}\subset h^{-1}(0)=\EuFrak{r}^{-1}_{0}(B)$ for $j=1,2,\dots,2g+b-1$. These curves $\tilde{C}_{j}$, $j=1,\dots,2g+b-1$, together represent a basis of the homology group $H_{1}(B,\mathbb{Z})\cong\mathbb{Z}^{2g+b-1}$. As $G$-orbits in $h^{-1}(0)$ contracts to points in $Fix(G)$, $H_{1}(h^{-1}(0),\mathbb{Z})\cong H_{1}(B,\mathbb{Z})$ and is generated by a lifting of the curves $\tilde{C}_{j}$ in $\tilde{T_{j}}$, $j=1,\dots,2g+b-1$. The tori $\tilde{T}_{j}$’s are also independent on $H_{2}(h^{-1}(0),\mathbb{Z})$. Also $\tilde{T_{j}}$’s and $\tilde{S}_{i}$’s are pairwise linearly independent in $H_{2}(h^{-1}(0),\mathbb{Z})$, and forming a basis for $H_{2}(h^{-1}(0),\mathbb{Z})\cong\mathbb{Z}^{2g+b-1+n}$. Therefore
    $H_{m}(M,\mathbb{Z})=\begin{cases}0&\text{ for }m>2,\\ \mathbb{Z}^{2g+b-1+n}\cong H_{2}(W_{n},\mathbb{Z})\oplus\mathbb{Z}^{2g+b-1}&\text{ for }m=2,\\ H_{1}(\Sigma=M_{red},\mathbb{Z})\cong\mathbb{Z}^{2g+b-1}&\text{ for }m=1.\end{cases}$
    In particular the homology groups of $(M,\omega,G)$ is completely determined by the triple $(k=n+1,g,b)$ with $k\geq 0$, $g\geq 0$, $b\geq 1$, and $g+b\geq 2$.
Parr 2: Stein structure on $(M,\omega=d\lambda,G)$
Below we study the existence of Stein structure on $(M,\omega=d\lambda,G)$. Recall the following theorem due to Eliashberg [2] (see also [3] Theorem 1.3 and [4] Theorem 2.3) about the existence of a Stein structure on a 4-manifold via handle attaching criteria, which can be stated as the following after Gompf ([4] Theorem 2.3):
Theorem 6.1 (Eliashberg.).
An oriented 4-manifold admits a Stein structure if and only if it is diffeomorphic to the interior of a handlebody whose handles all have index $\leq$ 2, and for which each 2-handle is attached along a Legendrian knot (in the standard contact structure on the relevant boundary 3-manifold) with framing obtained from the contact framing by adding one left twist.
Here a 4-dimensional 2-handle H is diffeomorphic to a product space modeled on $D^{\prime}\times D^{\prime\prime}_{\epsilon}\subset\mathbb{R}^{2}_{x_{1},x_{2}}\times i\mathbb{R}^{2}_{y_{1},y_{2}}\subset\mathbb{C}^{2}$, where $D^{\prime}=\{x_{1}^{2}+x_{2}^{2}\leq 1\}$, called the (Lagrangian) core disc of H, and $D^{\prime\prime}=\{y_{1}^{2}+y_{2}^{2}<\epsilon\}$ for some $\epsilon>0$. H can be viewed as a trivial $D^{2}$-bundle over $D^{\prime}$. The basis normal vector fields $\{\partial_{y_{1}},\partial_{y_{2}}\}$ over $D^{\prime}$, when restricted to $\partial D^{\prime}$, gives the canonical framing of the symplectic normal bundle of $\partial D^{\prime}\subset\textsf{H}$. In particular, we can take either $\partial_{y_{1}}$ or $\partial_{y_{2}}$ as the framing of the $D^{2}$-bundle over $\partial D^{\prime}$.
Below we will show that the handle decomposition of $B=D_{0}\cup(\cup_{i}\tilde{H}_{\tau_{i}})\cup(\cup_{j}H_{C_{j}})$ lifts to a corresponding handle decomposition of $M$ as a union of handles
$M=\textsf{H}_{0}\cup(\cup_{i}\textsf{H}_{V_{i}})\cup(\cup_{j}(\textsf{H}_{\gamma_{j}}\cup\textsf{H}_{U_{j}}))$
where $\textsf{H}_{0}=W_{0}=\EuFrak{r}^{-1}(D_{0})$ is a 0-handle, $\textsf{H}_{V_{i}}$ are 2-handles associated to $\tilde{H}_{\tau_{i}}$, and for each $j$, the pair $(\textsf{H}_{\gamma_{j}},\textsf{H}_{U_{j}})$, where $\textsf{H}_{\gamma_{j}}$ is a 1-handle and $\textsf{H}_{U_{j}}$ is a 2-handle, is associated to $H_{C_{j}}$. Moreover, all the 2-handles $\textsf{H}_{V_{i}}$ and $\textsf{H}_{U_{j}}$ of $M$ satisfy the contact framing condition, hence $M$ admits the structure of a Stein manifold.
Case 1: $\textsf{H}_{V_{i}}$.  Denote by $V_{i}:=\EuFrak{r}^{-1}_{0}(\tau_{i})$ the Lagrangian disc in $h^{-1}(0)$ with the boundary circle $\partial V_{i}$ attached to $\EuFrak{r}^{-1}_{0}(\partial D_{0})\subset\EuFrak{r}^{-1}_{G}(\partial D_{0})$. $V_{i}$ is the core disc of a 4-dimensional handle $\textsf{H}_{V_{i}}$ which can be identified with a subdomain of the cotangent bundle over $V_{i}$. We may assume that these handles are mutually disjoint when attached to $\EuFrak{r}^{-1}(\partial D_{0})$.
Pick any $V_{i}$ and denote it as $V$, also denote $\tau:=\tau_{i}=\EuFrak{r}_{0}(V)$, $p:=p_{i}$.
We may parametrize $V$ as
$V=\{(x_{1},x_{2})\mid x_{1}^{2}+x_{2}^{2}\leq 1\},\quad p=(0,0)\in Fix(G)$
and identify
$X|_{V}=x_{1}\partial_{x_{2}}-x_{2}\partial_{x_{1}}.$
Let $(y_{1},y_{2})$ be the fiber coordinates of the cotangent bundle $T^{*}V$ dual to $(x_{1},x_{2})$. Then $V$ can be identified with the core disc of the 4-dimensional 2-handle
$\textsf{H}_{V}:=\EuFrak{r}^{-1}_{G}(H_{\tau})\cong V\times D^{2}_{y}\subset T^{*}L,\quad D^{2}_{y}=\{(y_{1},y_{2})\mid y_{1}^{2}+y_{2}^{2}<\epsilon^{2}\}$
for some $0<\epsilon$, with $\partial_{-}\textsf{H}_{V}:=\partial V\times D^{2}_{\epsilon}$ attached to $\EuFrak{r}^{-1}(\partial D_{0})\subset\partial W_{0}$, where $W_{0}=\EuFrak{r}^{-1}(D_{0})$.
Identify $\textsf{H}_{V}$ as a subdomain of $\mathbb{C}^{2}$. Then on $\textsf{H}_{V}$ we may take $\omega=d\lambda$, with $\lambda=\frac{1}{2}\sum_{i=1}^{2}(x_{i}dy_{i}-y_{i}dx_{i})$, identify the $G$-action with the standard G-action on $\mathbb{C}^{2}$, and take the Euclidean metric as the $G$-invariant $\omega$-compatible Riemannian metric on $\textsf{H}_{V}$. Then
$\nabla h|_{\partial V}=-x_{1}\partial_{y_{2}}+x_{2}\partial_{y_{1}}.$
Note that both $X$ and $\nabla h$ are tangent to $\EuFrak{r}^{-1}(\partial D_{0})$ and hence to $\partial_{-}\textsf{H}_{V}$. The 1-form $\lambda$ restricts to a contact 1-form $\alpha$ on $\partial V\times D^{2}_{\epsilon}\subset\partial W_{0}$. $\lambda(X)=0$ along $\partial V$ so $\partial V$ is a Legendrian curve to the contact form. As the outward normal vector field to $\EuFrak{r}^{-1}(D_{0})=W_{0}$ near $\partial V$ is $-x_{1}\partial_{x_{1}}-x_{2}\partial_{x_{2}}\subset\ker\lambda|_{\partial V}$, the contact plane field along $\partial V$ is spanned by the symplectic pair $\{\nabla h,X\}$. So the contact framing along $\partial V$ (oriented by $X$) is given by the vector field $\nabla h|_{\partial V}=-x_{1}\partial_{y_{2}}+x_{2}\partial_{y_{1}}$.
Identify $V$ as the 0-section of the cotangent bundle $T^{*}V\supset\textsf{H}_{V}$. The normal bundle $N_{V}$ of $V\subset T^{*}V$ is trivial with $\{\partial_{y_{1}},\partial_{y_{2}}\}$ as the basis fields of $N_{V}$. $N_{V}$ restricted to $\partial V$ is the symplectic normal bundle $SN^{*}(\partial V)=\partial_{y_{1}}\wedge\partial_{y_{2}}|_{\partial V}$ of $\partial V$ with the natural trivialization/framing given by $\partial_{y_{1}}$.
$SN^{*}(\partial V)$ is then identified with the normal bundle $N_{\partial V/\partial W_{0}}$ of $\partial V\subset\partial W_{0}$ upon the attaching of $\partial_{-}\textsf{H}_{V}=\partial V\times D^{2}_{y}$ to $\partial W_{0}$ along $\partial V$. As we go once along $\partial V$ the contact framing $\nabla h$ makes one positive full rotation relative to the canonical framing $\partial_{y_{1}}$. In other words, the canonical framing $\partial_{y_{1}}$ along $\partial V$ is obtained from the contact framing by adding one left twist. So the natural framing along $\partial V$ given by $\partial_{y_{1}}$ is $-1$ relative to the contact framing along $\partial V$, which meets with the Stein condition on the attachment of 4-dimensional 2-handles.
The above result on the framing criteria along $\partial V\subset\partial_{-}\textsf{H}$ apply to $V_{i}$ for all $i=1,2,\dots,n$ as well. Each of the 2-handlebodies $\textsf{H}_{V_{i}}$ has its core disc $V_{i}$ attached along a Legendrian knot (in the standard contact structure on the relevant boundary 3-manifold) with framing obtained from the contact framing by adding one left twist. In particular, $W_{n}=\EuFrak{r}^{-1}(D)=\textsf{H}_{0}\cup(\cup_{i}\textsf{H}_{V_{i}})$ admits a Stein structure by Theorem 6.1.
Case 2: $\textsf{H}_{C_{j}}$.  Pick any one of the 2-dimensional handles $H_{C_{j}}$ and denote it as $H_{C}$, where $C$ is the core curve of $H_{C}$. We will see that $\EuFrak{r}^{-1}(H_{C})$ can be identified with the union of a 4-dimensional 1-handle $\textsf{H}_{\gamma}$ with core curve $\gamma\subset h^{-1}(0)$ and a 4-dimensional 2-handle $\textsf{H}_{U}$ with core disc $U\subset\cap h^{-1}(0)$, and $\partial U$ is attached to $\partial\textsf{H}_{\gamma}\cap h^{-1}(0)$.
The preimage $L:=\EuFrak{r}^{-1}_{0}(C)\subset h^{-1}(0)\cap W_{n}$ is a Lagrangian annulus $L\cong I\times S^{1}$ foliated by $G$-orbits generated by the Hamiltonian vector field $X$. We may parametrize $L$ as
$L=\{(x_{1},x_{2})\in\mathbb{R}\times\mathbb{R}/2\pi\mathbb{Z}\mid-1\leq x_{1}\leq 1,x_{2}\in\mathbb{R}\}$
so that $X=\partial_{x_{2}}$, and each integral curve of $\partial_{x_{1}}$ is a lifting of $C$ in $L$. The boundary of $L$, which is the pair of $G$-orbits given by the equations $x_{1}=\pm 1$ is attached to $\partial W_{n}$. Let
$\gamma:=\{(x_{1},x_{2})\mid-1\leq x_{1}\leq 1,\ x_{2}=0\}\subset L.$
$\gamma$ is a lifting of $C$ in $h^{-1}(0)\cap W_{n}$.
Let $(y_{1},y_{2})$ be the fiber coordinates of the cotangent bundle $T^{*}L$ dual to $(x_{1},x_{2})$. Then $\gamma$ can be identified with the core curve of the 4-dimensional 1-handle (a product of four intervals)
$\textsf{H}_{\gamma}:=\EuFrak{r}^{-1}(H_{C})\cong\gamma\times I^{3}\subset T^{*}L,\quad I^{3}=I_{|y_{1}|<\epsilon}\times I_{|x_{2}|<\epsilon}\times I_{y_{2}}$
for some $0<\epsilon\ll 1$, with $\partial_{-}\textsf{H}_{\gamma}:=\partial\gamma\times I^{3}$ attached to $\partial M_{0}$. Here $y_{2}$ parametrizes integral curves of $-\nabla h$, the negative gradient vector field of $h$ with respect to some $G$-invariant $\omega$-compatible Riemannian metric on $M$. Let
$I_{x}:=\gamma\times I_{|x_{2}|<\epsilon}=\{(x_{1},x_{2})\mid-1\leq x_{1}\leq 1,\ -\epsilon<x_{2}<\epsilon\}\subset L.$
Up to a smoothing of the corners, the complement
$U:=L\setminus I_{x}=[-1,1]\times[\epsilon,2\pi-\epsilon]\subset L$
is a Lagrangian disc in $h^{-1}(0)$ with boundary attached to $\partial W_{n}\cap h^{-1}(0)$. Observe that $U$ is the core 2-disc of the handle
$\textsf{H}_{U}:=U\times I_{|y_{1}|<\epsilon}\times I_{y_{2}},$
with $\partial_{-}\textsf{H}_{U}=\partial U\times I_{|y_{1}|<\epsilon}\times I_{y_{2}}$ attached to $\partial_{+}\textsf{H}_{\gamma}=\gamma\times\partial I^{3}$ so that
$\textsf{H}_{\gamma}\cup\textsf{H}_{U}=T^{*}L\cap\{|y_{1}|<\epsilon\}.$
Orient $\partial U=\gamma_{-}\cup\sigma_{+}\cup\gamma_{+}\cup\sigma^{-}$ counterclockwise, where
$\gamma_{-}$ =$=\{-1\leq x_{1}\leq 1,\ x_{2}=\epsilon\},$ $\quad\dot{\gamma}_{-}$ $=\partial_{x_{1}},$ (9)
$\sigma_{+}$ =$=\{x_{1}=1,\ \epsilon\leq x_{2}\leq 2\pi-\epsilon\},$ $\quad\dot{\sigma}_{+}$ $=\partial_{x_{2}}=X,$ (10)
$\gamma_{+}$ =$=\{-1\leq x_{1}\leq 1,\ x_{2}=2\pi-\epsilon\},$ $\quad\dot{\gamma}_{+}$ $=-\partial_{x_{1}},$ (11)
$\sigma_{-}$ =$=\{x_{1}=-1,\ -\epsilon\leq x_{2}\leq 2\pi-\epsilon\},$ $\quad\dot{\sigma}_{-}$ $=-\partial_{x_{2}}=-X.$ (12)
Recall the $G$-invariant primitive 1-form $\lambda$ of $\omega$. Observe that $\EuFrak{r}^{-1}(\partial D)\subset\partial W_{n}$ is a subdomain of $\partial W_{n}$, $\lambda$ restricted to $\partial W_{n}$ is a contact 1-form near $\partial L$, whose contact structure along $\partial L$ is spanned by the symplectic pair $\nabla h$ and $X$. So along $\sigma_{+}$ the contact structure is $X\wedge(-\nabla h)=\partial_{x_{2}}\wedge\partial_{y_{2}}$, and along $\sigma_{-}$ the contact structure is $-X\wedge\nabla h=-\partial_{x_{2}}\wedge(-\partial_{y_{2}})$. Along the boundary $(\partial_{+}\textsf{H}_{\gamma})\cap L=\gamma_{-}\cup\gamma_{+}$ the vector field $X$ is normal to $\partial_{+}\textsf{H}_{\gamma}$, hence along $\gamma_{-}$ the contact structure on $\partial_{+}\textsf{H}_{\gamma}$ is $\partial_{x_{1}}\wedge\partial_{y_{1}}$, spanned by the symplectic pair $\dot{\gamma}_{-}=\partial_{x_{1}}$ and $\partial_{y_{1}}$. Similarly along $\gamma_{+}$ the contact structure on $\partial_{+}\textsf{H}_{\gamma}$ is $(-\partial_{x_{1}})\wedge(-\partial_{y_{1}})$, spanned by the symplectic pair $\dot{\gamma}_{+}=-\partial_{x_{1}}$ and $-\partial_{y_{1}}$.
By smoothing the corners of $U\subset L$ we may identify $U$ with a unit disc $D_{U}=\{x_{1}^{2}+x_{2}^{2}\leq 1\}$ with boundary $\partial D_{U}$ attached to $\partial(\textsf{H}_{\gamma}\cup W_{n})$, so that along the legendrian circle $\ell:=\partial D_{U}=S^{1}=\mathbb{R}/2\pi\mathbb{Z}$,
˙$\begin{split}\dot{\ell}(\theta)&=\cos\theta\partial_{x_{2}}-\sin\theta\partial_{x_{1}},\quad\ell(0)=(1,0)\\ &=x_{1}\partial_{x_{2}}-x_{2}\partial_{x_{1}},\end{split}$
and the vector field
$\cos\theta\partial_{y_{2}}-\sin\theta\partial_{y_{1}}=x_{1}\partial_{y_{2}}-x_{2}\partial_{y_{1}}$
is the contact framing along $\ell$.
Similar to Case 1, the framing of the symplectic normal bundle $SN^{*}(\partial U)$ (which is spanned by the basis vector fields $\{\partial_{y_{1}},\partial_{y_{2}}\}$) is given by $\partial_{y_{1}}$. As we go once along $\ell$ the contact framing makes one positive full rotation relative to $\partial_{y_{1}}$. So the canonical framing $\partial_{y_{1}}$ is $-1$ relative to the contact framing along $\ell$. This result applies to all $\textsf{H}_{C_{i}}$. Combining the result from Case 1 we conclude that $M$ admits a Stein structure following Theorem 6.1. This completes the proof of Theorem 1.7.