### v4

# Special Hamiltonian $S^{1}$-actions on symplectic 4-manifolds

Abstract
In this paper we consider symplectic 4-manifolds $(M,\omega)$ with $c_{1}(M,\omega)=0$ which admit a Hamiltonian $S^{1}$-action together with a Maslov condition on orbits of the group action. We call such spaces special Hamiltonian $S^{1}$-spaces (SHam1-spaces) and denote them as $(M,\omega,\Phi,h)$ or $(M,\omega,\Phi)$, where $\Phi:S^{1}=\mathbb{R}/2\pi\mathbb{Z}\to Ham(M,\omega)$, $\Phi(t)=\phi_{t}$, $\phi_{0}=id_{M}$, is a group homomorphism, $h$ is the associated moment map. It turns out that there are no compact SHam1-spaces. To minimize topological complexity we assume that $(M,\omega=d\alpha)$ is an exact, connected open tame manifold with connected level sets $h^{-1}(c)$, $c\in\mathbb{R}$, the gradient vector field $\nabla h$ of the moment map $h$ is complete with respect to some $\omega$-compatible Riemannian metric on $M$, and all orbit spaces $h^{-1}(c)/\Phi$ are homeomorphic to a complete manifold. We also assume that the SHam1-action on $(M,\omega=d\alpha)$ is effective, semi-free, and the fixed point set is finite. SHam1-spaces with the above conditions are called simple. We classify all simple SHam1-spaces and show that all of these spaces admit the structure of a Stein surface. Moreover, we show that, a SHam1-action $\Phi$ on a 1-connected Stein surface $\mathcal{W}_{n}$ is linear near the $A_{n}$-string of Lagrangian spheres, up to a conjugation with a $\Phi$-equivariant diffeomorphism near the $A_{n}$-string of Lagrangian spheres. If $n=0,1$ the then linearity of a SHam1-action $\Phi$ can be further expanded to larger domains by employing suitable Liouville functions.
1 Introduction and main results
Hamiltonian $S^{1}$-actions on compact symplectic 4-manifolds have been classified by Karshon in [8] where she proved that all these Hamiltonian $S^{1}$-spaces are Kähler. In this article we consider a type of Hamiltonian $S^{1}$-action $\Phi:S^{1}\to Ham(M,\omega)$ on a symplectic 4-manifold $(M,\omega)$ with an extra property called the Maslov condition, which requests that along any nondegenerate orbit $C$, any $\Phi$-invariant Lagrangian subbundle in the symplectic normal bundle of $C$, treated as a loop of Lagrangian planes along $C$, has zero Maslov number. We call such a group action a special Hamiltonian $S^{1}$-action (SHam1-action) and the triple $(M,\omega,\Phi)$ a special Hamiltonian $S^{1}$-space (SHam1-space). See Definition 2.1 in §2 for the precise definition of SHam1-action. To ensure the Maslov condition is permitted we assume that $(M,\omega)$ is connected with the first Chern class $c_{1}(M,\omega)=0$, and that the SHam1-action is effective on $M$.
It turns out that the Maslov condition imposes restrictions on the topology of $M$ as well of the $S^{1}$-action.
Theorem 1.1.
Let $(M,\omega,\Phi)$ be a connected simple SHam1-space and $\Phi$ acts effectively on $M$ with moment map $h$, then
  1. (i).
    $M$ is not compact (Proposition 3.3), and
  2. (ii).
    the fixed point set $Fix(\Phi)$ is a discrete set of points (Proposition 3.4).
  3. (iii).
    Moreover, if $(M,\omega=d\lambda)$ is exact then $Fix(\Phi)\subset h^{-1}(0)$ up to addition of a constant to $h$ (Proposition 3.5).
Since a SHam1-space $(M,\omega,\Phi)$ is not compact, if we remove a codimension $\geq 1$ $\Phi$-invariant subset, e.g., any number of nonconstant $\Phi$-orbits, from $M$, then $\Phi$ restricts to a SHam1-action on the remaining symplectic manifold $(M^{\prime},\omega)$, so $(M^{\prime},\omega,\Phi)$ by itself is also a SHam1-space, but the topology of $M^{\prime}$ can be arbitrarily complicated. To avoid extra topological complexity of $(M,\omega,\Phi)$ beyond the presence of a SHam1-action, we assume that $(M,\omega=d\alpha)$ is simple, which is defined as follows:
Definition 1.2.
Let $(M,\omega=d\lambda,\Phi)$ be a connected open exact SHam1-space, and let $h$ denote the moment map associated to $\Phi$. $(M,\omega=d\lambda,\Phi)$ is simple if
  1. (i).
    $M$ is a complete manifold with respect to some Riemannian metric, and is homotopic to a finite handlebody, or equivalently, a finite CW-complex,
  2. (ii).
    the gradient vector field $\nabla h$ of the moment map $h$ is complete with respect to some $\omega$-compatible $\Phi$-invariant Riemannian metric on $M$,
  3. (iii).
    level sets $h^{-1}(c)$ are connected for all $c\in\mathbb{R}$, and all reduced spaces $h^{-1}(c)/\Phi$ are homeomorphic to a noncompact surface without punctures.
In this paper we consider the classification problem of all (connected) simple exact SHam1-spaces $(M,\omega=d\lambda,\Phi,h)$ with finite fixed point set $Fix(\Phi)$. All SHam1-sapces considered in this manuscript are assumed to be simple.
For the 1-connected case we obtain the following topological classification results.
Theorem 1.3.
Let $(M,\omega,\Phi)$ be a connected, 1-connected simple exact SHam1-space, and the $\Phi$-action is semi-free with $n+1$ fixed points for some integer $n\geq 0$. If $n>0$ then $(M,\omega)$ is symplectically the linear plumbing of $n$ cotangent bundles of spheres $S_{i}\subset M$, i=1,,ni=1,...,n, where the union $\cup_{i=1}^{n}S_{i}$ is an $A_{n}$-string of $\Phi$-invariant Lagrangian spheres. If $n=0$ then $(M,\omega)$ is diffeomorphic to the standard symplectic 4-space $(\mathbb{R}^{4},\omega_{0}=\sum_{i=1^{2}}(dx_{i}\wedge dy_{i})$.
For $n\in\mathbb{N}\cup\{0\}$ let $\mathcal{W}_{n}\subset\mathbb{C}^{3}$ denote the Stein surface defined by the equation
$z_{1}^{2}+z_{2}^{2}+z_{3}^{n+1}=1.$
Let $\omega_{n}$ denote the standard symplectic structure on $\mathbb{C}^{3}$ restricted to $\mathcal{W}_{n}$. It is known that for $n>0$ $\mathcal{W}_{n}$ is topologically the plumbing of $n$ copies of cotangent bundles $T^{*}S^{2}$ (of Lagrangian spheres) of type $A_{n}$ [16], and $\mathcal{W}_{0}$ is symplectomorphic to the standard symplectic 4-space.
Proposition 1.4.
Let $\tilde{\Theta}$ denote the $S^{1}$-action on $\mathbb{C}^{3}$ defined by
$(z_{1},z_{2},z_{3})\to(z_{1}\cos\theta-z_{2}\sin\theta,z_{1}\sin\theta+z_{2}\cos\theta,z_{3}),\quad\theta\in\mathbb{R}/2\pi\mathbb{Z}.$
The action of $\tilde{\Theta}$ preserves $\mathcal{W}_{n}$. We denote by $\Theta$ the induced action on $\mathcal{W}_{n}$. $\Theta$ acts on $(\mathcal{W}_{n},\omega_{n})$ as a SHam1-action with $n+1$ Fixed points, and $\textsf{h}:=x_{2}y_{1}-x_{1}y_{2}$ is the moment map with $Fix(\Theta)\subset\textsf{h}^{-1}(0)$.
Remark 1.5.
Note that $(\mathcal{W}_{n},\omega_{n}=d\alpha,\Theta^{-1})$ is also a SHam1-space, where $\Theta^{-1}$ is the inverse action of that of $\Theta$. More generally, if $(M,\omega,\Phi)$ is a SHam1-space, then so is $(M,\omega,\Phi^{-1})$. We will show in Proposition 3.5 that for any connected, exact, simple effective SHam1-space $(M,\omega,\Phi)$, the weight of the $\Phi$-action at its fixed points are either all equal to 1, or all equal to -1, relative to the Darboux charts at each of the fixed points.
Definition 1.6.
Two SHam1-spaces $(M_{i},\omega_{i},\Phi_{i})$, i=1,2=1,2 are symplectically equivariant if there exists a diffeomorphism $f:M_{1}\to M_{2}$ such that $f^{*}\omega_{2}=\omega_{1}$ and $^{*}\Phi_{2}=\Phi_{1}. ff$ is called an isomorphism between the two SHam-1 spaces as in [8].
Proposition 1.7.
Let $\Theta^{-1}$ denote the inverse SHam1-action of $\Theta$ on $(\mathcal{W}_{n},\omega_{n})$. Then there exists a Hamiltonian diffeomorphism $f$ of $(\mathcal{W}_{n},\omega_{n})$, such that $f^{-1}\circ\Theta^{-1}\circ f=\Theta$, i.e., actions of $\Theta$ and $\Theta^{-1}$ on $(\mathcal{W}_{n},\omega_{n})$ are equivariant up to a conjugation by a Hamiltonian diffeomorphism on $(\mathcal{W}_{n},\omega_{n})$. In particular, we can take
$f(x_{1},y_{1},x_{2},y_{2},x_{3},y_{3}):=f(-x_{1},-y_{1},x_{2},y_{2},x_{3},y_{3}),$
then $f=f^{-1}$ and $f$ preserves the Lefschetz fibers of $\Theta$ (hence of $\Theta^{-1})$. See Remark 2.3 and Remark 5.1 for detail.
Theorem 1.8.
A connected, 1-connected, simple SHam1-space $(M,\omega=d\lambda,\Phi)$ with $n+1$ fixed points is symplectomorphic to the Stein surface $(\mathcal{W}_{n},\omega_{n})$. In particular, the symplectic topology of $(M,\omega)$ is determined by the cardinality of the fixed point set $Fix(\Phi)$.
Moreover, together with the following theorem by Wu [16], any $A_{n}$-string of Lagrangian spheres in $(M,\omega)$ is Hamiltonian isotopic to any given $A_{n}$-string of Lagrangian spheres associated to the $\Theta$-action as in Theorem 1.8, up to a composition of Lagrangian Dehn twists along the latter:
Theorem 1.9 ([16]).
Any compactly supported symplectomorphism of $\mathcal{W}_{n}$ is Hamiltonian isotopic to a composition of Dehn twists along the standard spheres. In particular, $\pi_{0}(\text{Symp}_{c}(\mathcal{W}_{n}))=Br_{n+1}$. As a result, exact Lagrangians in $A_{n}$-surface singularities are isotopic to the zero section of a plumbed copy of $T^{*}S^{2}$, up to a composition of Lagrangian Dehn twists along the standard spheres.
We have the following result concerning the uniqueness of special Hamiltonian $S^{1}$-actions on $\mathcal{W}_{n}$.
Theorem 1.10.
(Linearity and symplectic equivariance of SHam1-actions) Let $\Phi$ be any semi-free SHam1-action on $(\mathcal{W}_{n},\omega_{n})$ with moment map $h$ with $n+1$ fixed points, $n\in\mathbb{Z}$, $n\geq 0$, and let $S=S_{\phi}=\cup_{i=1}^{n}S_{i}\subset\mathcal{W}_{n}$ be the associated $A_{n}$-string of Lagrangian spheres. Also let $\textsf{S}=\textsf{S}_{\Theta}$ denote the $A_{n}$-string of Lagrangian spheres in $W_{n}$ associated to $\Theta$.
  1. (i).
    There is a $\Phi$-invariant open neighborhood $U\subset\mathcal{W}_{n}$ of S (resp. the fixed point $(0,0,1)\in\mathcal{W}_{0}$ if $n=0$) and a $\Phi$-equivariant $C^{\infty}$-diffeomorphism $\EuFrak{e}_{\textsf{S}}:U\to V=\EuFrak{e}_{\textsf{S}}(U)\subset\mathcal{W}_{n}$ on which $\Phi=\EuFrak{e}_{\textsf{S}}\circ\Theta\circ\EuFrak{e}_{\textsf{S}}^{-1}$, $\omega^{\prime}:=\EuFrak{e}_{\textsf{S}}^{*}\omega_{n}$ is symplectic and $\Phi$-invariant, the $\Phi$-action on $U$ is linear with respect to the symplectic form $\omega^{\prime}$. Then by Definition 1.6 $(U,\omega^{\prime},\Phi,h)$ and $(V,\omega_{n},\Theta,\textsf{h})$ are symplectically equivariant via the symplectomorphism $\EuFrak{e}_{\textsf{S}}:(U,\omega^{\prime})\to(V,\omega_{n})$.
  2. (ii).
    If $n=0$, any semi-free SHam1-action $\Phi$ on $(\mathbb{C}^{2},\omega)$ with $Fix_{\Phi}=\{0\}$ is linear up to conjugation and symplectically equivariant to the standard linear action of $\Theta$ on open 4-ball $B_{r}(0)=\{q\in\mathbb{C}^{2}\mid|q|^{2}<r\}$ centered at 0 with any prescribed radius $\sqrt{r}$.
  3. (iii).
    If $n=1$, then $\Phi$ is symplectically equivariant to $\Theta$ on any compactly supported $\Phi$-invariant open domain containing the $A_{1}$-sphere $S\cong S^{2}$ associated to the $\Theta$-action on $\mathcal{W}_{1}$.
For the case that $(M,\omega=d\lambda,\Phi,h)$ is not 1-connected we assume that $Fix(\Phi)$ is not empty (otherwise the $\Phi$-action is free and the Maslov condition is vacuous). Then the orbit space $h^{-1}(0)/\Phi$ is topologically an oriented noncontractible surface of genus $g\geq 0$ with $b\geq 1$ boundary components, $g+b\geq 2$, and $k:=|Fix(\Phi)|=n+1\geq 1$ marked points.
Theorem 1.11.
Let $(M,\omega,\Phi)$ be a connected exact SHam1-space with $c_{1}(M,\omega)=0$ and $n+1$ fixed points, $n\geq 0$. Assume also that $M$ is not 1-connected. Then $M$ admits the structure of a Stein surface, which is topologically obtained by attaching (2g+b1)(2g+b-1)-pairs of 1- and Lagrangian 2-handlebodies to the boundary of a Stein domain $W_{n}\subset\mathcal{W}_{n}$ diffeomorphic to $\mathcal{W}_{n}$. The homology groups of $M$ are
[…]
In particular, $H_{2}(M,\mathbb{Z})$ is generated by $n$ Lagrangian spheres and 2g+b12g+b-1 monotone Lagrangian tori.
Remark 1.12 (Relation with complexity one Hamiltonian space).
The 4-dimensional exact SHam1-spaces $(M,\omega=d\lambda,\Phi)$ that we consider here are also centered complexity one spaces as defined in [9] (see also [10, 11, 12]) except that the moment maps of these spaces are not proper unless restricted to a $\Phi$-invariant domain $D$ of $M$, so that for any compact interval $a,b]\subset\mathbb{R}$ the preimage $h^{-1}([a,b])$ of the moment map is compact.
The rest of this paper is organized as follows: In §2 we give the definition of a SHam1-action on a symplectic 4-manifold $(M,\omega)$. As an example we show that an $S^{1}$-subgroup of $SU(2)$ acts on the standard symplectic 4-space $(\mathbb{R}^{4},\omega_{0})$ special Hamiltonianly. Proof of Theorem 1.1 is given in §3 where we also establish various symplectic topological properties of SHam1-spaces. Theorem 1.3 is proved in §4. Proposition 1.4 and Theorem 1.8 are verified in §5 where the $\Theta$-action is analyzed and an $A_{n}$-string of Lagrangian spheres is given. This is followed by the proof of Theorem 1.10. Theorem 1.11 is proved in §6.
2 Special Hamiltonian $S^{1}$-action
Maslov index of a loop of Lagrangian planes (see e.g. Ch.2 of [13]). Let $L$ be a Lagrangian plane field defined along an oriented closed curve $C$ in a symplectic 4-manifold $(M,\omega)$. Assume that there exists a smooth map $\psi:D\to M$ from the unit 2-disc $D$ into $M$ such that $C=\psi(\partial D)$. Fix a symplectic trivialization $\Phi$ of the pullback bundle $\psi^{*}TM$ over $D$, with which the Lagrangian bundle $\psi^{*}L$ over $\partial D$ can be identified with a loop of Lagrangian planes in the symplectic space $\mathbb{R}^{4}\cong\mathbb{C}^{2}$ or equivalently, the image of a map
$\tau:S^{1}\to\Lambda(\mathbb{C}^{2}),$
where $\Lambda(\mathbb{C}^{2})$ is the space of unoriented Lagrangian planes in $\mathbb{C}^{2}$.
The unitary group $U(2)$ acts transitively on $\Lambda(\mathbb{C}^{2})$ with isotropy group $O(2)$, i.e., $\Lambda(\mathbb{C}^{2})$ is conjugate to $U(2)/O(2)$. By picking a unitary basis field $L$ can be represented by a loop (or half of a loop if $L$ not orientable) of unitary matrices $A$ in $U(2)$. The Maslov index $\mu(\tau)=\mu(L,\psi)$ is then defined to be the degree of the composed map
(1) ${\det}^{2}\circ\tau:S^{1}\overset{\tau}{\to}U(2)\overset{\det^{2}}{\to}U(1)\cong S^{1}.$
Several remarks on $\mu(L,\psi)$ are in order:
  1. (i).
    If $M$ is 1-connected then $\mu(L,\psi)$ is always defined since $C$ is the boundary of some disc.
  2. (ii).
    $\mu(L,\psi)$ is independent of the choice of $\psi$.
  3. (iii).
    If $c_{1}(M,\omega)=0$ Then $\mu(L,\psi)$ is independent of the class $\psi]\in\pi_{2}(M,C)$ and hence the choice of $\psi$.
Definition 2.1 (Special Hamiltonian $S^{1}$-action).
Let $\Phi$ be an $S^{1}$-group of Hamiltonian diffeomorphisms acting on $M$ effectively with nonempty fixed point set $Fix(\Phi)$. Let $h\in C^{\infty}(M,\mathbb{R})$ denote the moment map of $\Phi$, and $X=X_{h}$ the Hamiltonian vector field on $M$ which generates the $\Phi$-action so that
$\omega(X,\cdot)=-dh.$
$h:M\to\mathbb{R}$ is also called a Hamiltonian function of $(M,\omega)$. For $p\not\in Fix(\Phi)$ let $C:=Orb_{\Phi}(p)$ denote the $\Phi$-orbit of $p$. Take a nonzero $v\in T_{p}M/T_{p}C$ so that the plane $X|_{p}\wedge v$ spanned by $X|_{p}$ and $v$ is Lagrangian. Extend $X|_{p}\wedge v$ to an oriented $\Phi$-invariant Lagrangian plane field $L_{p}$ along $C$ via the $G$-action. The Maslov index $\mu(L_{p})$ is independent of the choice of nonzero $v\in(T_{p}M/T_{p}C)$ with $\omega_{p}(X,v)=0$. Also $\mu(L_{p})=\mu(L_{q})$ for $q\in Orb_{\Phi}(p)$. We say that the $\Phi$ is special if
$\mu(L_{p})=0,\quad p\not\in Fix(\Phi).$
Example 2.2.
Let $M=\mathbb{R}^{4}=\mathbb{C}^{2}$ with the standard symplectic structure $\omega_{0}=\sum_{j=1}^{2}dx_{j}\wedge dy_{j}=d\lambda_{0}$, $\lambda_{0}=\frac{1}{2}(\sum_{j=1}^{2}x_{j}dy_{j}-y_{j}dx_{j})$ and the standard Euclidean inner product. We claim that
$\Theta:=\Big\{\Theta_{\theta}=\begin{pmatrix}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{pmatrix}\Big|\theta\in\mathbb{R}/2\pi\mathbb{Z}\Big\}\subset SU(2)$
is a semi-free special Hamiltonian $S^{1}$-group acting on $\mathbb{C}^{2}$ with $Fix(\Theta)=\{0\}$.
It is easy to see that this $\Theta$-action is semi-free with $Fix(\Theta)=\{0\in\mathbb{C}^{2}\}$, and is Hamiltonian with moment map
$\textsf{h}:\mathbb{C}^{2}\to\mathbb{R},\quad\textsf{h}(z_{1},z_{2})=\Im(z_{1}\bar{z}_{2})=x_{2}y_{1}-x_{1}y_{2},$
the corresponding Hamiltonian vector field is
$X=-x_{2}\partial_{x_{1}}+x_{1}\partial_{x_{2}}-y_{2}\partial_{y_{1}}+y_{1}\partial_{y_{2}}.$
Maslov condition. We claim that $\Theta$ is special: Since the set of nontrivial orbits of $\Theta$ is connected and the Maslov index is discrete, it suffices to show that the Maslov condition holds for a typical orbit of $\Theta$, say the loop $C:=Orb_{\Theta}(p)$ with $p=(1,0)\in\mathbb{C}^{2}$,
$C(t)=(\cos t,\sin t)\subset\mathbb{C}^{2},t\in\mathbb{R}/2\pi\mathbb{Z}.$
Then $\dot{C}(0)=X_{\textsf{h}}|_{p}=\partial_{x_{2}}$. Take $v=-\partial_{x_{1}}|_{p}\in T_{p}\mathbb{C}^{2}$. Apply $\Theta$ to $X_{\textsf{h}}|_{p}\wedge v=\partial_{x_{2}}\wedge(-\partial_{x_{1}})|_{p}$. The resulting Lagrangian plane field along $C$ is $L:=\partial_{x_{1}}\wedge\partial_{x_{2}}$. The pair $(\partial_{x_{1}},\partial_{x_{2}})$ is a unitary basis field of the tangent bundle $T\mathbb{C}^{2}$, also a unitary basis field of $L$. Therefore under the map $\tau$ as in (1) $L$ can be represented by the constant loop $\begin{bmatrix}1&0\\ 0&1\end{bmatrix}\subset U(2)$ with constant determinant $1$, hence the Maslov index of $L$ is $\mu(L)=0$. This confirms that the $\Theta$-action on $\mathbb{C}^{2}$ is special. $\Box$
Morse index of h at 0. Note that $d\textsf{h}$ vanishes at $\{0\}=Fix(\Theta)$. The Hessian matrix of h at the point 0, with respect to the basis $(\partial_{x_{1}},\partial_{x_{2}},\partial_{y_{1}},\partial_{y_{2}})$, is
$\text{Hess}_{0}(\textsf{h})=\begin{bmatrix}0&0&0&-1\\ 0&0&1&0\\ 0&1&0&0\\ -1&0&0&0\end{bmatrix}$,
which is nondegenerate with eigenvalues 1,1,1,1-1,-1,$1,$ so h is a Morse function on $\mathbb{C}^{2}$, and the Morse index of h at $0\in\mathbb{C}^{2}$ is 2. The corresponding (-1)- and 1-eigenspaces are
$E_{-1}=Span\{\partial_{x_{1}}+\partial_{y_{2}},\ \partial_{y_{1}}-\partial_{x_{2}}\},\quad E_{1}=Span\{\partial_{x_{1}}-\partial_{y_{2}},\ \partial_{y_{1}}+\partial_{x_{2}}\}$
respectively. Also the two 2-dimensional vector spaces $E_{-1}$ and $E_{1}$ can be identified with a pair of complex lines (and hence $\omega_{0}$-symplectic planes) in $\mathbb{C}^{2}$ intersecting transversally at $0\in\mathbb{C}^{2}$:
$E_{-1}=\{z_{1}+iz_{2}=0\},\quad E_{1}=\{z_{1}-iz_{2}=0\}.$
Stable/unstable submanifolds of h. Let $\nabla\textsf{h}$ denote the gradient vector field of h relative to the standard Euclidean metric on $\mathbb{R}^{4}\cong\mathbb{C}^{2}$, which is both $\omega_{0}$-compatible and G-invariant. Then
$\nabla\textsf{h}=-y_{2}\partial_{x_{1}}+x_{2}\partial_{y_{1}}+y_{1}\partial_{x_{2}}-x_{1}\partial_{y_{2}}.$
Let $W_{\textsf{h}}^{s}(0)$ and $W_{\textsf{h}}^{u}(0)$ denote the stable and unstable submanifolds of of the flow of $\nabla\textsf{h}$ at $0\in\mathbb{R}^{4}\cong\mathbb{C}^{2}$ respectively. We claim that
$W_{\textsf{h}}^{s}(0)=E_{-1}=\{z_{1}+iz_{2}=0\},\quad W_{\textsf{h}}^{u}(0)=E_{1}=\{z_{1}-iz_{2}=0\}.$
Consider the point $(1,i)\in\mathbb{C}^{2}$ on $E_{-1}$, the integral curve of $\nabla\textsf{h}$ passing through the point $(1,i)$ at time $t=0$ is $\ell_{(1,i)}(t)=(e^{-t},ie^{-t})\subset E_{-1}$, the half $\mathbb{R}$-space spanned by $\nabla\textsf{h}(1,i)=-(\partial_{x_{1}}+\partial_{y_{2}})$. The $\Theta$-action preserves the space of integral curves of $\nabla\textsf{h}$. In particular $\Theta_{\pi/2}(\ell_{(1,i)})=\ell_{(-i,1)}\subset E_{-1}$ is the integral curve of $\nabla\textsf{h}$ passing through the point (i,1)(-i,1) at time $t=0$, also a half $\mathbb{R}$-space spanned by $\nabla\textsf{h}(-i,1)=\partial_{y_{1}}-\partial_{x_{2}}\subset E_{-1}$. Moreover, $Orb_{\Theta}(\ell_{(1,i)})=E_{-1}\setminus\{0\}$. So $W_{\textsf{h}}^{s}(0)=E_{-1}=\{z_{1}+iz_{2}=0\}$.
The case for $W_{\textsf{h}}^{u}(0)$ can be verified in a similar way. Consider the point $(1,-i)\in\mathbb{C}^{2}$ on $E_{1}$, the integral curve of $\nabla\textsf{h}$ passing through (1,i)(1,-i) at time $t=0$ is $\ell_{(1,-i)}(t)=(e^{t},-ie^{t})\subset E_{1}$, the half $\mathbb{R}$-space spanned by $\nabla\textsf{h}(1,-i)=\partial_{x_{1}}-\partial_{y_{2}}$. $\Theta_{\pi/2}(\ell_{(1,-i)})=\ell_{(i,1)}\subset E_{1}$ is the integral curve of $\nabla\textsf{h}$ passing through the point $(i,1)$ at time $t=0$, also a half $\mathbb{R}$-space spanned by $\nabla\textsf{h}(i,1)=\partial_{y_{1}}+\partial_{x_{2}}\subset E_{1}$. Moreover, $Orb_{\Theta}(\ell_{(1,-i)})=E_{1}\setminus\{0\}$. So $W_{\textsf{h}}^{u(0)}=E_{1}=\{z_{1}-iz_{2}=0\}$. $\Box$
Remark 2.3.
Observe that $\Theta^{-1}$ is the inverse action of $\Theta$ on $\mathbb{C}^{2}$, with the pair $(-X,-\textsf{h})$ as its corresponding Hamiltonian vector field and moment map instead, and orbits of the $\Theta^{-1}$-action are precisely the orbits of $\Theta$-action but with the opposite orientation. Nevertheless, $\Theta$ and $\Theta^{-1}$ act on $\mathbb{C}^{2}$ symplectically equivariant. For example, consider the symplectic diffeomorphism $f_{0}:\mathbb{C}^{2}\to\mathbb{C}^{2}$ given by
$f_{0}(x_{1},y_{1},x_{2},y_{2})=(-x_{1},-y_{1},x_{2},y_{2}),$
$f_{0}$ is a linear symplectic map which can be represented by the unitary matrix
$\begin{bmatrix}-1&0\\ 0&1\end{bmatrix}$,
and $f_{0}=f_{0}^{-1}$. A direct computation shows that
$f_{0}^{-1}\circ\Theta^{-1}\circ f_{0}=\begin{bmatrix}-1&0\\ 0&1\end{bmatrix}\begin{bmatrix}\cos\theta&\sin\theta\\ -\sin\theta&\cos\theta\end{bmatrix}\begin{bmatrix}-1&0\\ 0&1\end{bmatrix}=\begin{bmatrix}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{bmatrix}=\Theta.$
So $\Theta$ and $\Theta^{-1}$ are equivariant SHam-1 actions on $\mathbb{C}^{2}$. $\Box$
3 Some topological constraints imposed by Maslov condition
It turns out that at every isolated fixed point $p\in Fix(\Phi)$ of a special Hamiltonian $S^{1}$-space $(M,\omega,\Phi)$, the $\Phi$-action is Hamiltonian isotopic to the standard $\Theta$-action in a Darboux chart of $p$.
Proposition 3.1.
Let $(M,\omega,\Phi)$ be a SHam1-space with moment map $h$, and $p\in Fix(\Phi)$ an isolated fixed point of the $\Phi$-action.
  1. (i).
    At $p$ the $\Phi$-action induces a linear symplectic $S^{1}$-action on the tangent space $T_{p}M$ which can be identified with $\Theta\subset SU(2)$ with respect to a suitable symplectic Darboux chart of $p\in M$.
  2. (ii).
    There is a $\Phi$-invariant chart $(V,x_{1},x_{2},y_{1},y_{2})$ centered at $p$ and $\Phi$-equivariant with respect to a linear action of $\Phi$ on $\mathbb{R}^{4}\cong\mathbb{C}^{2}$ such that $\omega|_{V}=\sum_{k=1}^{2}dx_{k}\wedge dy_{k}$.
  3. (iii).
    All isolated critical points of $h$ are nondegenerate with Morse index $2$.
Proof.
(i). Parametrize $\Phi$ by $t\in\mathbb{R}/2\pi\mathbb{Z}$ so that $\Phi_{0}$ is the identity map. Fix $p\in Fix(\Phi)$. By Darboux-Weinstein Theorem we can symplectically identify a small open neighborhood $V\subset M$ of $p$ with a open neighborhood of the origin $0\in(\mathbb{R}^{4},\omega_{0})$ of the standard symplectic $\mathbb{R}^{4}$ via a local diffeomorphism $\phi:(V,\omega)\to(\phi(V)\subset\mathbb{R}^{4},\omega_{0})$ with $\phi(p)=0$ and $\phi^{*}\omega_{0}=\omega$. Then the linear map $A_{t}:=(\Phi_{t})_{*}|_{p}:T_{p}M\to T_{p}M$ is symplectic, i.e., $A_{t}\in Sp(4,\mathbb{R})$ for all $t$, with $A_{0}$ being the identity map. Since $\Phi_{t}\circ\Phi_{s}=\Phi_{t+s}=\Phi_{s}\circ\Phi_{t}$ for all $s,t$, we have $A_{t}A_{s}=A_{s}A_{t}$ for all $s,t$ as a result. I.e., $A_{t}$, $t\in\mathbb{R}/2\pi\mathbb{Z}$, is an $S^{1}$-subgroup of $Sp(4,\mathbb{R})$. Up to conjugation (or equivalently a choice of $\omega_{0}$-compatible complex structure) we may assume that $A$ is an $S^{1}$-subgroup of the unitary group $U(2)=Sp(4,\mathbb{R})\cap O(4,\mathbb{R})$ with respect to the standard complex structure $J_{0}$ on $\mathbb{R}^{4}\cong\mathbb{C}^{2}$. Now $A$ is either the centralizer $S^{1}$-group $C$ of $U(2)$ or an $S^{1}$-subgroup of $SU(2)$.
If $A=C$ then $A=\Big\{A_{t}=\begin{bmatrix}e^{it}&0\\ 0&e^{it}\end{bmatrix}\Big|t\in\mathbb{R}/2\pi\mathbb{Z}\Big\}$. The linear $A$-action is generated by the Hamiltonian vector field $X^{A}:=\sum_{k=1}^{2}(-y_{k}\partial_{x_{k}}+x_{k}\partial_{y_{k}})$. To check the Maslov condition, pick a point $q\not\in Fix(A)$ say, $q=(-i,0)\in\mathbb{C}^{2}$ where $X^{A}_{q}=\partial_{x_{1}}$. Let $v_{q}=\partial_{x_{2}}$. The Lagrangian plane field $L=Orb_{A}(\pm\partial_{x_{1}}\wedge\pm\partial_{x_{2}})$ along $\gamma:=Orb_{A}(q)$ is $L(t)=(\cos t\partial_{x_{1}}+\sin t\partial_{y_{1}})\wedge(\cos t\partial_{x_{2}}+\sin t\partial_{y_{2}})$, $t\in\mathbb{R}/2\pi\mathbb{Z}$. With its unitary basis field $(\cos t\partial_{x_{1}}+\sin t\partial_{y_{1}},\cos t\partial_{x_{2}}+\sin t\partial_{y_{2}})$ $L$ can be represented by the loop of unitary matrices $\begin{bmatrix}e^{it}&0\\ 0&e^{it}\end{bmatrix}$. By (1) the Maslov index of $L$ is $\mu(L)=4\neq 0$. So $A\neq C$. Similarly $A\neq C^{-1}$ as otherwise we would have $\mu(Orb_{A})(\partial_{x_{1}}\wedge\partial_{x_{2}})=-4\neq 0$. So $A$ is an $S^{1}$-subgroup of $SU(2)$. Since all $S^{1}$-subgroups of $SU(2)$ are conjugate we have $A=\Theta$ or $\Theta^{-1}$ up to a change of the complex coordinates of $\mathbb{C}^{2}$ via a special unitary linear map given by some element of $SU(2)$.
As the set $Fix(\Phi)$ is discrete without accumulation points we may choose for each $p\in Fix(\Phi)$ a symplectic Darboux chart $U_{p}$ such that $U_{p}\cap U_{q}=\emptyset$ for $p,q\in Fix(\Phi)$ with $p\neq q$. The above result about $A$ applies to every $p\in Fix(\Phi)$ as a result. This finishes the proof of (i).
(ii) It follows from (i) that an equivariant version of Darboux Theorem [1] (see also [6] Theorem 22.2) applies to every isolated fixed point $p\in Fix(\Phi)$. Namely, for each isolated fixed point $p\in Fix(\Phi)$, there exists a $\Phi$-invariant chart $V$ centered at $p\in M$ and $\Phi$-equivariant with respect to a linear action of $\Phi$ on $\mathbb{R}^{4}$ so that $\omega|_{V}=\omega_{0}$ is the standard symplectic form $\omega_{0}=\sum_{k=1}^{2}dx_{k}\wedge dy_{k}$ on $\mathbb{R}^{4}$, and $f^{-1}\circ\Phi\circ f=\Theta\subset SU(2)$.
For (iii), observe that $dh|_{p}=-\omega(X_{h}|_{p},\cdot)$ for $p\in M$. Since $\omega$ is nondegenerate, $\omega(X_{h}|_{p},\cdot)=0$ iff $X_{h}$ vanishes at $p$, i.e., iff $p\in Fix(\Phi)$, i.e., $Crit(h)=Fix(\Phi)$. By (ii) we may assume that $\Phi=\Theta$ and hence $h=\textsf{h}$ near each isolated critical point $p\in Fix(\Phi)$ of $h$, therefore all isolated critical points of $h$ are nondegenerate with Morse index $2$. ∎
Forgetting the Maslov condition, 4-dimensional SHam1-spaces are by themselves (not necessarily compact) Hamiltonian $S^{1}$-spaces. Recall that 4-dimensional compact Hamiltonian $S^{1}$-spaces have been classified by Karshon [8]:
Theorem 3.2 ([8], Theorem 6.3).
Every compact 4-dimensional Hamiltonian $S^{1}$-space $M$ can be obtained by a sequence of $S^{1}$-equivariant symplectic blow-ups from one of the following two types of ”minimal model spaces” at their fixed points:
  1. (i).
    the complex projective plane $\mathbb{C}P^{2}$ or a Hirzebruch surface, with a symplectic form that comes from a Kähler form, if $M$ has at most one fixed surface (a symplectic sphere); or
  2. (ii).
    a ruled manifold, i.e. a smooth manifold which is topologically an $S^{2}$-bundle over a closed surface, with a compatible $S^{1}$-action that fixes the base surfaces and rotates the fibers ([8] Definition 6.13, Lemma 6.15). A ruled manifold admits a compatible Kähler structure, has two fixed surfaces and no interior fixed points .
Recall that if $(M,\omega,\Phi)$ is a 4-dimensional compact Hamiltonian $S^{1}$-space then each connected component of $Fix(\Phi)$ is either a single point or a symplectic surface, and the maximum and minimum of the moment map $h$ is each attained on exactly one component of $Fix(\Phi)$ (see e.g. [8] Appendix A).
The following two propositions together verify Theorem 1.1.
Proposition 3.3.
A closed compact symplectic 4-manifold $(M,\omega)$ does not admit a SHam1-action.
Proof.
Assume in contrary that $(M,\omega)$ allows a special SHam1-action $\Phi:S^{1}\to Ham(M,\omega)$. We may assume that the action is effective. Let $h$ denote its moment map. Then $Fix(\Phi)=Crit(h)$. $M$ is compact, $h$ is bounded, and the sets $M_{-}:=\{p\in M\mid h(p)=\min h\}$ and $M_{+}:=\{p\in M\mid h(p)=\max h\}$ are both nonempty. Since every isolated fixed point of the $\Phi$-action has Morse index 2, the set $M_{-}$ contains no connected component with isolated fixed points, and neither does $M_{+}$. Then both $M_{-}$ and $M_{+}$ are connected compact closed symplectic surfaces and, following [8], $M$ can be obtained from a ruled manifold with two fixed surfaces and no interior fixed points by a sequence of equivalent blowups at fixed points that are not minima for the moment map of the ruled manifold.
Suppose that $M$ is a ruled manifold, then the orbit $\gamma_{q}:=Orb_{G}(q)$ of $q\not\in Fix(G)$ is contained some symplectic $S^{2}$-fiber $S$ of $M$. But $c_{1}(M)([S])=S\cdot S+e(S)=0+2\neq 0$, $c_{1}(M.\omega)\neq 0$. Indeed, let $F_{p}$ denote the $S^{2}$ fiber over $p\in M_{-}$. Fix any $q\in F_{p}\setminus(M_{+}\cup M_{-})$. $Orb_{\Phi}(q)$ divides $F_{p}\cong S^{2}$ into two discs $D_{-}\cup D_{+}$ with $D_{-}\cap M_{-}=\{p\}$. We may identity a $\Phi$-invariant symplectic Darboux neighborhood of $p\in M$ as $D^{\prime}\times D_{\epsilon}\subset\mathbb{C}_{z_{1}}\times\mathbb{C}_{z_{2}}$ so that $D_{-}\subset D^{\prime}\times\{0\}$, and $D_{\epsilon}$ is a coordinate chart of $F_{p}$ centered at $p$. Up to a choice of orientation of $\gamma_{q}$ we may assume that the tangent vector field along $\gamma_{q}$ is $X:=-y_{1}\partial_{x_{1}}+x_{1}\partial_{y_{1}}$. Let $v:=\partial_{x_{2}}|_{\gamma_{q}}$ be the vector field $\partial_{x_{2}}$ restricted to $\gamma_{q}$. Then $L:=X\wedge v$ is a Lagrangian plane field along $\gamma_{q}$ with Maslov index $\mu(L,D_{-})=2$. On the other hand, if we choose $D_{+}$ as the disc bounded by $\gamma_{q}$, then the orientation of $\gamma_{q}$ as the boundary of $D_{+}$ is given by X-X, and the corresponding Maslov index becomes $\mu(L=(-X)\wedge v,D_{+})=-2$. Either way the Maslov condition is not satisfied. So $M$ cannot be a ruled manifold, nor can it be obtained from any ruled manifold by a sequence of equivalent blowups at fixed points that are not minima for the moment map of the ruled manifold. We conclude that a 4-dimensional compact symplectic manifold $(M,\omega)$ does not admit an effective special Hamiltonian $S^{1}$-action. ∎
Proposition 3.4 (Fixed point set is discrete).
Let $(M,\omega)$ be a connected noncompact symplectic 4-manifold. Suppose that there is a special Hamiltonian $S^{1}$-action $\Phi:S^{1}\to Ham(M,\omega)$ on $M$ with contractible orbits and nonempty fixed point set $Fix(\Phi)$, then $Fix(\Phi)$ consists of isolated points.
Proof.
Let $\Phi:S^{1}\to Ham(M,\omega)$ be a semi-free SHam1-action on $(M,\omega)$. Denote by $X$ the Hamiltonian vector field on $M$ which generates the $\Phi$-action, and $h:M\to\mathbb{R}$ the corresponding moment map defined by $\omega(X,\cdot)=-dh$. By assumption all $\Phi$-orbits are contractable and $Fix(\Phi)$ is not empty. It is known that every connected component of $Fix(\Phi)$ is either a single point or a symplectic surface ([8], Appendix A). If $Fix(\Phi)$ contains a connected symplectic surface $\Sigma$ then $h$ is constant on $\Sigma$ and $h(\Sigma)$ is a local extremum of $h$.
Fix a point $p\in\Sigma$ and an open neighborhood $U_{p}\subset\Sigma$ of $p$ so that $\Phi$ acts nontrivially near $U_{p}$ except on $U_{p}$. There is a $\Phi$-invariant tubular neighborhood of $U_{p}$ which is symplectomorphic to a trivial disc bundle $E\cong D\times U_{p}$ over $U_{p}$ on which $\Phi$ acts nontrivially as rotations along fibers of $E$, fixing $U_{p}$ pointwise. Since $\Sigma$ is symplectic we may identify $E=D\times U_{p}$ with $D_{1}\times D_{2}\subset\mathbb{C}\times\mathbb{C}$ with $U_{p}=\{0\}\times D_{2}$ being the base disc centered at $p=(0,0)$, and $D_{1}\times\{pt\}$ as fibers, with $\Phi$ acts $D_{1}\times D_{2}$ by $t\cdot(z_{1},z_{2})=(e^{\pm int}z_{1},z_{2})$, $t\in\Phi\cong\mathbb{R}/2\pi\mathbb{Z}$.
Let $\gamma$ be a $\Phi$-orbit in the disc fiber over $p$ and $D_{\gamma}\subset D_{1}\times\{0\}$ the fiber disc over $p$ bounding $\gamma$. Assume that $h(\Sigma)$ is a local minimum of $h$ then $h=\frac{1}{2}|z_{1}|^{2}+constant$ on $E$ and the tangent vector field of $\gamma$ is $X=-y_{1}\partial_{x_{1}}+x_{1}\partial_{y_{1}}$. Let $v:=\partial_{x_{2}}|_{\gamma}$. Then $L:=X\wedge v$ is a loop of Lagrangian planes along $\gamma$ with maslov index $\mu(L,D_{\gamma})=2\neq 0$. On the other hand, if $h(\Sigma)$ is a local maximum of $h$ then $h=-\frac{1}{2}|z_{1}|^{2}+constant$ on $E$ and the tangent vector field of $\gamma$ is $X=y_{1}\partial_{x_{1}}-x_{1}\partial_{y_{1}}$. In this case $\mu(L,D_{\gamma})=-2\neq 0$. Either way the Maslov condition is not met. So $Fix(\Phi)$ does not contain any symplectic surface. We conclude that $Fix(\Phi)$ is a discrete set of points. ∎
Proposition 3.5.
Let $(M,\omega=d\lambda,\Phi)$ be a connected, exact, simple SHam1-space, and let $\lambda$ be a $\Phi$-invariant primitive 1-form of $\omega$. Then $Fix(\Phi)\subset h^{-1}(0)$ up to addition of a constant to $h$,
Proof.
Let $X$ be the Hamiltonian vector field of the moment map $h$ associated to the $\Phi$-action. Observe that
(2) $dh=d\lambda(X_{h},\cdot)=L_{X_{h}}\lambda-d(\lambda(X_{h}))=-d(\lambda(X_{h}))$
with $L_{X_{h}}\lambda=0$ since $\lambda$ is $\Phi$-invariant, so $\lambda(X_{h})-h$ is a constant. Up to an addition of a constant to $h$ we may assume that
(3) $h=\lambda(X_{h})\quad\text{and hence \quad$\lambda(X_{h})=0$ on $h^{-1}(0)$},$
then
$Fix(\Phi)\subset h^{-1}(0)$
since $\lambda(X_{h})$ is constant on level sets of $h$. ∎
4 Simply-connected Exact Simple SHam1-spaces
The rest of this paper will focus on the classification of connected exact simple SHam1-spaces. We assume that $Fix(\Phi)$ is a nonempty finite set as before. Let
$M_{c}:=\{h^{-1}(c)\},\quad\EuFrak{r}_{c}=\EuFrak{r}|_{h^{-1}(c)}:h^{-1}(c)/\Phi,\quad c\in\mathbb{R}.$
Proposition 4.1.
Let $(M,\omega=d\lambda,\Phi)$ be a connected, 1-connected exact simple SHam1-space with $c_{1}(M,\omega)=0$, $\lambda$ being $\Phi$-invariant, $\Phi$ acts semi-freely on $M$ with moment map $h:M\to\mathbb{R}$, and $Fix(\Phi)\subset h^{-1}(0)$ is a finite set. Then the followings are true.
  1. (i).
    $h^{-1}(c)/\Phi$ is 1-connected hence homeomorphic to a disc.
  2. (ii).
    $\lambda(X_{h})=h$ on $M$. Moreover $d\lambda(X_{h},\cdot)=0$ on $h^{-1}(0)$. In particular, for $q\in h^{-1}(0)\setminus Fix(\Phi)$, the symplectic normal space $T^{\omega}_{q}(h^{-1}(0)):=\{v\in T_{q}M\mid\omega_{q}(v,u)=0,\ \forall u\in T_{q}(h^{-1}(0))\}$ is generated by $(X_{h})_{q}$.
Proof. (i) Recall that $h^{-1}(c)$ is connected for all $c$. Observe that the gradient flows of $\pm\nabla h$ induce a deformation retract from $M$ to $h^{-1}(0)$, keeping $h^{-1}(0)$ fixed all the time. So $M$ is homotopic to $h^{-1}(0)$. Since $M$ is 1-connected, so is $h^{-1}(0)$. $\EuFrak{r}_{0}:h^{-1}(0)\to h^{-1}(0)/\Phi$ is a singlar fibration over $h^{-1}(0)/\Phi$, hence $h^{-1}(0)/\Phi$ is 1-connected. Moreover, the gradient flows of $\pm\nabla h$ induce a homeomorphism between the quotient spaces $h^{-1}(c)/\Phi$ and $h^{-1}(0)/\Phi$ for all $c\in\mathbb{R}$, so $h^{-1}(c)/\Phi$ is 1-connected for all $c$ provided that $M$ is 1-connected.
(ii) $\lambda$ is a $\Phi$-invariant primitive 1-form of $\omega$, hence $0=L_{X_{h}}\lambda=d(\lambda(X_{h}))+d\lambda(X_{h},\cdot)=d(\lambda(X_{h}))-dh$. Since $d(\lambda(X_{h}))-dh=0$ on $h^{-1}(0)$, $\lambda(X_{h})$ is constant on $h^{-1}(0)$. $h^{-1}(0)$ is connected and $\lambda(X_{h})=0$ at $p\in Fix(\Phi)$ so $\lambda(X_{h})=0$ and hence $d\lambda(X_{h},\cdot)=0$ on $h^{-1}(0)$. Therefore $(X_{h})_{q}\in T^{\omega}_{q}(h^{-1}(0))$ all $q\in h^{-1}(0)\setminus Fix(\Phi)$. Since $\dim(T^{\omega}_{q}(h^{-1}(0))+\dim T_{q}(h^{-1}(0))=\dim T_{q}M=4$, $X_{h}\subset T(h^{-1}(0))$ generates the field of symplectic normal spaces of $T(h^{-1}(0)\setminus Fix(\Phi))$. $\Box$
Remark 4.2.
Property (ii) above implies that for any embedded smooth curve $\gamma\subset M_{0}/\Phi$, its preimage $L_{\gamma}:=\EuFrak{r}_{0}^{-1}(\gamma)\subset h^{-1}(0)$ is an immersed monotone Lagrangian surface of $(M,\omega)$. For example, if $\gamma$ is a circle missing the set $(Fix(\Phi)$ then $L_{\gamma}$ is an embedded monotone Lagrangian torus. If $\gamma$ is a circle which meets the set $\EuFrak{r}_{0}(Fix(\Phi))$ at exactly one point, then $L_{\gamma}$ is a Lagrangian Whitney sphere, i.e., an immersed monotone Lagrangian torus with one transversal self-intersection point. If $\gamma$ is an embedded arc which connects exactly two distinct points in $\EuFrak{r}_{0}(Fix(\Phi))$ and at its endpoint, then $L_{\gamma}$ is an embedded Lagrangian sphere.
It turns out that a 1-connected exact SHam1-space $(M,\omega,\Phi)$ with finite $Fix(\Phi)$ is symplectically the plumbing of a finite number of $T^{*}S^{2}$.
Proof of Theorem 1.3.
Let $\gamma:(-\infty,\infty)\to h^{-1}(0)/\Phi\cong\mathbb{C}$ be a smoothly embedded curve so that the complement $(h^{-1}(0)/\Phi)\setminus\gamma$ consists of two connected components, and $\gamma(t_{i})=p_{i}$ for some $0=t_{0}<t_{1}<t_{2}<\cdots<t_{n}<\infty$, where $\{p_{0},p_{1},\dots,p_{n}\}=Fix(\Phi)$. Let $\gamma_{i}:=\gamma([t_{i-1},t_{i}])$, $\gamma_{0}=\gamma((-\infty,0])$, $\gamma_{n+1}=\gamma([t_{n},\infty))$. Let $S_{i}:=\EuFrak{r}_{0}^{-1}(\gamma_{i})\subset h^{-1}(0)$, then by Proposition 4.1 (ii) $S_{i}$ is a $\Phi$-invariant Lagrangian sphere which is smooth except perhaps at the two points $p_{i-1}$ and $p_{i}$.
Let $U_{i}\subset M$ be a $\Phi$-invariant symplectic Darboux chart centered at $p_{i}$, $U_{i}\cap U_{j}=\emptyset$ for $i\neq j$. Let $U:=\cup_{i=0}^{n}U_{i}$. By shrinking $U_{i}$ if necessary we may assume that $\Phi=\Theta$ on $U$. Then $h=\textsf{h}$ and $X_{h}=X_{\textsf{h}}$ on $U$. Each $U_{i}$ is symplectically identified with a open neighborhood of $0\in\mathbb{C}^{2}$ with $p_{i}$ identified with the origin $0\in\mathbb{C}^{2}$. Recall $\EuFrak{r}_{0}=\EuFrak{r}|_{h^{-1}(0)}/\Phi$. We may assume that $\EuFrak{r}(z_{1},z_{2})=\frac{1}{2}(z_{1}^{2}+z_{2}^{2})+\gamma(t_{i})$ when restricted to each $U_{i}$, and, by perturbing $\gamma$ if necessary, that $\EuFrak{r}_{0}^{-1}(\gamma_{i})\cap(U_{i-1}\cap h^{-1}(0))=\{y_{1}=0=y_{2}\}\cap U_{i-1}$ and $\EuFrak{r}_{0}^{-1}(\gamma_{i})\cap(U_{i}\cap h^{-1}(0))=\{x_{1}=0=x_{2}\}\cap U_{i}$ are Lagrangian discs for each $i$. Then
$S_{i}:=\EuFrak{r}_{0}^{-1}(\gamma_{i}),\quad i=1,...,n,$
are smoothly embedded Lagrangian spheres,
$S_{i}\pitchfork S_{i+1}=\{p_{i}\}\ \text{ for }i=1,...,n-1,\quad S_{i}\cap S_{j}=\emptyset\ \text{ if }|i-j|\neq 1.$
$S_{i}$ can be oriented so that the intersection number is $S_{i}\cdot S_{i+1}=1$ for i=1,,n1i=1,...,n-1. The union $\mathcal{S}:=\cup_{i=1}^{n}S_{i}$ is an $A_{n}$-string of Lagrangian spheres.
By Lagrangian Neighborhood Theorem [13, 15], a tubular neighborhood $V_{i}\subset M$ of the Lagrangian sphere $S_{i}$ is symplectomorphic to a tubular neighborhood of the 0-section of the cotangent bundle $(T^{*}S_{i},\omega_{can})$ of $S_{i}$ equipped with the canonical symplectic structure. $T^{*}S_{i}$ and $T^{*}S_{i+1}$ are plumbed together near $p_{i}$ so that tangent and cotangent spaces of $S_{i}$ are identified with the cotangent and tangent spaces of $S_{i+1}$ respectively for i=1,,n1i=1,...,n-1:
$T_{p_{i}}S_{i}=\partial_{x_{1}}\wedge\partial_{x_{2}}=T^{*}_{p_{i}}S_{i+1},\quad T^{*}_{p_{i}}S_{i}=\partial_{y_{1}}\wedge\partial_{y_{2}}=T_{p_{i}}S_{i+1},$
with $T^{*}S_{i}$ and $T^{*}S_{i+1}$ plumbed near $p_{i}$ via the identification $T_{p_{i}}S_{i}\cong T^{*}_{p_{i}}S_{i+1}$ with the following correspondence on local coordinates at $p_{i}$:
$T^{*}_{p_{i}}S_{i}\ni(x,y)\to(-y,x)\in T^{*}_{p_{i}}(S_{i+1}),$
where $x=(x_{1},x_{2})$ are coordinates of $S_{i}$ near $p_{i}$, $y=(y_{1},y_{2})$ are the fiber coordinates of the cotangent space $T^{*}_{p_{i}}S_{i}$.
Indeed we can take an open cover of $\EuFrak{r}_{0}=h^{-1}(0)/\Phi\cong\mathbb{R}^{2}$ by 1-connected domains $V_{i}\subset\EuFrak{r}_{0}$, $i=1,\cdots,n$ so that $\EuFrak{r}_{0}=\cup_{i}V_{i}$, $\gamma_{i}\subset V_{i}$, $V_{i}\cap V_{j}=\emptyset$ if |ij|>1|i-j|>1, and $V_{i}\cap V_{i+1}$ is 1-connected. Then each $\EuFrak{r}^{-1}(V_{i})$ is symplectomorphic to an open neighborhood in $T^{*}S_{i}$ of the zero section $S_{i}$ of $T^{*}S_{i}$, and $V_{i}\cap V_{i+1}$ is where the plumbing of $T^{*}S_{i}$ and $T^{*}S_{i+1}$ takes place. This completes the proof of Theorem 1.3 if $|Fix(\Phi)|=n+1\geq 2$.
If $|Fix(\Phi)|=1$ then $M$ is diffeomorphic to a symplectic 4-space, and we may take the fixed point to be $0\in\mathbb{R}^{4}$. $\Box$
5 A special Hamiltonian $S^{1}$-action on $\mathcal{W}_{n}$
We start with a verification of Proposition 1.4:
For each $n\in\mathbb{N}\cup\{0\}$ consider the Stein surface in $\mathbb{C}^{3}$ given by
$\mathcal{W}_{n}:z_{1}^{2}+z_{2}^{2}+z_{3}^{n+1}=1,\quad(z_{1},z_{2},z_{3})\in\mathbb{C}^{3}.$
The standard SHam1-action $\Theta$ on $\mathbb{C}^{2}$ can be lifted to a SHam1-action on $\mathbb{C}^{3}$ which acts trivially on the $z_{3}$-coordinate:
$\tilde{\Theta}:=\Big\{\tilde{\Theta}_{\theta}=\begin{pmatrix}\cos\theta&-\sin\theta&0\\ \sin\theta&\cos\theta&0\\ 0&0&1\end{pmatrix}\mid\theta\in\mathbb{R}/2\pi\mathbb{Z}\Big\}\subset SU(3).$
The action of $\tilde{\Theta}$ preserves $\mathcal{W}_{n}$ and restricts to a special Hamiltonian $S^{1}$-action on $\mathcal{W}_{n}$, denoted as $\Theta$ for simplicity, with moment map
$\textsf{h}=\Im(z_{1}\bar{z}_{2})=x_{2}y_{1}-x_{1}y_{2}\in C^{\infty}(\mathbb{C}^{3}),$
the corresponding Hamiltonian vector field is
$X=-x_{2}\partial_{x_{1}}+x_{1}\partial_{x_{2}}-y_{2}\partial_{y_{1}}+y_{1}\partial_{y_{2}},$
and the fixed point set is
$Fix(\Theta)=\{p_{k}:=(0,0,\xi^{k})|k=0,...,n\}\subset\textsf{h}^{-1}(0),\quad\xi:=e^{\frac{2\pi i}{n+1}}.$
The tangent space of $\mathcal{W}_{n}$ at $p_{k}$ is $T_{p_{k}}\mathcal{W}_{n}=\mathbb{C}^{2}_{z_{1}z_{2}}$ for each $k$. Since h is independent of $z_{3}$, the Morse index of h at $p_{k}$ is $2$ as in Example 2.2. It follows that the $\Theta$-action on $\mathcal{W}_{n}$ is special Hamiltonian. This completes the proof of Proposition 1.4.
Remark 5.1.
(Equivarience between the $\Theta$-action and its inverse $\Theta^{-1}$ on $(\mathcal{W}_{n},\omega_{n})$.)
On $(\mathcal{W}_{n},\omega_{n})$ consider the symplectic diffeomorphism $f:\mathcal{W}_{n}\to\mathcal{W}_{n}\subset\mathbb{C}^{3}$
$f(x_{1},y_{1},x_{2},y_{2},x_{3},y_{3})=(-x_{1},-y_{1},x_{2},y_{2},x_{3},y_{3}).$
$f$ can be represented by the unitary matrix
$\begin{bmatrix}-1&0&0\\ 0&1&0\\ 0&0&1\end{bmatrix},\quad f^{-1}=f.$
Then
$f^{-1}\circ\Theta^{-1}\circ f=\begin{bmatrix}-1&0&0\\ 0&1&0\\ 0&0&1\end{bmatrix}\begin{bmatrix}\cos\theta&\sin\theta&0\\ -\sin\theta&\cos\theta&0\\ 0&0&1\end{bmatrix}\begin{bmatrix}-1&0&0\\ 0&1&0\\ 0&0&1\end{bmatrix}$
$\begin{bmatrix}\cos\theta&-\sin\theta&0\\ \sin\theta&\cos\theta&0\\ 0&0&1\end{bmatrix}=\Theta.$
So $\Theta$ and $\Theta^{-1}$ are symplectically equivariant SHam-1 actions on $\mathcal{W}_{n}$. $\Box$
Consider the standard projection
$\pi^{\prime}:\mathcal{W}_{n}\subset\mathbb{C}^{3}\to\mathbb{C},\quad\pi^{\prime}(z_{1},z_{2},z_{3})=z_{3}\quad\text{ restricted to }\mathcal{W}_{n}.$
Observe that the value of $z_{3}$ is constant along any integral curve $\ell$ of $\nabla\textsf{h}$ and hence on the cylinder $Orb_{\Theta}(\ell)$ degenerated or not. So for $z\in\mathbb{C}_{z_{3}}$ the preimage $C_{z}:=(\pi^{\prime})^{-1}(z)$ is a $\Theta$-invariant cylinder degenerate only at point $p_{k}=(0,0,\xi^{k})$ if $z=\xi^{k}$ for some $k=0,1,\dots,n$, with tangent space $T_{p}C_{z}=span\{X,\nabla\textsf{h}\}$ when $p\neq p_{k}$.
Consider the level set $\textsf{h}^{-1}(0)$. Without loss of generality, we may restrict to a $\Theta$-invariant Stein domain
$W_{n}\subset\mathcal{W}_{n},\quad W_{n}=(\pi^{\prime})^{-1}(D)\cap h^{-1}([-c,c])$
for some $c>0$, $D=\{|z|\leq r\}$ for some $r>1$ and $Fix(\Theta)\subset D$.
On $W_{n}$ consider the Riemann surface
$\Sigma=\{z_{2}=0\}\cap W_{n}\subset\textsf{h}^{-1}(0)$
which contains $Fix(\Theta)$ and is transversal to $\Theta$-orbits except at $Fix(\Theta)$. Let $\Sigma_{\theta}:=\Theta_{\theta}(\Sigma)$, then $\Sigma_{\theta+\pi}=\Sigma_{\theta}$ for each $\theta$, and $\Sigma_{\theta}\cap\Sigma_{\theta^{\prime}}=Fix(\Theta)$ if $\theta\neq\theta^{\prime}\mod\pi$. Moreover,
$\textsf{h}^{-1}(0)=\cup_{\theta\in\mathbb{R}/\pi\mathbb{Z}}\Sigma_{\theta}.$
For each $\theta$ the symplectic reduction map $\EuFrak{r}_{0}|_{\Sigma_{\theta}}:\Sigma_{\theta}\to\mathbb{C}$ is a $2:1$ branched covering map branching at $Fix(\Theta)$, with branching locus $\{\xi\in\mathbb{C}\mid\xi^{n+1}=1\}\subset\mathbb{C}$. Let $g$ and $b$ be the genus and boundary components of $\Sigma$ (hence each $\Sigma_{\theta}$), then 22gb=2(n+1)2-2g-b=2-(n+1). Observe that the subgroup $K:=\{\pm Id\}\subset\Theta$ acts on $\partial\Sigma$, $\partial\Sigma/K$ is a simple closed circle which is diffeomorphic to $\partial D$ under the map $\EuFrak{r}_{0}$. Therefore $b=1$ or $2$, $\Sigma$ is a surface of genus $g=\lfloor\frac{n}{2}\rfloor$ with b=(n+1)2gb=(n+1)-2g boundary components. More explicitly,
[…]
To see the $A_{n}$-string of Lagrangian spheres in $W_{n}\subset\mathcal{W}_{n}$, let
(4) $\gamma_{k}=\Big\{e^{i\theta}\in\mathbb{C}_{z_{3}}\Big|\frac{2(k-1)\pi}{n+1}\leq\theta\leq\frac{2k\pi}{n+1}\Big\},\quad k=1,\dots,n,$
$\gamma_{k}$ is an arc between $\xi^{k-1}$ and $\xi^{k}$. Orient $\gamma_{k}$ so that its tangents point from $\xi^{k-1}$ to $\xi^{k}$. Then $\textsf{S}_{k}:=(\EuFrak{r}_{0})^{-1}(\gamma_{k})$ is a $\Theta$-invariant Lagrangian sphere, $\hat{\gamma}_{k}:=\textsf{S}_{k}\cap\Sigma_{0}$ a simple closed curve, and $\textsf{S}_{k}\cap\Sigma_{\theta}=\Theta_{\theta}(\hat{\gamma}_{k})$, $\textsf{S}_{k}=\cup_{\theta\in\mathbb{R}/\pi\mathbb{Z}}\Theta_{\theta}(\hat{\gamma}_{k})$. The union $\cup_{k=1}^{n}\textsf{S}_{k}$ is an $A_{n}$-string of Lagrangian spheres. $\mathcal{W}_{n}$ is an $A_{n}$-surface singularity as in [16].
Proof of Theorem 1.8
Let $(M,\omega,\Phi)$ be a connected and 1-connected exact SHam1-space with $n+1$ fixed points, $n\in\mathbb{N}\cup\{0\}$, and $X$ the Hamiltonian vector field associated to $h$. Let $S=S_{1}\cup S_{2}\cdots\cup S_{n}$ be an $A_{n}$-string of Lagrangian spheres in $h^{-1}(0)\subset(M,\omega,\Phi)$ associated with $(\Phi,h)$, and $\textsf{S}:=\textsf{S}_{1}\cup\textsf{S}_{2}\cdots\cup\textsf{S}_{n}$ the $A_{n}$-string of Lagrangian spheres in $(\mathcal{W}_{n},\omega_{n},\Theta)$, say, induced from (4). There exists a diffeomorphism $\phi:S\to\textsf{S}$ so that $\phi(S_{i})=\textsf{S}_{i}$ for $i=1,\cdots,n$, and $\Theta\circ\phi=\phi\circ\Phi$ on $S$. Identify $\mathcal{W}_{n}=\cup_{i=1}^{n}T^{*}\textsf{S}_{i}=\cup_{i=1}^{n}T^{*}S_{i}$ with the plumbing of $n$ copies of the cotangent bundle $T^{*}S^{2}$ of Lagrangian $2$-spheres. By Darboux theorem we may extend $\phi$ to a symplectomorphism, also denoted as $\phi$, $\phi:U\to\phi(U)\subset\mathcal{W}_{n}$, from a neighborhood $U$ of $S_{i}\cup S_{i+1}$ to a neighborhood $\phi(U)$ of $\textsf{S}_{i}\cup\textsf{S}_{i+1}$ for each $i$. Since a diffeomorphism between two Lagrangian surfaces always extends to a symplectomorphism between their cotangent bundles, we can extend $\phi$ to a symplectomorphism $\phi:M\to\mathcal{W}_{n}$, sending $T^{*}S_{i}$ to $T^{*}\textsf{S}_{i}$ so that $\phi^{*}\omega_{n}=\omega$. $\Box$
Proof of Theorem 1.10 (i)
Recall that the $n=0$ case follows from Proposition 3.1, so we may assume that $n\geq 1$. With $\phi$ in the proof of Theorem 1.8 we may identify $(M,\omega,\Phi,h)$ with $(\mathcal{W}_{n},\omega_{n},\Phi,h)$, so that $\Phi$ acts on $(\mathcal{W}_{n},\omega_{n}=d\alpha_{n})$ as an $S^{1}$-group of Hamiltonian diffeomorphisms $\Phi_{\theta}\in Ham(\mathcal{W}_{n},\omega_{n})$, $\Phi_{0}=id$.
Both $\Phi$ and $\Theta$ actions preserve $\textsf{S}=\textsf{S}_{1}\cup\cdots\cup\textsf{S}_{n}$ with $Fix(\Phi)=Fix(\Theta)=\{p_{0},p_{1},\cdots,p_{n}\}$, $p_{0}\in\textsf{S}_{1}$, $p_{i}=\textsf{S}_{i}\cap\textsf{S}_{i+1}$ for $i=1,\cdots,n-1$, and $p_{n}\in\textsf{S}_{n}$. Since both $\Phi$ and $\Theta$ act on each $\textsf{S}_{i}$ as rotations with $p_{i-1}$ and $p_{i}$ fixed, by modifying $\Phi$ if necessary we may assume that $\Phi_{\theta}=\Theta_{\theta}$ on S for all $\theta$. The Euclidean metric $g_{0}$ on $\mathbb{C}^{3}$ restricts to a $\Theta$-invariant, $\omega_{n}$-compatible Riemannian metric on $\mathcal{W}_{n}$. Also let $g$ denote a $\omega_{n}$-compatible metric on $\mathcal{W}_{n}$ which is also $\Phi$-invariant. We may assume that $g=g_{0}$ on $T_{\textsf{S}_{i}}\mathcal{W}_{n}$ the tangent bundle of $\mathcal{W}_{n}$ over $\textsf{S}_{i}$ for all $i$. Let $N(\textsf{S}_{i})$ denote the normal bundle of $\textsf{S}_{i}\subset\mathcal{W}_{n}$ determined by the metric $g_{0}$, which can also be identified with the cotangent bundle $T^{*}\textsf{S}_{i}$. Note that
(5) $N_{p}(\textsf{S}_{i})=T_{p}\textsf{S}_{j},\qquad T_{p}\textsf{S}_{i}=N_{p}(\textsf{S}_{j})\ \text{ at $p=\textsf{S}_{i}\cap\textsf{S}_{j}$ when $|i-j|=1$}.$
Recall that the action of $\Theta$ on $\mathcal{W}_{n}$ is linear. Since $\Phi_{\theta}=\Theta_{\theta}$ on $\textsf{S}_{i}$, the linearized action $\Phi_{*}$ of $\Phi$ acts on $N(S_{i})$ in the same way as $\Theta=\Theta_{*}$ does. Denote $x_{\theta}:=\Phi_{\theta}(x)=\Theta_{\theta}(x)$ and $v_{\theta}:=(\Phi_{\theta})_{*}(v)=\Theta_{\theta}(v)\in N_{x_{\theta}}(\textsf{S}_{i})$ for $x\in\textsf{S}_{i}$, $v\in N_{x}(\textsf{S}_{i})$. We have for $v\in N_{x}(\textsf{S}_{i})$
$(\Phi_{\theta})_{*}=\Theta_{\theta}:N_{x}(\textsf{S}_{i})\to N_{x_{\theta}}(\textsf{S}_{i}),\quad v\to v_{\theta}=\Theta_{\theta}(v).$
At $p_{i}\in Fix(\Phi)$ we have
$(\Phi_{\theta})_{*}=\Theta_{\theta}:T_{p_{i}}\mathcal{W}_{n}\to T_{p_{i}}\mathcal{W}_{n}$
and the action preserves the splitting $T_{p_{i}}\mathcal{W}_{n}=T_{p_{i}}\textsf{S}_{i}\oplus N_{p_{i}}(\textsf{S}_{i})$. Below we consider exponential maps associated to $\textsf{S}\subset\mathcal{W}_{n}$.
For each $i$ consider the exponential map over $\textsf{S}_{i}$ relative to $g$, so that for $x\in\textsf{S}_{i}$ and $v\in N_{x}(\textsf{S}_{i})$,
$\EuFrak{e}_{i}:=\exp_{i}:\ N(\textsf{S}_{i})\to\mathcal{W}_{n},\quad v\to\gamma_{v}(1)$
where $\gamma_{v}(1)$ is the time $t=1$ image of the geodesic $\gamma_{v}(t)$ relative to $g$ with $\gamma_{v}(0)=x$ and $\dot{\gamma}_{v}(0)=v$.
If $n>1$ then at each $p_{i}\in Fix(\Phi)$ with $1\leq i<n$ we also include the exponential map
$\EuFrak{e}_{p_{i}}:=\exp_{p_{i}}:T_{p_{i}}\mathcal{W}_{n}\to\mathcal{W}_{n}$
relative to any $\Phi$-invariant $\omega_{n}$-compatible Riemannian metric $g$: for $v\in T_{p_{i}}\mathcal{W}_{n}$
$\EuFrak{e}_{p_{i}}(v):=\gamma_{v}(1)$
is the time $t=1$ image of the geodesic $\gamma_{v}(t)$ relative to the metric $g$ with $\gamma_{v}(0)=p_{i}$ and $\dot{\gamma}_{v}(0)=v$.
With (5) the $\EuFrak{e}_{p_{i}}$’s and $\EuFrak{e}_{i}$’s piece together well to form a differentiable map $\EuFrak{e}_{\textsf{S}}$ from the cotangent/normal bundle
$N(\textsf{S}):=\begin{cases}N(\textsf{S}_{1})&\text{if $n=1$}\\ \cup_{i=1}^{n}N(\textsf{S}_{i})\cup_{i=1}^{n-1}T_{p_{i}}\mathcal{W}_{n}&\text{if $n>1$}\end{cases}$
into $\mathcal{W}_{n}$. There exists $\epsilon>0$ so that the map
$\EuFrak{e}_{\textsf{S}}:=\exp_{\textsf{S}}:\ V:=\{v\in N(\textsf{S})\mid|v|<\epsilon\}\to\mathcal{W}_{n}$
is a diffeomorphism from $V$ onto the open set $U:=\EuFrak{e}_{\textsf{S}}(V)$. We can also consider the exponential map $\EuFrak{e}^{0}_{\textsf{S}}:=\exp^{0}_{\textsf{S}}$ relative to the metric $g_{0}$, and for $v\in N(\textsf{S})$ based at $x\in\textsf{S}$, we write $\gamma^{0}_{v}(t)$ as the geodesic relative to $g_{0}$ with $\gamma^{0}_{v}(0)=x$ and $\dot{\gamma}^{0}_{v}(0)=v$.
Since $\Phi_{*}=\Theta$ and $g=g_{0}$ on $T_{\textsf{S}}\mathcal{W}_{n}$, $\gamma_{v}$ and $\gamma^{0}_{v}$ are tangent at the base point $x$ of $v$. For $v\in N(\textsf{S})$ denote $v_{\theta}:=(\Phi_{\theta})_{*}(v)$, $v_{0}=v$, then
$(\Phi_{\theta})_{*}(tv)=tv_{\theta},\quad(\Phi_{\theta_{1}+\theta_{2}})_{*}(v)=(\Phi_{\theta_{1}})_{*}(\Phi_{\theta_{2}})_{*}(v)=(\Phi_{\theta_{1}})_{*}(v_{\theta_{2}})=v_{\theta_{1}+\theta_{2}}.$
As $g$ is $\Phi$-invariant, $\Phi_{\theta}$ maps $g$-geodesics to $g$-geodesics, preserving the arc length. Also, $\Phi_{\theta}=\Theta_{\theta}$ when restricted to S, we have $\Phi_{\theta}(\gamma_{v})=\gamma_{v_{\theta}}$, and for $v\in N(\textsf{S})$ based at $x\in\textsf{S}$, the map $\EuFrak{e}_{\textsf{S}}$ takes $\gamma^{0}_{v}(1)$ to $\gamma_{v}(1)$, and $\gamma^{0}_{v_{\theta}}(1)$ to $\gamma_{v_{\theta}}(1)$.
The domain $U=\EuFrak{e}_{\textsf{S}}(V)$ is $\Phi$-invariant since $V$ is $\Phi_{*}=\Theta$-invariant and $g=g^{0}$ on $V\subset N(\textsf{S})$. Also we may identify $(V\subset N(\textsf{S}),\omega_{n})$ with its image $(\EuFrak{e}_{\textsf{S}}(V)=U,\omega_{n})$. Then on $U$
$\Phi=\EuFrak{e}_{\textsf{S}}\circ\Theta\circ\EuFrak{e}_{\textsf{S}}^{-1}.$
Let
$\omega^{\prime}:=(\EuFrak{e}_{\textsf{S}}^{-1})^{*}\omega_{n}=(\EuFrak{e}_{\textsf{S}}^{-1})^{*}\sum_{j=1}^{3}dx_{j}\wedge dy_{j}=\sum_{j=1}^{3}dx^{\prime}_{j}\wedge dy^{\prime}_{j},$
where $x^{\prime}_{j}:=x_{j}\circ\EuFrak{e}_{\textsf{S}}^{-1}$ and $y^{\prime}_{j}:=y_{j}\circ\EuFrak{e}_{\textsf{S}}^{-1}$. $\omega^{\prime}$ is symplectic on $U$, invariant under $\Phi$-action, and the action of $\Phi$ on $U$ is linear with respect to $\omega^{\prime}=\sum_{j=1}^{3}dx^{\prime}_{j}\wedge dy^{\prime}_{j}$. Let the pair $(\textsf{X},\textsf{h})$ denote the Hamiltonian vector field and corresponding Hamiltonian function associated to the $\Theta$-action, then
$(X:=(\EuFrak{e}_{\textsf{S}})_{*}\textsf{X},\ h:=\textsf{h}\circ\EuFrak{e}^{-1}_{\textsf{S}})$
is the the corresponding Hamiltonian vector field and Hamiltonian function associated to the $\Phi$-action on $U$. We may extend $\EuFrak{e}_{\textsf{S}}$ to a diffeomorphism $\EuFrak{e}^{\prime}_{\textsf{S}}:\mathcal{W}_{n}\to\mathcal{W}_{n}$ with $\Phi$-invariant compact support containing $U$. Let
$\Phi^{\prime}:=\EuFrak{e}^{\prime}_{\textsf{S}}\circ\Theta\circ{\EuFrak{e}^{\prime}_{\textsf{S}}}^{-1}=p.$
$\Phi^{\prime}$ acts as a SHam1–action $\mathcal{W}_{n}$, with $\Phi^{\prime}=\Phi$ on $U$ and outside the compact support of $\EuFrak{e}^{\prime}_{\textsf{S}}$.
Note that $(\EuFrak{e}_{\textsf{S}})_{*}:=d\EuFrak{e}_{\textsf{S}}=id$ when restricted to S, so $\omega^{\prime}=\omega_{n}$ on S. Both $\omega^{\prime}=d\alpha^{\prime}$ and $\omega_{n}=d\alpha_{n}$ are $\Phi$-invariant exact symplectic forms. Up to averaging via the $\Phi$-action we may assume that both primitive 1-forms $\alpha_{n}$ and $\alpha^{\prime}$ are $\Phi$-invariant. Consider the smooth family of differential forms $\omega_{t}:=(1-t)\omega_{n}+t\omega^{\prime}$, $0\leq t\leq 1$. Recall that $\omega^{\prime}=\omega_{n}$ on S. Then similar to the proof of Proposition 3.1 (ii), by shrinking $U$ if necessarily, we may assume that $\omega_{t}$ are symplectic on $U$ for all $0\leq t\leq 1$, and there exists a smooth isotopy $\phi_{t}:U\to M$, $t\in[0,1]$, with $\phi_{0}=id$, $\phi_{t}=id$ on S for all $t$, and
$\phi_{1}^{*}\omega^{\prime}=\omega_{n}.$
Therefore, any effective SHam1-action on $(\mathcal{W}_{n},\omega_{n})$ is linear near S up to conjugation with a smooth isotopy $\phi_{t}$ with $\phi^{-1}_{1}\circ\Phi\circ\phi_{1}=\Theta$ on $U$. In addition $(U,\omega^{\prime},\Phi)$ and $(V,\omega_{n},\Theta)$ are symplectically equivariant via the symplectomorphism $\EuFrak{e}_{\textsf{S}}:(U,\omega^{\prime})\to(V,\omega_{n})$. $\Box$
In the following we will show that on $(\mathcal{W}_{n},\omega_{n})$ with $n=0,1$, any semi-free SHam1-action, which we denote it as $\Phi$, up to conjugations is linear and, following1.10(i), is symplectically equivariant to the $\Theta$-action on any prescribed relatively compact domain containing (i) the fixed point of the $\Theta$-action if $n=0$, and (ii) a $\Theta$-invariant Lagrangian sphere if $n=1$.
Proof of Theorem 1.10 (ii)(iii)
(ii): $n=0$. We denote $q=(x,y)$ for $q\in(\mathbb{C}^{2},\omega=\sum_{j=1}^{2}dx_{j}\wedge dy_{j})$, where $x=(x_{1},x_{2})$, $y=(y_{1},y_{2})$, also $|x|=\sqrt{x_{1}^{2}+x_{2}^{2}}$, $|y|=\sqrt{y_{1}^{2}+y_{2}^{2}}$.
Recall that $\mathcal{W}_{0}:\ z_{1}^{2}+z_{2}^{2}+z_{3}=1$. The embedding
$\iota:\mathbb{C}^{2}\to\mathbb{C}^{3},\quad\iota(z_{1},z_{2})=(z_{1},z_{2},1-z_{1}^{2}-z_{2}^{2})$
is a symplectomorphism between $(\mathbb{C}^{2},\omega=\sum_{j=1}^{2}dx_{j}\wedge dy_{j})$ and its image $(\mathcal{W}_{0},\omega_{0})$, $\mathcal{W}_{0}=\iota(\mathbb{C}^{2})$. Conversely the projection map $\pi:\mathcal{W}_{0}\to\mathbb{C}^{2}$, $\pi(z_{1},z_{2},z_{3})=(z_{1},z_{2})$ is the inverse symplectomorphism, $\pi\circ\iota=id_{\mathbb{C}^{2}}$, $\iota\circ\pi=id_{\mathcal{W}_{0}}$, $\pi^{*}\omega=\omega_{0}$.
Consider The Liouville vector field $Y:=\frac{1}{2}\sum_{j=1}^{2}x_{j}\partial_{x_{j}}+y_{j}\partial_{y_{j}}$ on $\mathbb{C}^{2}$. The differential $\iota_{*}:T\mathbb{C}^{2}\to T\mathcal{W}_{0}$ induces a Liouville vector field $Y_{0}$ on $\mathcal{W}_{0}$ given by $Y_{0}:=\iota_{*}Y$. Both $Y$ and $Y_{0}$ are $\Theta$-invariant. Recall from Proposition 3.1 $\Phi$ and $\Theta$ are symplectically equivariant when restricted to some open neighborhood $U$ of the fixed point. Since $\mathbb{C}_{2}$ and $\mathcal{W}_{0}$ are symplectomorphic $\Theta$-spaces, and any semi-free SHam1-action on $\mathcal{W}_{0}$ descends to one in $\mathbb{C}^{2}$, so it suffices to compare $\Phi$ and $\Theta$ in $(\mathbb{C}^{2},\omega)$. Then by Proposition 3.1 (ii) up to conjugation $\Phi$-action is equivariant to $\Theta$-action and hence is linear on some open neighborhood $U_{0}$ of $0\in\mathbb{C}^{2}$, i.e., there exists a diffeomorphism $\EuFrak{e}_{0}:U_{0}\to U$ from a $\Theta$-invariant open neighborhood of $0\in\mathbb{C}^{2}$ onto a $\Phi$-invariant open neighborhood $U$ of $0\in\mathbb{C}^{2}$, $\EuFrak{e}_{0}(0)=0\in\mathbb{C}^{2}$, and
$\Phi_{\theta}=\EuFrak{e}_{0}\circ\Theta_{\theta}\circ\EuFrak{e}_{0}^{-1}\quad\text{on $U$},\quad\forall\theta,$
or equivalently,
$\Theta_{\theta}=\EuFrak{e}_{0}^{-1}\circ\Phi_{\theta}\circ\EuFrak{e}_{0}\quad\text{on $U_{0}$},\quad\forall\theta.$
Let $\rho:[0,\infty)\to[0,1]$ be a $C^{\infty}$-function so that $\rho^{\prime}(t)\leq 0$ for $t\in[0,\infty)$, and for some $0<r_{1}<r_{0}<\infty$, $\rho(t)=1$ if $t\leq r_{1}$, $\rho(t)=0$ if $t\geq r_{0}$. Let $\tilde{\rho}:\mathbb{C}^{2}\to[0,1]$ be the $C^{\infty}$-function given by $\tilde{\rho}(q)=\rho(|q|^{2})$. Let $\tilde{Y}:=\tilde{\rho}Y$ and $\psi_{t}$ be the smooth isotopy on $\mathbb{C}^{2}$ associated to $\tilde{Y}$, $\psi_{0}=id$, $\frac{d\psi_{t}}{dt}=\tilde{Y}\circ\psi_{t}$. For $q=(x,y)\in\mathbb{C}^{2}$,
$\psi_{t}(q)=(e^{\frac{t}{2}\tilde{\rho}(q)}x,e^{\frac{t}{2}\tilde{\rho}(q)}y).$
For $r>0$ and $r<r_{1}$ there exists $t_{r}<0$ such that whenever $|q|^{2}\leq r$,
$\psi_{t_{r}}(x,y)=(e^{t_{r}/2}x,e^{t_{r}/2}y)\in U_{0}.$
Let $\Psi_{r}:=\EuFrak{e}_{0}\circ\psi_{t_{r}}$, then for $q=(x,y)$ with $|q|^{2}\leq r$,
(6) $\begin{split}(\Psi_{r}^{-1}\circ\Phi_{\theta}\circ\Psi_{r})(x,y)&=\psi_{t_{r}}^{-1}\circ\EuFrak{e}_{0}^{-1}\circ\Phi_{\theta}\circ\EuFrak{e}_{0}\circ\psi_{t_{r}}(x,y)\\ &=\psi_{t_{r}}^{-1}\circ\EuFrak{e}_{0}^{-1}\circ\Phi_{\theta}\circ\EuFrak{e}_{0}(e^{t_{r}/2}x,e^{t_{r}/2}y)\\ &=\psi_{t_{r}}^{-1}\circ\Theta_{\theta}(e^{t_{r}/2}x,e^{t_{r}/2}y)\\ &=\psi_{t_{r}}^{-1}(e^{t_{r}/2}x_{\theta},e^{t_{r}/2}y_{\theta})\\ &=(x_{\theta},y_{\theta})=\Theta_{\theta}(x,y)\end{split}$
where
(7) $\begin{split}x_{\theta}&:=(x_{1}\cos\theta+x_{2}\sin\theta,-x_{1}\sin\theta+x_{2}\cos\theta),\\ y_{\theta}&:=(y_{1}\cos\theta+y_{2}\sin\theta,-y_{1}\sin\theta+y_{2}\cos\theta).\end{split}$
So
$\Psi_{r}^{-1}\circ\Phi\circ\Psi_{r}=\Theta\quad\text{on }\{(q=(x,y)\in\mathbb{C}^{2}||q|^{2}\leq r<r_{1}\}\subset\mathbb{C}^{2}.$
By allowing $r_{1}$ and $r_{0}$ to be arbitrarily large we conclude the that any semi-free SHam1-action on $(\mathbb{C}^{2},\omega)$ with $Fix_{\Phi}=\{0\}$ is linear up to conjugation and symplectically equivariant to the standard linear action of $\Theta$ on open 4-ball $B_{r}(0)=\{q\in\mathbb{C}^{2}\mid|q|^{2}<r\}$ centered at 0 with any prescribed radius $\sqrt{r}$. This completes the proof of Case 1.
(iii): $n=1$. Consider $(\mathcal{W}_{1},\omega_{1})$ where $\mathcal{W}_{1}:\{z_{1}^{2}+z_{2}^{2}+z_{3}^{2}=1\}\subset\mathbb{C}^{3}$ is given by the equations
$\sum_{j=1}^{3}x_{j}^{2}=1+\sum_{j=1}^{3}y_{j}^{2},\quad\sum_{j=1}^{3}x_{j}y_{j}=0.$
For convenience we use the notations $x:=(x_{1},x_{2},x_{3})$, $y:=(y_{1},y_{2},y_{3})$, $x^{2}:=\sum_{j=1}^{3}x_{j}^{2}$ and $y^{2}:=\sum_{j=1}^{3}y_{j}^{2}$.
$\mathcal{W}_{1}$ is indeed the cotangent bundle of the sphere $S:=\{\sum_{j=1}^{3}x_{j}^{2}=1\}\subset\mathbb{R}^{3}$ which is Lagrangian viewed as a submanifold of $(\mathcal{W}_{1},\omega_{1})$, $\Theta$ acts on $S$ with two fixed points $(0,0,\pm 1)$. Consider the vector field on $\mathcal{W}_{1}$:
$Y:=\sum_{j=1}^{3}y_{j}\partial_{y_{j}}.$
The 1-form
$\alpha:=\iota_{Y}\omega_{1}=-\sum_{j=1}^{3}y_{j}\partial_{x_{j}}$
satisfies $d\alpha=\omega_{1}$, is a $\Theta$-invariant primitive 1-form of $\omega_{1}$, so $Y$ is a Liouville vector field on $\mathcal{W}_{1}$ which vanishes precisely on the Lagrangian sphere $S$. Moreover, $Y$ is invariant under the $\Theta$-action, which can be checked by direct computation.
Remark 5.2.
The vector field $Y:=\sum_{j=1}^{3}y_{j}\partial_{y_{j}}$ on $\mathcal{W}_{1}$ is indeed the gradient vector field $\nabla f$ of the $C^{\infty}$-function $f(x,y)=\sum_{j=1}^{3}y_{j}\partial_{y_{j}}$ with respect to the Riemannian metric which is the pullback of the Euclidean metric on $\mathbb{C}^{3}$ associated to the inclusion map $\iota:\mathcal{W}_{1}\to\mathbb{C}^{3}$.
It is known that all Lagrangian spheres in the cotangent bundle $T^{*}S^{2}$ are hamiltonian isotopic [7], so are those in $\mathcal{W}_{1}$. Recall from proof of Theorem 1.10 (i) that up to conjugation any semi-free SHam1-action on $\mathcal{W}_{1}$ are symplectically equivalent to the standard $\Theta$-action when restricted to some open neighborhood $U$ of the Lagrangian sphere $S$. We may assume that $(x,y)\in U$ whenever $|y|:=\sqrt{y^{2}}<\delta$ for some $\delta>0$.
Let $\eta_{t}$ be the time $t$ map of the flow of $Y$ with $\eta_{0}=id_{\mathcal{W}_{1}}$. $\eta_{t}$ is $\Theta$-invariant for all $t$. For $(x,y)\in\mathcal{W}_{1}$
$\eta_{t}(x,y)=(e^{s}x,e^{t}y),\quad e^{2s}x^{2}=1+e^{2t}y^{2},\quad s,t\in\mathbb{R}.$
Consider a $C^{\infty}$-function $\rho:[0,\infty)\to[0,1]$ such that $\rho^{\prime}\leq 0$, and for some $0<r_{1}<r_{0}<\infty$, $\rho(r)=1$ if $0\leq r\leq r_{1}$, $\rho(r)=0$ if $r_{0}\leq r$. let $\tilde{\rho}:\mathcal{W}_{1}\to[0,1]$ be defined as $\tilde{\rho}(x,y):=\rho(|y|^{2})$. Let $Y_{\rho}:=\tilde{\rho}Y$.
Let $\psi_{t}$ be the time $t$ map of the flow of $Y_{\rho}$. $\frac{d\psi_{t}}{dt}=Y_{\rho}(\psi_{t})$. $\phi_{0}=id_{\mathcal{W}_{1}}$. $\frac{d\psi_{t}}{dt}=0$ when restricted to $S$. $\psi_{t}$ is $\Theta$-equivariant for all $t$. For $(x,y)\in\mathcal{W}_{1}$,
$\psi_{t}(x,y)=(e^{s}x,e^{\rho(|y|^{2})t}y),\quad e^{2s}x^{2}=1+e^{2\rho(|y|^{2})t}y^{2}.$
For $r>0$ and $r<r_{1}$ there exists $t_{r}<0$ such that $\psi_{t_{r}}(x,y)\in U$ whenever $|y|\leq r$,
$\psi_{t_{r}}(x,y)=(e^{s_{r}}x,e^{t_{r}}y)\quad e^{2s_{r}}x^{2}=1+e^{2t_{r}}y^{2}.$
As in the proof of Theorem 1.10(i) we may assume that $\Phi=\Theta$ when restricted to the Lagrangian sphere $S:x^{2}=1,y=0$ in $\mathcal{W}_{1}$, and consider the exponential map $\EuFrak{e}_{S}$ over $S$ relative to a $\Phi$-invariant $\omega_{1}$-compatible Riemannian metric $g$ on $\mathcal{W}_{1}$:
$\EuFrak{e}_{S}:=\exp_{S}:N(S)\to\mathcal{W}_{1},\quad\EuFrak{e}_{S}(v)=\gamma_{v}(1).$
There exists $\epsilon>0$ so
$\EuFrak{e}_{S}:V=\{v\in N(S)||v|<\epsilon\}\to\mathcal{W}_{1}$
is a diffeomorphism from $V$ onto the open set $U:=\EuFrak{e}_{S}(V)$. Then on $U$
$\Phi_{\theta}=\EuFrak{e}_{S}\circ\Theta_{\theta}\circ(\EuFrak{e}_{S})^{-1},\quad\forall\theta.$
Let $F_{r}:=\EuFrak{e}_{S}\circ\psi_{t_{r}}$. Similar to the $n=0$ case, we have
$F_{r}^{-1}\circ\Phi_{\theta}\circ F_{r}=\Theta_{\theta}\quad\text{on $U_{r}:=\{(x,y)\in\mathcal{W}_{1}|\ |y|^{2}\leq r\}$}.$
By allowing $r_{1}$ and $r_{0}$ to be arbitrarily large we conclude that any semi-free special Hamiltonian $S^{1}$-action on $(\mathcal{W}_{1},\omega_{1})$ is linearly up to conjugation and symplectically equivariant to the standard linear action of $\Theta$ on $U_{r}$ for any prescribed $r>0$. This completes the proof of Theorem 1.10. $\Box$
6 Non-simply connected cases
Now we consider the case that $(M,\omega)$ is a connected exact symplectic 4-manifold with $c_{1}(M,\omega)=0$ but not 1-connected. Assume that $(M,\omega)$ equips with a semi-free special Hamiltonian $S^{1}$-action which we denote as $\Phi\subset\text{Ham}(M,\omega)$.
Since the Maslov condition applies only when nonconstant orbits of $\Phi$ are homologically trivial, we assume that $Fix(\Phi)$ is nonempty and finite to accommodate the Maslov condition . Let
$k=n+1:=|Fix(\Phi)|\geq 1$
denote the number of fixed points of the $\Phi$-action, $h:M\to\mathbb{R}$ the moment map associated to $\Phi$, and $X=X_{h}$ the Hamiltonian vector field of $h$ defined by $\omega(X,\cdot)=-dh$.
Fix any $\lambda\in\Omega^{1}(M)$ with $d\lambda=\omega$. As in the 1-connected case we may assume that $\lambda$ is $\Phi$-invariant and $Fix(\Phi)\subset h^{-1}(0)$, then $\lambda$ satisfies $\lambda(X)=h$. recall
$\EuFrak{r}_{c}:=h^{-1}(c)/\Phi,\quad\text{ also \ let}\quad\EuFrak{r}^{-1}_{[-c,c]}:=h^{-1}([-c,c])$
associated to the $\Phi$-action on $M$. $h^{-1}(c)/\Phi$ is topologically an oriented surface of genus $g$ with $k$ marked points (which are the fixed points of the $\Phi$-action) and $b\geq 1$ connected boundary components at infinity. Since $\Phi$ acts freely on $h^{-1}(c)$ for all $c\neq 0$, $\EuFrak{r}_{c}=h^{-1}(c)/\Phi$ is not 1-connected, hence
$g+b\geq 2,\quad g\geq 0,\ b\geq 1.$
Proof of Theorem 1.11
Recall that $\EuFrak{r}_{c}=h^{-1}(c)/\Phi$ and and $\EuFrak{r}_{c^{\prime}}=h^{-1}(c^{\prime})/\Phi$ are homeomorphic topological surfaces for $c,c^{\prime}\in\mathbb{R}$. Let $B_{0}:=\EuFrak{r}_{0}$. Up to homotopy equivalence we may identify $M$ with the corresponding $\Phi$-invariant subdomain $B=\EuFrak{r}^{-1}_{[-c,c]}B_{0}$ for sone $c>0$, $B$ is homotopic to $M$.
Part 1: Handle decomposition of $B$ and Homology of $M$. For each of the $k=1+n$ fixed points $p_{i}$, $0\leq i\leq n$, of the $\Phi$-action on $M$, denote by $q_{i}:=\EuFrak{r}_{0}(p_{i})\in B_{0}$ the image of $p_{i}$ under the map $\EuFrak{r}_{0}:M_{0}\to M_{0}/\Phi=B_{0}$. In $B_{0}$ pick $n+1$ open discs $D_{i}$ centered at $q_{i}$, $i=0,1,\dots,n$, such that the closures of $D_{i}$’s are pairwise disjoint closed discs in $B_{0}$. As $B_{0}$ can be constructed by attaching 2-dimensional 1-handles to 2-dimensional 0-handles, we may view $D_{i}$ as the set of 2-dimensional 0-handles of $B_{0}$, and with $q_{i}$ as the core point of $D_{i}$. Then $B_{0}$ can be obtained by attaching the following two types of 1-handles to $\cup_{i}D_{i}$:
  1. (i).
    A disjoint union of $n$ 1-handles which connect $\cup_{i=1}^{n}D_{i}$ to $D_{0}$:
    Let $\tau_{i}\subset B_{0}\setminus D_{0}$, $i=1,\dots,n$, be a set of pairwise disjoint embedded arcs with endpoints such that $\tau_{i}$ is disjoint from $q_{l}$ if $l\neq i$, $q_{i}=\partial_{+}\tau_{i}$ is one endpoint of $\tau_{i}$, and $\tau_{i}$ intersects transversally with $\partial D_{0}$ at the other endpoint $\partial_{-}\tau_{i}$ of $\tau_{i}$. Thicken each of $\tau_{i}$ a bit to get $n$ mutually disjoint 2-dimensional 1-handles $H_{\tau_{i}}$ with $\tau_{i}$ as the core curve of $H_{\tau_{i}}$, $H_{\tau_{i}}\cap D_{l}=\emptyset$ if $l\neq i$ or 0, and $H_{\tau_{i}}\cap D_{0}$ is the end interval of $H_{\tau_{i}}$ which contains the endpoint $\partial_{-}\tau_{i}$ of $\tau_{i}$. WLOG we may assume that that $D_{i}\cup H_{\tau_{i}}$ is 1-connected for $i=1,2,\dots,n$.
    Let
    $D:=D_{0}\cup(\cup_{i=1}^{n}\tilde{H}_{\tau_{i}}),\quad\tilde{H}_{\tau_{i}}:=H_{\tau_{i}}\cup D_{i},$
    $D$ is diffeomorphic to a 2-disc. Note that $Fix(\Phi)\subset\EuFrak{r}^{-1}(D)$, and $\EuFrak{r}^{-1}_{[-c,c]}(D)$ can be identified with a 1-connected Stein domain $W_{n}\subset\mathcal{W}_{n}$.
  2. (ii).
    A disjoint union (2g+b1)(2g+b-1) 1-handles with both ends attached to $\partial D_{0}$, missing all of the handles $H_{\tau_{i}}$ in (i):
    Denote these 1-handles as $H_{C_{j}}$ for $j=1,2,\dots,2g+b-1$, where $C_{j}$ is the core arc of $H_{C_{j}}$. Denote the boundary of $H_{C_{j}}$ as $\partial H_{C_{j}}=\partial_{+}H_{C_{j}}\cup\partial_{-}H_{C_{j}}$, $H_{C_{j}}$ is attached to $\partial D$ along $\partial_{\pm}H_{C_{j}}$, where $\partial_{\pm}H_{C_{j}}$ is the disjoint union of two short intervals containing $\partial C_{j}$. By handle sliding along $\partial D$ (and thinning $\tilde{H}_{\tau_{i}}$ and $H_{C_{j}}$ if necessary) we may assume that all $\partial H_{C_{j}}$ are attached to $\partial D_{0}$ and not touching any part of $\partial H_{\tau_{i}}$.
Then the union
$D\cup\Big(\sum_{j=1}^{2g+b-1}H_{C_{j}}\Big)=B_{0}$
up to a diffeomorphism.
Recall $\EuFrak{r}_{0}:h^{-1}(0)\to h^{-1}(0)/\Phi\cong B_{0}$ the standard projection. Fix a $\Phi$-invariant $\omega$-compatible Riemannian metric on $M$ and let $\nabla h$ denote the gradient vector field of the moment map $h$ with respect to the metric. $M$ deformation retracts to the hypersurface $h^{-1}(0)$ via flows of gradients $\pm\nabla h$ of $h$, therefore $H_{i}(M,\mathbb{Z})\cong H_{i}(h^{-1}(0),\mathbb{Z})$ for all $i$.
Now $h^{-1}(0)=\EuFrak{r}_{0}^{-1}(B_{0})$. As $\Phi$ acts freely on $\EuFrak{r}_{0}^{-1}(B_{0})\setminus Fix(\Phi)$, we have the following results.
  1. (i).
    $\EuFrak{r}_{0}^{-1}(\tau_{i})\subset h^{-1}(0)$ is Lagrangian disc centered at $p_{i}$, $1\leq i\leq n$. If we extend $\tau_{i}$ to an embedded arc $\tilde{\tau}_{i}\subset D$ with $q_{i}$ and $q_{0}$ as endpoints, then $\tilde{S}_{i}:=\EuFrak{r}_{0}^{-1}(\tilde{\tau}_{i})$ is a Lagrangian 2-sphere with points $p_{i}$ and $p_{0}$ as poles for $i=1,2,\dots n$. These $n$ spheres are homologically independent in $\EuFrak{r}^{-1}_{0}(D)\subset h^{-1}(0)$. Indeed $\EuFrak{r}^{-1}_{0}(D)$ is homotopic to the bouquet $\cup_{i=1}^{n}\tilde{S}_{i}$ of $n$ 2-spheres.
  2. (ii).
    $\EuFrak{r}_{0}^{-1}(C_{j})\subset h^{-1}(0)$ is Lagrangian cylinder with boundary attached to $\EuFrak{r}^{-1}_{0}(\partial D)$. If we extend $C_{j}$ to a simple closed curve $\tilde{C}_{j}\subset D\setminus\{q_{i}\mid i=0,\dots n\}$ then we get an embedded Lagrangian torus $\tilde{T}_{j}\subset h^{-1}(0)=\EuFrak{r}^{-1}_{0}(B_{0})$ for $j=1,2,\dots,2g+b-1$. These curves $\tilde{C}_{j}$, $j=1,\dots,2g+b-1$, together represent a basis of the homology group $H_{1}(h^{-1}(0))=H_{1}(B,\mathbb{Z})\cong\mathbb{Z}^{2g+b-1}$. As $\Phi$-orbits in $h^{-1}(0)$ contracts to points in $Fix(\Phi)$, $H_{1}(h^{-1}(0),\mathbb{Z})\cong H_{1}(B,\mathbb{Z})$ and is generated by a lifting of the curves $\tilde{C}_{j}$ in $\tilde{T_{j}}$, $j=1,\dots,2g+b-1$. The tori $\tilde{T}_{j}$’s are also independent on $H_{2}(h^{-1}(0),\mathbb{Z})$. Also $\tilde{T_{j}}$’s and $\tilde{S}_{i}$’s are pairwise linearly independent in $H_{2}(h^{-1}(0),\mathbb{Z})$, and forming a basis for $H_{2}(h^{-1}(0),\mathbb{Z})\cong\mathbb{Z}^{2g+b-1+n}$. Therefore
[…]
    In particular the homology groups of $(M,\omega,\Phi)$ is completely determined by the triple $(k=n+1,g,b)$ with $k\geq 0$, $g\geq 0$, $b\geq 1$, and $g+b\geq 2$.
Parr 2: Stein structure on $(M,\omega=d\lambda,\Phi)$
Below we study the existence of Stein structure on $(M,\omega=d\lambda,\Phi)$. Recall the following theorem due to Eliashberg [2] (see also [3] Theorem 1.3 and [4] Theorem 2.3) about the existence of a Stein structure on a 4-manifold via handle attaching criteria, which can be stated as the following after Gompf ([4] Theorem 2.3):
Theorem 6.1 (Eliashberg.).
An oriented 4-manifold admits a Stein structure if and only if it is diffeomorphic to the interior of a handlebody whose handles all have index $\leq$ 2, and for which each 2-handle is attached along a Legendrian knot (in the standard contact structure on the relevant boundary 3-manifold) with framing obtained from the contact framing by adding one left twist.
Here a 4-dimensional 2-handle H is diffeomorphic to a product space modeled on $D^{\prime}\times D^{\prime\prime}_{\epsilon}\subset\mathbb{R}^{2}_{x_{1},x_{2}}\times i\mathbb{R}^{2}_{y_{1},y_{2}}\subset\mathbb{C}^{2}$, where $D^{\prime}=\{x_{1}^{2}+x_{2}^{2}\leq 1\}$, called the (Lagrangian) core disc of H, and $D^{\prime\prime}=\{y_{1}^{2}+y_{2}^{2}<\epsilon\}$ for some $\epsilon>0$. H can be viewed as a trivial $D^{2}$-bundle over $D^{\prime}$. The basis normal vector fields $\{\partial_{y_{1}},\partial_{y_{2}}\}$ over $D^{\prime}$, when restricted to $\partial D^{\prime}$, gives the canonical framing of the symplectic normal bundle of $\partial D^{\prime}\subset\textsf{H}$. In particular, we can take either $\partial_{y_{1}}$ or $\partial_{y_{2}}$ as the framing of the $D^{2}$-bundle over $\partial D^{\prime}$.
Below we will show that the handle decomposition of $B_{0}=D_{0}\cup(\cup_{i=1}^{n}\tilde{H}_{\tau_{i}})\cup(\cup_{j}H_{C_{j}})$ lifts to a corresponding handle decomposition of $M$ as a union of handles
$M=\textsf{H}_{0}\cup(\cup_{i}\textsf{H}_{V_{i}})\cup(\cup_{j}(\textsf{H}_{\gamma_{j}}\cup\textsf{H}_{U_{j}}))$
where $\textsf{H}_{0}=\EuFrak{r}^{-1}(D_{0})$ is a 0-handle, $\textsf{H}_{V_{i}}$ are 2-handles associated to $\tilde{H}_{\tau_{i}}$, and for each $j$, the pair $(\textsf{H}_{\gamma_{j}},\textsf{H}_{U_{j}})$, where $\textsf{H}_{\gamma_{j}}$ is a 1-handle and $\textsf{H}_{U_{j}}$ is a 2-handle, is associated to $H_{C_{j}}$. Moreover, all the 2-handles $\textsf{H}_{V_{i}}$ and $\textsf{H}_{U_{j}}$ of $M$ satisfy the contact framing condition, hence $M$ admits the structure of a Stein manifold.
Case 1: $\textsf{H}_{V_{i}}$. Denote by $V_{i}:=\EuFrak{r}^{-1}_{0}(\tau_{i})$ the Lagrangian disc in $h^{-1}(0)$ with the boundary circle $\partial V_{i}$ attached to $\EuFrak{r}^{-1}_{0}(\partial D_{0})$. $V_{i}$ is the core disc of a 4-dimensional handle $\textsf{H}_{V_{i}}$ which can be identified with a subdomain of the cotangent bundle over $V_{i}$. We may assume that these handles are mutually disjoint when attached to $\EuFrak{r}^{-1}(\partial D_{0})$.
Pick any $V_{i}$ and denote it as $V$, also denote $\tau:=\tau_{i}=\EuFrak{r}(V)$, $p:=p_{i}$. We may parametrize $V$ as
$V=\{(x_{1},x_{2})\mid x_{1}^{2}+x_{2}^{2}\leq 1\},\quad p=(0,0)\in Fix(\Phi)$
and identify
$X|_{V}=x_{1}\partial_{x_{2}}-x_{2}\partial_{x_{1}}.$
Let $(y_{1},y_{2})$ be the fiber coordinates of the cotangent bundle $T^{*}V$ dual to $(x_{1},x_{2})$. Then $V$ can be identified with the core disc of the 4-dimensional 2-handle
$\textsf{H}_{V}:=\EuFrak{r}^{-1}_{0}(H_{\tau})\cong V\times D^{2}_{y}\subset T^{*}L,\quad D^{2}_{y}=\{(y_{1},y_{2})\mid y_{1}^{2}+y_{2}^{2}<\epsilon^{2}\}$
for some $0<\epsilon$, with $\partial_{-}\textsf{H}_{V}:=\partial V\times D^{2}_{\epsilon}$ attached to $\EuFrak{r}^{-1}(\partial D_{0})\subset\partial W_{0}$, where $W_{0}=\EuFrak{r}^{-1}(D_{0})$.
Identify $\textsf{H}_{V}$ as a subdomain of $\mathbb{C}^{2}$. Then on $\textsf{H}_{V}$ we may take $\omega=d\lambda$, with $\lambda=\frac{1}{2}\sum_{i=1}^{2}(x_{i}dy_{i}-y_{i}dx_{i})$, identify the $\Phi$-action with the standard $\Theta$-action on $\mathbb{C}^{2}$, and take the Euclidean metric as the $\Phi$-invariant $\omega$-compatible Riemannian metric on $\textsf{H}_{V}$. Then
$\nabla h|_{\partial V}=-x_{1}\partial_{y_{2}}+x_{2}\partial_{y_{1}}.$
Note that both $X$ and $\nabla h$ are tangent to $\EuFrak{r}^{-1}(\partial D_{0})$ and hence to $\partial_{-}\textsf{H}_{V}$. The 1-form $\lambda$ restricts to a contact 1-form $\alpha$ on $\partial V\times D^{2}_{\epsilon}$. $\lambda(X)=0$ along $\partial V$ so $\partial V$ is a Legendrian curve to the contact form. As the outward normal vector field to $\EuFrak{r}^{-1}(D_{0})=W_{0}$ near $\partial V$ is $x_{1}\partial_{x_{1}}-x_{2}\partial_{x_{2}}\subset\ker\lambda|_{\partial V}$, the contact plane field along $\partial V$ is spanned by the symplectic pair $\{\nabla h,X\}$. So the contact framing along $\partial V$ (oriented by $X$) is given by the vector field $\nabla h|_{\partial V}=-x_{1}\partial_{y_{2}}+x_{2}\partial_{y_{1}}$.
Identify $V$ as the 0-section of the cotangent bundle $T^{*}V\supset\textsf{H}_{V}$. The normal bundle $N_{V}$ of $V\subset T^{*}V$ is trivial with $\{\partial_{y_{1}},\partial_{y_{2}}\}$ as the basis fields of $N_{V}$. $N_{V}$ restricted to $\partial V$ is the symplectic normal bundle $SN^{*}(\partial V)=\partial_{y_{1}}\wedge\partial_{y_{2}}|_{\partial V}$ of $\partial V$ with the natural trivialization/framing given by $\partial_{y_{1}}$.
$SN^{*}(\partial V)$ is then identified with the normal bundle $N_{\partial V/\partial W_{0}}$ of $\partial V\subset\partial W_{0}$ upon the attaching of $\partial_{-}\textsf{H}_{V}=\partial V\times D^{2}_{y}$ to $\partial W_{0}$ along $\partial V$. As we go once along $\partial V$ the contact framing $\nabla h$ makes one positive full rotation relative to the canonical framing $\partial_{y_{1}}$. In other words, the canonical framing $\partial_{y_{1}}$ along $\partial V$ is obtained from the contact framing by adding one left twist. So the natural framing along $\partial V$ given by $\partial_{y_{1}}$ is 1-1 relative to the contact framing along $\partial V$, which meets with the Stein condition on the attachment of 4-dimensional 2-handles.
The above result on the framing criteria along $\partial V\subset\partial_{-}\textsf{H}$ apply to $V_{i}$ for all $i=1,2,\dots,n$ as well. Each of the 2-handlebodies $\textsf{H}_{V_{i}}$ has its core disc $V_{i}$ attached along a Legendrian knot (in the standard contact structure on the relevant boundary 3-manifold) with framing obtained from the contact framing by adding one left twist. In particular, $\textsf{H}_{0}\cup(\cup_{i}\textsf{H}_{V_{i}})$ admits a Stein structure by Theorem 6.1. $\Box$
Case 2: $\textsf{H}_{C_{j}}$. Pick any one of the 2-dimensional handles $H_{C_{j}}$ and denote it as $H_{C}$, where $C$ is the core curve of $H_{C}$. We will see that $\EuFrak{r}^{-1}(H_{C})$ can be identified with the union of a 4-dimensional 1-handle $\textsf{H}_{\gamma}$ with core curve $\gamma\subset h^{-1}(0)$ and a 4-dimensional 2-handle $\textsf{H}_{U}$ with core disc $U\subset h^{-1}(0)$, and $\partial U$ is attached to $\partial\textsf{H}_{\gamma}\cap h^{-1}(0)$.
The preimage $L:=\EuFrak{r}^{-1}_{0}(C)\subset h^{-1}(0)$ is a Lagrangian annulus $L\cong I\times S^{1}$ foliated by $G$-orbits generated by the Hamiltonian vector field $X$. We may parametrize $L$ as
$L=\{(x_{1},x_{2})\in\mathbb{R}\times\mathbb{R}/2\pi\mathbb{Z}\mid-1\leq x_{1}\leq 1,x_{2}\in\mathbb{R}\}$
so that $X=\partial_{x_{2}}$, and each integral curve of $\partial_{x_{1}}$ is a lifting of $C$ in $L$. The boundary of $L$, which is the pair of $\Phi$-orbits given by the equations $x_{1}=\pm 1$ is attached to $\partial W_{n}$. Let
$\gamma:=\{(x_{1},x_{2})\mid-1\leq x_{1}\leq 1,\ x_{2}=0\}\subset L.$
$\gamma$ is a lifting of $C$ in $h^{-1}(0)$.
Let $(y_{1},y_{2})$ be the fiber coordinates of the cotangent bundle $T^{*}L$ dual to $(x_{1},x_{2})$. Then $\gamma$ can be identified with the core curve of the 4-dimensional 1-handle (a product of four intervals)
$\textsf{H}_{\gamma}:=\EuFrak{r}^{-1}(H_{C})\cong\gamma\times I^{3}\subset T^{*}L,\quad I^{3}=I_{|y_{1}|<\epsilon}\times I_{|x_{2}|<\epsilon}\times I_{y_{2}}$
for some $0<\epsilon\ll 1$, with $\partial_{-}\textsf{H}_{\gamma}:=\partial\gamma\times I^{3}$ attached to $\partial M_{0}$. Here $y_{2}$ parametrizes integral curves of $\nabla h$, the negative gradient vector field of $h$ with respect to some $G$-invariant $\omega$-compatible Riemannian metric on $M$. Let
$I_{x}:=\gamma\times I_{|x_{2}|<\epsilon}=\{(x_{1},x_{2})\mid-1\leq x_{1}\leq 1,\ -\epsilon<x_{2}<\epsilon\}\subset L.$
Up to a smoothing of the corners, the complement
$U:=L\setminus I_{x}=[-1,1]\times[\epsilon,2\pi-\epsilon]\subset L$
is a Lagrangian disc in $h^{-1}(0)$ with boundary attached to $\partial W_{n}\cap h^{-1}(0)$. Observe that $U$ is the core 2-disc of the handle
$\textsf{H}_{U}:=U\times I_{|y_{1}|<\epsilon}\times I_{y_{2}},$
with $\partial_{-}\textsf{H}_{U}=\partial U\times I_{|y_{1}|<\epsilon}\times I_{y_{2}}$ attached to $\partial_{+}\textsf{H}_{\gamma}=\gamma\times\partial I^{3}$ so that
$\textsf{H}_{\gamma}\cup\textsf{H}_{U}=T^{*}L\cap\{|y_{1}|<\epsilon\}.$
Orient $\partial U=\gamma_{-}\cup\sigma_{+}\cup\gamma_{+}\cup\sigma^{-}$ counterclockwise, where
(8) $\gamma_{-}$ $=\{-1\leq x_{1}\leq 1,\ x_{2}=\epsilon\},$ $\quad\dot{\gamma}_{-}$ $=\partial_{x_{1}},$
(9) $\sigma_{+}$ $=\{x_{1}=1,\ \epsilon\leq x_{2}\leq 2\pi-\epsilon\},$ $\quad\dot{\sigma}_{+}$ $=\partial_{x_{2}}=X,$
(10) $\gamma_{+}$ $=\{-1\leq x_{1}\leq 1,\ x_{2}=2\pi-\epsilon\},$ $\quad\dot{\gamma}_{+}$ $=-\partial_{x_{1}},$
(11) $\sigma_{-}$ $=\{x_{1}=-1,\ -\epsilon\leq x_{2}\leq 2\pi-\epsilon\},$ $\quad\dot{\sigma}_{-}$ $=-\partial_{x_{2}}=-X.$
Recall the $\Phi$-invariant primitive 1-form $\lambda$ of $\omega$. Observe that $\EuFrak{r}^{-1}(\partial D)\subset\partial W_{n}$ is a subdomain of $\partial W_{n}$, $\lambda$ restricted to $\partial W_{n}$ is a contact 1-form near $\partial L$, whose contact structure along $\partial L$ is spanned by the symplectic pair $\nabla h$ and $X$. So along $\sigma_{+}$ the contact structure is $X\wedge(-\nabla h)=\partial_{x_{2}}\wedge\partial_{y_{2}}$, and along $\sigma_{-}$ the contact structure is $X\wedge\nabla h=-\partial_{x_{2}}\wedge(-\partial_{y_{2}})$. Along the boundary $(\partial_{+}\textsf{H}_{\gamma})\cap L=\gamma_{-}\cup\gamma_{+}$ the vector field $X$ is normal to $\partial_{+}\textsf{H}_{\gamma}$, hence along $\gamma_{-}$ the contact structure on $\partial_{+}\textsf{H}_{\gamma}$ is $\partial_{x_{1}}\wedge\partial_{y_{1}}$, spanned by the symplectic pair $\dot{\gamma}_{-}=\partial_{x_{1}}$ and $\partial_{y_{1}}$. Similarly along $\gamma_{+}$ the contact structure on $\partial_{+}\textsf{H}_{\gamma}$ is $(-\partial_{x_{1}})\wedge(-\partial_{y_{1}})$, spanned by the symplectic pair $\dot{\gamma}_{+}=-\partial_{x_{1}}$ and $\partial_{y_{1}}$.
By smoothing the corners of $U\subset L$ we may identify $U$ with a unit disc $D_{U}=\{x_{1}^{2}+x_{2}^{2}\leq 1\}$ with boundary $\partial D_{U}$ attached to $\partial(\textsf{H}_{\gamma}\cup W_{n})$, so that along the legendrian circle $\ell:=\partial D_{U}=S^{1}=\mathbb{R}/2\pi\mathbb{Z}$,
$\begin{split}\dot{\ell}(\theta)&=\cos\theta\partial_{x_{2}}-\sin\theta\partial_{x_{1}},\quad\ell(0)=(1,0)\\ &=x_{1}\partial_{x_{2}}-x_{2}\partial_{x_{1}},\end{split}$
and the vector field
$\cos\theta\partial_{y_{2}}-\sin\theta\partial_{y_{1}}=x_{1}\partial_{y_{2}}-x_{2}\partial_{y_{1}}$
is the contact framing along $\ell$.
Similar to Case 1, the framing of the symplectic normal bundle $SN^{*}(\partial U)$ (which is spanned by the basis vector fields $\{\partial_{y_{1}},\partial_{y_{2}}\}$) is given by $\partial_{y_{1}}$. As we go once along $\ell$ the contact framing makes one positive full rotation relative to $\partial_{y_{1}}$. So the canonical framing $\partial_{y_{1}}$ is 1-1 relative to the contact framing along $\ell$. This result applies to all $\textsf{H}_{C_{i}}$. Combining the result from Case 1 we conclude that $M$ admits a Stein structure following Theorem 6.1. This completes the proof of Theorem 1.11. $\Box$